Jefferson Lab HALL A Logbook for November 10 - Full Listing

  • Main INDEX

    Tuesday

  • 342144 11/30/10 23:59 hyao Swing shift summary
  • 342145 11/30/10 23:59 adaq Beam Time Accounting
  • 342143 11/30/10 23:44 Carlos Bleedthrough measurement
  • 342142 11/30/10 23:39 hyao hall a tool and bpm screen
  • 342141 11/30/10 23:28 a-onl Start_Run_9128,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 3uA
  • 342140 11/30/10 23:23 mhuang Cryotarget screens
  • 342139 11/30/10 23:21 a-onl End_of_Run_9127
  • 342138 11/30/10 23:21 pzhu roc,rc9127
  • 342137 11/30/10 23:20 pzhu scaler 9127
  • 342136 11/30/10 22:54 pzhu analysis normal for run 9127
  • 342135 11/30/10 22:53 Carlos Slot 0 of 2006 frame
  • 342134 11/30/10 22:43 hyao The pic is after fix
  • 342133 11/30/10 22:28 pzhu DVCS high votage problem
  • 342132 11/30/10 22:28 hyao dvcs 2006 hv tripped
  • 342131 11/30/10 22:27 a-onl Start_Run_9127,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 3uA
  • 342130 11/30/10 22:27 a-onl End_of_Run_9122
  • 342129 11/30/10 22:23 a-onl End_of_Run_9126
  • 342128 11/30/10 22:23 pzhu scaler9126
  • 342127 11/30/10 22:23 pzhu roc,rc9126
  • 342126 11/30/10 21:36 a-onl Start_Run_9126,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 3uA
  • 342125 11/30/10 21:15 pzhu scaler9125
  • 342124 11/30/10 21:15 a-onl End_of_Run_9125
  • 342123 11/30/10 21:15 pzhu 9125roc,rc
  • 342122 11/30/10 21:13 Carlos Kin3-high rates
  • 342121 11/30/10 21:07 Carlos Golden run updated to 9124
  • 342120 11/30/10 21:01 a-onl Start_Run_9125,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 3uA
  • 342119 11/30/10 21:00 pzhu scaler9124
  • 342118 11/30/10 21:00 a-onl End_of_Run_9124
  • 342117 11/30/10 21:00 pzhu 9124roc,rc
  • 342116 11/30/10 21:00 mhuang Cryotarget screens
  • 342115 11/30/10 20:15 a-onl Start_Run_9124,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 3uA
  • 342114 11/30/10 20:11 pzhu scaler 9123
  • 342113 11/30/10 20:11 pzhu roc, runcontrol 9123
  • 342112 11/30/10 20:10 a-onl End_of_Run_9123
  • 342111 11/30/10 19:28 auto_snap Radiation and Ozone snapshot
  • 342110 11/30/10 19:28 auto_snap Miscellaneous snapshot
  • 342109 11/30/10 19:28 auto_snap Hall A Tools snapshot
  • 342108 11/30/10 19:28 auto_snap Compton Polarimeter snapshot
  • 342107 11/30/10 19:27 auto_snap Beam Position and Current snapshot
  • 342106 11/30/10 19:26 a-onl Start_Run_9123,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 2uA
  • 342105 11/30/10 19:23 a-onl Start_Run_9122,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=2,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production,current 2uA
  • 342104 11/30/10 19:22 mhuang target moved to loop2 15cm
  • 342103 11/30/10 19:17 a-onl End_of_Run_9121
  • 342102 11/30/10 18:46 a-onl Start_Run_9121,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=optics
  • 342101 11/30/10 18:33 hyao Compton chicane is off
  • 342100 11/30/10 18:30 mhuang Target moved to optics
  • 342099 11/30/10 18:20 a-onl End_of_Run_9120
  • 342098 11/30/10 18:09 hyao MCC ask to turn off compton chicane
  • 342097 11/30/10 17:50 a-onl Start_Run_9120,Run_type=Other,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger efficiency run
  • 342096 11/30/10 17:48 a-onl End_of_Run_9119
  • 342095 11/30/10 17:48 a-onl End_of_Run_9118
  • 342094 11/30/10 17:47 Carlos Rates run 9118
  • 342093 11/30/10 17:46 Carlos Coda screens run 9118
  • 342092 11/30/10 17:22 efuchey Compton rates for today
  • 342091 11/30/10 17:06 a-onl Start_Run_9118,Run_type=Production,target_type=LH2,beam_e0=5.5,beam_raster=off,beam_current=3,leftarm_p=2.591,leftarm_theta=21.49,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, kin-3 high
  • 342090 11/30/10 16:58 mhuang Target moved to Loop2 15cm
  • 342089 11/30/10 16:52 hyao harp Scan result from MCC
  • 342088 11/30/10 16:36 musson FIlter Schematic
  • 342087 11/30/10 16:33 musson Downstream BCM
  • 342086 11/30/10 16:27 mhuang Cryotarget screens
  • 342085 11/30/10 16:09 efuchey QWP scan for 11/30/2010
  • 342083 11/30/10 16:01 A. Deur Day shift summary
  • 342084 11/30/10 16:00 adaq Beam Time Accounting
  • 342082 11/30/10 15:16 a-onl End_of_Run_9117
  • 342081 11/30/10 15:15 a-onl End_of_Run_9115
  • 342080 11/30/10 15:07 a-onl Start_Run_9117,Run_type=Pedestal,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ARS Pedestal HV Off
  • 342079 11/30/10 15:04 a-onl End_of_Run_9116
  • 342078 11/30/10 15:03 a-onl Start_Run_9116,Run_type=Pedestal,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ARS Pedestal HV Off
  • 342077 11/30/10 14:58 a-onl Start_Run_9115,Run_type=Pedestal,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ARS Pedestal HV Off
  • 342076 11/30/10 14:58 a-onl End_of_Run_9114
  • 342075 11/30/10 14:54 Carlos BCM calibration
  • 342074 11/30/10 14:53 a-onl Start_Run_9114,Run_type=Pedestal,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ARS Pedestal HV Off
  • 342073 11/30/10 14:52 a-onl Start_Run_9113,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test Ben's firmware - LED on block 8 775 V
  • 342072 11/30/10 13:40 a-onl End_of_Run_9113
  • 342071 11/30/10 13:40 A. Camsonne Production firmware for trigger loaded
  • 342070 11/30/10 13:39 A. Camsonne Restored HV on block 8 to 533.2V
  • 342069 11/30/10 13:29 hyde BCM VTOF double pulsing
  • 342068 11/30/10 13:19 a-onl End_of_Run_9112
  • 342067 11/30/10 13:08 A. Deur Hall A tools snapshot for the new kinematics
  • 342066 11/30/10 12:59 a-onl Start_Run_9110,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test Ben's firmware - LED on block 8 775 V
  • 342065 11/30/10 12:58 a-onl End_of_Run_9109
  • 342064 11/30/10 12:09 A. Deur setting LHRS to 2.591 GeV
  • 342063 11/30/10 11:42 hyde Run 9105 livetime beginning of run
  • 342062 11/30/10 11:30 A. Camsonne Stop amplitude
  • 342061 11/30/10 11:27 auto_snap Radiation and Ozone snapshot
  • 342060 11/30/10 11:27 auto_snap Miscellaneous snapshot
  • 342059 11/30/10 11:27 auto_snap Hall A Tools snapshot
  • 342058 11/30/10 11:27 auto_snap Compton Polarimeter snapshot
  • 342057 11/30/10 11:27 auto_snap Beam Position and Current snapshot
  • 342056 11/30/10 11:05 hyde Beginning of run9105 TS livetime
  • 342055 11/30/10 10:59 a-onl Start_Run_9109,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test Ben's firmware - LED on block 8 775 V
  • 342054 11/30/10 10:52 A. Deur target snapshots +L1 sensor
  • 342053 11/30/10 10:49 a-onl End_of_Run_9108
  • 342052 11/30/10 10:47 A. Camsonne Changed HV of block 8 for LED tests
  • 342051 11/30/10 09:56 hyde Run 9105 livetime deadtime
  • 342050 11/30/10 09:35 a-onl End_of_Run_9107
  • 342049 11/30/10 09:24 a-onl Start_Run_9107,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test Ben's firmware - no validation on ARS
  • 342048 11/30/10 09:24 A. Camsonne FPGA Trigger loaded with test firmware
  • 342047 11/30/10 09:23 a-onl End_of_Run_9106
  • 342046 11/30/10 09:23 a-onl End_of_Run_9102
  • 342043 11/30/10 08:13 A. Deur Controled access requested. Moved target to empty position
  • 342042 11/30/10 08:09 A. Deur target snapshots + chamber vacuum
  • 342041 11/30/10 07:59 adaq Beam Time Accounting
  • 342040 11/30/10 07:55 rakithab Shift Summary
  • 342039 11/30/10 07:36 rakithab No beam during the rest of the OWL shift
  • 342038 11/30/10 06:52 rakithab Beam not available during last 1 hour
  • 342037 11/30/10 06:06 hoyoung Target screens
  • 342036 11/30/10 05:59 rakithab run number should be 9105
  • 342035 11/30/10 05:59 rakithab run number should be 9105
  • 342034 11/30/10 05:53 a-onl End_of_Run_9105
  • 342033 11/30/10 05:53 rakithab Hall A Scaler data during the run 9104
  • 342032 11/30/10 05:53 rakithab CODA status during the run 9104
  • 342031 11/30/10 05:40 a-onl Start_Run_9105,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Kin1 Low 3 pass
  • 342030 11/30/10 05:39 a-onl End_of_Run_9104
  • 342029 11/30/10 05:35 rakithab Hall A Scaler data during the run 9104
  • 342028 11/30/10 05:34 rakithab CODA status during the run 9104
  • 342027 11/30/10 04:41 a-onl Start_Run_9104,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Kin1 Low 3 pass
  • 342026 11/30/10 04:39 a-onl End_of_Run_9103
  • 342025 11/30/10 04:35 rakithab CODA status during the run 9103
  • 342024 11/30/10 04:35 rakithab Hall A Scaler data during the run 9103
  • 342023 11/30/10 04:03 hoyoung Target screens
  • 342022 11/30/10 03:39 A. Camsonne Production trigger fimware loaded - Run 9103 production
  • 342021 11/30/10 03:38 a-onl Start_Run_9103,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production - Kin1 Low 3 pass
  • 342020 11/30/10 03:34 A. Camsonne Run 9102 test new logic module/ trigger firmware - no prescales
  • 342019 11/30/10 03:33 A. Camsonne Test new module logic no prescale
  • 342018 11/30/10 03:33 a-onl Start_Run_9102,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70 , pedestal for trigger hallA1x600.txt with beam, no prescales
  • 342017 11/30/10 03:31 rakithab Beam recovery completed
  • 342016 11/30/10 03:31 A. Camsonne Test run with new trigger version Run 9101
  • 342015 11/30/10 03:30 a-onl End_of_Run_9101
  • 342014 11/30/10 03:30 A. Camsonne Scalers with new logic modules
  • 342013 11/30/10 03:29 a-onl Start_Run_9101,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70 , pedestal for trigger hallA1x600.txt with beam
  • 342012 11/30/10 03:27 hoyoung Target moved to 15cm LH2
  • 342011 11/30/10 03:27 auto_snap Radiation and Ozone snapshot
  • 342010 11/30/10 03:27 auto_snap Miscellaneous snapshot
  • 342009 11/30/10 03:27 auto_snap Hall A Tools snapshot
  • 342008 11/30/10 03:27 auto_snap Compton Polarimeter snapshot
  • 342007 11/30/10 03:27 auto_snap Beam Position and Current snapshot
  • 342006 11/30/10 02:51 A. Camsonne Config TrigDVCSBen
  • 342005 11/30/10 02:39 a-onl End_of_Run_9100
  • 342004 11/30/10 02:34 a-onl Start_Run_9100,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70 , pedestal for trigger hallA1x600.txt
  • 342003 11/30/10 02:29 a-onl Start_Run_9099,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70
  • 342002 11/30/10 01:53 a-onl Start_Run_9097,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70
  • 342001 11/30/10 01:52 a-onl End_of_Run_9096
  • 342000 11/30/10 01:50 a-onl Start_Run_9096,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70
  • 341999 11/30/10 01:20 a-onl Start_Run_9095,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70
  • 341998 11/30/10 01:19 hoyoung Target screens
  • 341997 11/30/10 01:13 a-onl End_of_Run_9094
  • 341996 11/30/10 01:13 a-onl Start_Run_9094,Run_type=LED,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Ben's firmware, Delay 0x70
  • 341995 11/30/10 01:00 a-onl Start_Run_9093,Run_type=Other,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Junk
  • 341994 11/30/10 00:57 a-onl End_of_Run_9092
  • 341993 11/30/10 00:56 a-onl Start_Run_9092,Run_type=Other,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Junk
  • 341992 11/30/10 00:55 a-onl End_of_Run_9091
  • 341991 11/30/10 00:54 a-onl Start_Run_9091,Run_type=Other,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Junk
  • 341990 11/30/10 00:53 a-onl End_of_Run_9090
  • 341989 11/30/10 00:50 a-onl Start_Run_9090,Run_type=Other,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Junk
  • 341988 11/30/10 00:49 a-onl Start_Run_9089,Run_type=Other,target_type=Empty loop,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test trigger - Junk
  • 341987 11/30/10 00:46 A. Camsonne adaql2 rebooted
  • Monday

  • 341986 11/29/10 23:59 a-onl End_of_Run_9088
  • 341985 11/29/10 23:59 a-onl End_of_Run_9083
  • 341984 11/29/10 23:32 adaq Beam Time Accounting
  • 341983 11/29/10 23:31 Xiaohui Swing Shift Summary
  • 341982 11/29/10 23:31 mhuang Cryotarget screens
  • 341981 11/29/10 23:30 Xiaohui Hall in controlled access
  • 341980 11/29/10 21:47 Kalyan No beam in Hall-A until 3am due to RF separator problem
  • 341979 11/29/10 20:40 mhuang Move to empty target
  • 341978 11/29/10 20:15 mhuang Cryotarget screens
  • 341977 11/29/10 19:27 auto_snap Radiation and Ozone snapshot
  • 341976 11/29/10 19:27 auto_snap Miscellaneous snapshot
  • 341975 11/29/10 19:27 auto_snap Hall A Tools snapshot
  • 341974 11/29/10 19:27 auto_snap Compton Polarimeter snapshot
  • 341973 11/29/10 19:27 auto_snap Beam Position and Current snapshot
  • 341972 11/29/10 19:13 Xiaohui Beam off
  • 341971 11/29/10 18:41 pzhu 9087 scaler
  • 341970 11/29/10 18:41 a-onl End_of_Run_9087
  • 341969 11/29/10 18:41 pzhu 9087 roc,rc
  • 341968 11/29/10 18:09 pzhu scaler data screenshot for run 9086
  • 341967 11/29/10 18:09 pzhu roc,rc screenshot for run 9086
  • 341966 11/29/10 18:08 a-onl Start_Run_9087,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341965 11/29/10 18:08 a-onl End_of_Run_9086
  • 341964 11/29/10 16:59 a-onl Start_Run_9086,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341963 11/29/10 16:54 pzhu scaler data window for run 9085
  • 341962 11/29/10 16:54 pzhu roc,rc screenshot for run 9085
  • 341961 11/29/10 16:53 a-onl End_of_Run_9085
  • 341960 11/29/10 16:32 mhuang cryotarget screens
  • 341959 11/29/10 16:29 mhuang note of compton run 24354
  • 341958 11/29/10 16:27 a-onl Start_Run_9085,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341957 11/29/10 16:26 a-onl End_of_Run_9084
  • 341956 11/29/10 16:25 pzhu scaler data for run 9084
  • 341955 11/29/10 16:25 pzhu ROC,Runcontrol window for run 9084
  • 341954 11/29/10 16:17 jpchen Day shift summary
  • 341953 11/29/10 16:15 adaq Beam Time Accounting
  • 341952 11/29/10 16:04 jpchen cryotarget screens
  • 341951 11/29/10 15:15 a-onl Start_Run_9084,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341950 11/29/10 15:14 a-onl End_of_Run_9082
  • 341949 11/29/10 15:13 n.muangma scaler 9083
  • 341948 11/29/10 15:13 n.muangma coda roc 9083
  • 341947 11/29/10 15:05 Carlos DVCS IOC rebooted
  • 341946 11/29/10 14:42 a-onl Start_Run_9083,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341945 11/29/10 14:33 jpchen Compton HV trip
  • 341944 11/29/10 14:17 n.muangma scaler 9082
  • 341943 11/29/10 14:17 n.muangma 9082 coda roc
  • 341942 11/29/10 13:12 a-onl Start_Run_9082,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341941 11/29/10 13:11 a-onl End_of_Run_9081
  • 341940 11/29/10 13:10 n.muangma scaler 9081
  • 341939 11/29/10 13:09 n.muangma coda rocs 9081
  • 341938 11/29/10 11:53 a-onl Start_Run_9081,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341937 11/29/10 11:53 a-onl End_of_Run_9080
  • 341936 11/29/10 11:53 jpchen beam back
  • 341935 11/29/10 11:45 n.muangma scaler 9080
  • 341934 11/29/10 11:44 n.muangma coda rocs 9080
  • 341933 11/29/10 11:39 jpchen beam off
  • 341932 11/29/10 11:27 auto_snap Radiation and Ozone snapshot
  • 341931 11/29/10 11:27 auto_snap Miscellaneous snapshot
  • 341930 11/29/10 11:26 auto_snap Hall A Tools snapshot
  • 341929 11/29/10 11:26 auto_snap Compton Polarimeter snapshot
  • 341928 11/29/10 11:26 auto_snap Beam Position and Current snapshot
  • 341927 11/29/10 10:46 a-onl Start_Run_9080,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341926 11/29/10 10:09 a-onl End_of_Run_9079
  • 341925 11/29/10 10:08 n.muangma 9079 scaler
  • 341924 11/29/10 10:08 jpchen Beam off
  • 341923 11/29/10 10:08 n.muangma coda rocs 9079
  • 341922 11/29/10 09:21 jpchen Beam back.
  • 341921 11/29/10 09:21 a-onl Start_Run_9079,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341920 11/29/10 09:13 jpchen Beam away to set up spliter.
  • 341919 11/29/10 09:12 a-onl End_of_Run_9078
  • 341918 11/29/10 09:07 jpchen MCC will work on beam spliter in ~ 1 hour
  • 341917 11/29/10 09:03 a-onl Start_Run_9078,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341916 11/29/10 09:03 a-onl End_of_Run_9077
  • 341915 11/29/10 09:02 n.muangma scaler
  • 341914 11/29/10 09:01 n.muangma coda, roc, 9077
  • 341913 11/29/10 08:06 jroche Owl shift summary
  • 341912 11/29/10 08:03 adaq Beam Time Accounting
  • 341911 11/29/10 07:41 a-onl Start_Run_9077,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341910 11/29/10 07:40 a-onl End_of_Run_9076
  • 341909 11/29/10 07:39 itard entries end of run 9076
  • 341908 11/29/10 07:09 jroche Compton HV trip
  • 341907 11/29/10 06:31 a-onl Start_Run_9076,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341906 11/29/10 06:31 a-onl End_of_Run_9075
  • 341905 11/29/10 06:30 itard entries end of run 9075
  • 341904 11/29/10 06:25 jroche cerenkov 2 gain back to normal
  • 341903 11/29/10 05:33 a-onl Start_Run_9075,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341902 11/29/10 05:29 a-onl End_of_Run_9074
  • 341901 11/29/10 05:28 a-onl End_of_Run_9062
  • 341900 11/29/10 05:28 itard entries end of run 9074
  • 341899 11/29/10 04:28 a-onl Start_Run_9074,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341898 11/29/10 04:27 a-onl End_of_Run_9073
  • 341897 11/29/10 04:26 itard entries end of run 9073
  • 341896 11/29/10 03:56 jroche changed coefficents for charge compuationin end of run entries
  • 341895 11/29/10 03:41 jroche Scaler results for the BCM calibration performed on November 24
  • 341894 11/29/10 03:26 auto_snap Radiation and Ozone snapshot
  • 341893 11/29/10 03:26 auto_snap Miscellaneous snapshot
  • 341892 11/29/10 03:26 auto_snap Hall A Tools snapshot
  • 341891 11/29/10 03:26 auto_snap Compton Polarimeter snapshot
  • 341890 11/29/10 03:26 auto_snap Beam Position and Current snapshot
  • 341889 11/29/10 03:24 a-onl Start_Run_9073,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341888 11/29/10 03:16 a-onl End_of_Run_9072
  • 341887 11/29/10 03:13 jroche target to LH2 15 cm
  • 341886 11/29/10 03:09 a-onl Start_Run_9072,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, Kin1 Low Energy
  • 341885 11/29/10 03:08 jroche MCC reports trouble with the RF separator
  • 341884 11/29/10 03:06 a-onl End_of_Run_9071
  • 341883 11/29/10 03:05 itard entries end of run 9071
  • 341882 11/29/10 02:47 a-onl Start_Run_9071,Run_type=Production,target_type=Optics: multi 1mm C foils,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=efficiency run, Kin1 Low Energy
  • 341881 11/29/10 02:43 a-onl End_of_Run_9070
  • 341880 11/29/10 02:41 itard entries end of run 9070
  • 341879 11/29/10 02:29 a-onl Start_Run_9070,Run_type=Production,target_type=Optics: multi 1mm C foils,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=efficiency run, Kin1 Low Energy
  • 341878 11/29/10 02:27 jroche target to Optics
  • 341877 11/29/10 02:23 a-onl End_of_Run_9069
  • 341876 11/29/10 02:22 itard entires end of run 9069
  • 341875 11/29/10 02:02 jroche Gain drift for cerenkov 2
  • 341874 11/29/10 01:54 a-onl Start_Run_9069,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=efficiency run, Kin1 Low Energy
  • 341873 11/29/10 01:51 a-onl End_of_Run_9068
  • 341872 11/29/10 01:49 itard entries end of run 9068
  • 341871 11/29/10 00:49 a-onl Start_Run_9068,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341870 11/29/10 00:48 a-onl End_of_Run_9067
  • 341869 11/29/10 00:02 hyde Evening shift summary 28 Nov 2010
  • 341868 11/29/10 00:00 adaq Beam Time Accounting
  • Sunday

  • 341867 11/28/10 23:43 a-onl Start_Run_9067,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341866 11/28/10 23:42 a-onl End_of_Run_9066
  • 341865 11/28/10 23:36 rakithab CODA status during the run 9066
  • 341864 11/28/10 23:36 rakithab Hall A scaler data during the run 9066
  • 341863 11/28/10 22:53 hyde Run 9065 livetime deadtime
  • 341862 11/28/10 22:43 a-onl Start_Run_9066,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341861 11/28/10 22:42 a-onl End_of_Run_9065
  • 341860 11/28/10 22:38 rakithab CODA status during the run 9065
  • 341859 11/28/10 22:38 rakithab Hall A scaler data during the run 9065
  • 341858 11/28/10 21:42 a-onl Start_Run_9065,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341857 11/28/10 21:41 a-onl End_of_Run_9064
  • 341856 11/28/10 21:38 rakithab CODA status run 9064
  • 341855 11/28/10 21:38 rakithab Hall A Scaler data during the run 9064
  • 341854 11/28/10 21:07 hyde run 9063 livetime deadtime
  • 341853 11/28/10 20:42 hyde compton laser irregular
  • 341852 11/28/10 20:42 a-onl Start_Run_9064,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341851 11/28/10 20:41 a-onl End_of_Run_9063
  • 341850 11/28/10 20:35 rakithab Hall A Scaler data run 9063
  • 341849 11/28/10 20:34 rakithab CODA status run 9063
  • 341848 11/28/10 20:11 hyde DVCS IOC rebooted
  • 341847 11/28/10 20:10 hyde calo dc levels
  • 341846 11/28/10 20:00 a-onl Start_Run_9063,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341845 11/28/10 19:50 hyde End of run entries are late
  • 341844 11/28/10 19:42 a-onl End_of_Run_9061
  • 341843 11/28/10 19:37 rakithab Hall A scaler data run 9062
  • 341842 11/28/10 19:36 rakithab CODA status run 9062
  • 341841 11/28/10 19:34 hyde gas shed alarm
  • 341840 11/28/10 19:26 auto_snap Radiation and Ozone snapshot
  • 341839 11/28/10 19:26 auto_snap Miscellaneous snapshot
  • 341838 11/28/10 19:26 auto_snap Hall A Tools snapshot
  • 341837 11/28/10 19:26 auto_snap Compton Polarimeter snapshot
  • 341836 11/28/10 19:26 auto_snap Beam Position and Current snapshot
  • 341835 11/28/10 19:14 hyde Compton rates
  • 341834 11/28/10 19:12 hyde DVCS Calo ADC
  • 341833 11/28/10 18:58 a-onl Start_Run_9062,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341832 11/28/10 18:45 hyde livetime and deadtime run 9059
  • 341831 11/28/10 18:37 rakithab Hall A scaler data run 9061
  • 341830 11/28/10 18:36 rakithab CODA status run 9061
  • 341829 11/28/10 18:30 hyde livetime and deadtime run 9059
  • 341828 11/28/10 17:57 hyde Compton rates
  • 341827 11/28/10 17:41 a-onl Start_Run_9061,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341826 11/28/10 17:39 a-onl End_of_Run_9060
  • 341825 11/28/10 17:38 rakithab Hall A Scaler data run 9060
  • 341824 11/28/10 17:37 rakithab CODA status screen shot run 9060
  • 341823 11/28/10 17:14 hyde compton rates
  • 341822 11/28/10 16:43 rakithab Hall A Scaler rates run 9059
  • 341821 11/28/10 16:39 a-onl Start_Run_9060,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341820 11/28/10 16:38 a-onl End_of_Run_9059
  • 341819 11/28/10 16:38 hyde end of run screens for 9059
  • 341818 11/28/10 16:07 adaq Beam Time Accounting
  • 341817 11/28/10 15:58 jroche Day shift summary
  • 341816 11/28/10 15:38 a-onl Start_Run_9059,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341815 11/28/10 15:37 buddhini Scaler rates before run 9058
  • 341814 11/28/10 15:37 buddhini CODA windows before end of run 9058
  • 341813 11/28/10 15:36 a-onl End_of_Run_9058
  • 341812 11/28/10 15:32 jroche The enegry feedback is now back on the Hall C beamline
  • 341811 11/28/10 15:21 jroche target screens
  • 341810 11/28/10 15:18 jroche General Tools
  • 341809 11/28/10 15:18 jroche Calo DC currents
  • 341808 11/28/10 15:09 jroche energy fastfeed back locked on hall A beam a few minutes ago
  • 341807 11/28/10 14:28 a-onl Start_Run_9058,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341806 11/28/10 14:28 buddhini Scaler rates at the end of 9058
  • 341805 11/28/10 14:26 buddhini CODA windows before end of run 9057
  • 341804 11/28/10 14:26 a-onl End_of_Run_9057
  • 341803 11/28/10 13:34 jroche Comton Hv back to 1750V during Compton run 24132
  • 341802 11/28/10 13:23 a-onl Start_Run_9057,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341801 11/28/10 12:46 buddhini Scaler rates after end of run 9056
  • 341800 11/28/10 12:46 buddhini CODA windows after end of run 9056
  • 341799 11/28/10 12:44 a-onl End_of_Run_9056
  • 341798 11/28/10 12:38 jroche lowered HV for compton to 1675V
  • 341797 11/28/10 12:05 jroche Compton rates fluctuations: CEBAF energy instable: MCC running Krest
  • 341796 11/28/10 11:56 buddhini Scaler rates at the bedining of run 9056
  • 341795 11/28/10 11:53 a-onl Start_Run_9056,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341794 11/28/10 11:52 buddhini Rebooted ROC18 and restarted CODA
  • 341793 11/28/10 11:45 buddhini Failed to end run 9055
  • 341792 11/28/10 11:41 buddhini CODA windows before end of run 9055
  • 341791 11/28/10 11:26 auto_snap Radiation and Ozone snapshot
  • 341790 11/28/10 11:26 auto_snap Miscellaneous snapshot
  • 341789 11/28/10 11:26 auto_snap Hall A Tools snapshot
  • 341788 11/28/10 11:25 auto_snap Compton Polarimeter snapshot
  • 341787 11/28/10 11:25 auto_snap Beam Position and Current snapshot
  • 341786 11/28/10 10:48 a-onl Start_Run_9055,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341785 11/28/10 10:47 buddhini Scaler rates at the end of run 9054
  • 341784 11/28/10 10:47 buddhini CODA windows before end of run 9054
  • 341783 11/28/10 10:46 a-onl End_of_Run_9054
  • 341782 11/28/10 10:12 jroche excursion of the Compton rates at the end of run 24306
  • 341781 11/28/10 09:44 a-onl Start_Run_9054,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341780 11/28/10 09:44 buddhini Scaler rates at the end of run 9054
  • 341779 11/28/10 09:43 buddhini CODA windows before end of run 9054
  • 341778 11/28/10 09:43 a-onl End_of_Run_9053
  • 341777 11/28/10 08:46 a-onl Start_Run_9053,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341776 11/28/10 08:46 buddhini Scaler rates at the end of run 5092
  • 341775 11/28/10 08:41 buddhini CODA windows at the end of run 9052
  • 341774 11/28/10 08:41 a-onl End_of_Run_9052
  • 341773 11/28/10 08:07 cshloka updated compton plots
  • 341772 11/28/10 07:59 adaq Beam Time Accounting
  • 341771 11/28/10 07:52 Carlos Owl shift summary
  • 341770 11/28/10 07:32 a-onl Start_Run_9052,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341769 11/28/10 07:32 a-onl End_of_Run_9051
  • 341768 11/28/10 07:31 cshloka scalers for run 9051
  • 341767 11/28/10 07:30 cshloka roc, coda for run 9051
  • 341766 11/28/10 07:22 Carlos DVCS IOC rebooted
  • 341765 11/28/10 06:50 a-onl Start_Run_9051,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341764 11/28/10 06:32 a-onl End_of_Run_9043
  • 341763 11/28/10 06:31 cshloka scalers for run 9050
  • 341762 11/28/10 06:30 cshloka rocs, run control screens for run 9050
  • 341761 11/28/10 05:57 Carlos BCM u/d readings
  • 341760 11/28/10 05:32 a-onl Start_Run_9050,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, Kin1 Low Energy
  • 341759 11/28/10 05:31 a-onl End_of_Run_9049
  • 341758 11/28/10 05:31 a-onl Start_Run_9049,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Efficiency Run
  • 341757 11/28/10 05:29 a-onl End_of_Run_9048
  • 341756 11/28/10 05:28 cshloka scalers for run 9048
  • 341755 11/28/10 05:28 cshloka rocs coda screens for run 9048
  • 341754 11/28/10 05:00 a-onl Start_Run_9048,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Efficiency Run
  • 341753 11/28/10 04:58 a-onl End_of_Run_9047
  • 341752 11/28/10 04:58 cshloka scalres for run 9047
  • 341751 11/28/10 04:57 cshloka roc, coda for run 9047
  • 341750 11/28/10 03:58 a-onl Start_Run_9047,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341749 11/28/10 03:57 a-onl End_of_Run_9046
  • 341748 11/28/10 03:57 cshloka scalers for run 9046
  • 341747 11/28/10 03:56 cshloka roc, coda for run 9046
  • 341746 11/28/10 03:25 auto_snap Radiation and Ozone snapshot
  • 341745 11/28/10 03:25 auto_snap Miscellaneous snapshot
  • 341744 11/28/10 03:25 auto_snap Hall A Tools snapshot
  • 341743 11/28/10 03:25 auto_snap Compton Polarimeter snapshot
  • 341742 11/28/10 03:25 auto_snap Beam Position and Current snapshot
  • 341741 11/28/10 02:55 a-onl Start_Run_9046,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341740 11/28/10 02:54 a-onl End_of_Run_9045
  • 341739 11/28/10 02:53 cshloka scalers for run 9045
  • 341738 11/28/10 02:53 cshloka roc, coda screens for run 9045
  • 341737 11/28/10 01:53 a-onl Start_Run_9045,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341736 11/28/10 01:52 a-onl End_of_Run_9044
  • 341735 11/28/10 01:52 cshloka scalers for run 9044
  • 341734 11/28/10 01:51 cshloka roc, coda screens for run 9044
  • 341733 11/28/10 01:08 a-onl Start_Run_9044,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341732 11/28/10 01:05 Carlos DVCS IOC rebooted
  • 341731 11/28/10 00:51 cshloka scalers for run 9043
  • 341730 11/28/10 00:49 cshloka roc and coda screens for run 9043
  • 341729 11/28/10 00:29 adaq Beam Time Accounting
  • 341728 11/28/10 00:16 jroche Swing shift summary
  • 341727 11/28/10 00:10 jroche First results from the BCM calibration of November 24 (EPICS VALUES)
  • Saturday

  • 341726 11/27/10 23:51 jroche new BCM downstream calibration factor for EPICS
  • 341725 11/27/10 23:48 a-onl Start_Run_9043,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341724 11/27/10 23:47 a-onl End_of_Run_9042
  • 341723 11/27/10 23:44 a-onl Start_Run_9042,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341722 11/27/10 23:43 a-onl End_of_Run_9041
  • 341721 11/27/10 23:39 punjabi roc, coda screens, run 9041
  • 341720 11/27/10 22:42 a-onl Start_Run_9041,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341719 11/27/10 22:41 a-onl End_of_Run_9040
  • 341718 11/27/10 22:41 punjabi coda screen, run 9040
  • 341717 11/27/10 22:40 punjabi roc screens, run 9040
  • 341716 11/27/10 21:49 jroche target screens
  • 341715 11/27/10 21:44 jroche General Tools (shift check list)
  • 341714 11/27/10 21:41 a-onl Start_Run_9040,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341713 11/27/10 21:41 a-onl End_of_Run_9039
  • 341712 11/27/10 21:40 jroche Calo DC current
  • 341711 11/27/10 21:39 punjabi coda screen, run 9039
  • 341710 11/27/10 21:39 punjabi roc screens, run 9039
  • 341709 11/27/10 20:42 a-onl Start_Run_9039,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341708 11/27/10 20:41 a-onl End_of_Run_9038
  • 341707 11/27/10 20:40 punjabi coda screen, run 9038
  • 341706 11/27/10 20:39 punjabi roc screens, run 9038
  • 341705 11/27/10 19:42 a-onl Start_Run_9038,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341704 11/27/10 19:41 a-onl End_of_Run_9037
  • 341703 11/27/10 19:38 punjabi coda screen
  • 341702 11/27/10 19:35 punjabi roc screens, run 9037
  • 341701 11/27/10 19:25 auto_snap Radiation and Ozone snapshot
  • 341700 11/27/10 19:25 auto_snap Miscellaneous snapshot
  • 341699 11/27/10 19:25 auto_snap Hall A Tools snapshot
  • 341698 11/27/10 19:25 auto_snap Compton Polarimeter snapshot
  • 341697 11/27/10 19:25 auto_snap Beam Position and Current snapshot
  • 341696 11/27/10 18:41 a-onl Start_Run_9037,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341695 11/27/10 18:38 punjabi coda screen
  • 341694 11/27/10 18:37 punjabi roc screens
  • 341693 11/27/10 17:53 a-onl Start_Run_9036,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341692 11/27/10 17:52 a-onl End_of_Run_9035
  • 341691 11/27/10 17:45 Punjabi cada screen
  • 341690 11/27/10 17:44 Punjabi Roc screens
  • 341689 11/27/10 17:31 jroche Comptom Photon detector tabl toy=43mm
  • 341688 11/27/10 16:51 a-onl Start_Run_9035,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341687 11/27/10 16:49 jroche harp scan results
  • 341686 11/27/10 16:47 jroche target moved to LH2 15cm
  • 341685 11/27/10 16:42 Ramesh Half wave plate IN
  • 341684 11/27/10 16:07 efuchey 11-27-day shift summary
  • 341683 11/27/10 15:59 adaq Beam Time Accounting
  • 341682 11/27/10 15:55 efuchey Beam recovery procedure
  • 341681 11/27/10 15:41 efuchey MCC sending tune beam
  • 341680 11/27/10 15:35 efuchey Target moved to empty
  • 341679 11/27/10 15:32 efuchey Bema update
  • 341678 11/27/10 15:24 efuchey ABU, BANU for MCC
  • 341677 11/27/10 13:58 efuchey Beam down (again)
  • 341676 11/27/10 13:44 efuchey Calo base currents
  • 341675 11/27/10 13:36 efuchey Compton rates after BCR
  • 341674 11/27/10 13:30 a-onl Start_Run_9034,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin1-low
  • 341673 11/27/10 13:28 efuchey target move to LH2
  • 341672 11/27/10 13:25 efuchey Beam is back (again !)
  • 341671 11/27/10 13:07 efuchey Hall A screenshots
  • 341670 11/27/10 13:04 efuchey Target charts
  • 341669 11/27/10 13:03 efuchey Taget screenshots
  • 341668 11/27/10 11:40 efuchey target moved to empty.
  • 341667 11/27/10 11:24 efuchey beam down again for about one hour.
  • 341666 11/27/10 11:19 efuchey target moved to LH2 15 cm
  • 341665 11/27/10 11:15 efuchey Beam is back
  • 341664 11/27/10 10:57 efuchey Target moved to EMPTY
  • 341663 11/27/10 10:53 efuchey Beam coming back soon.
  • 341662 11/27/10 10:23 efuchey QWP scan for 11-27-2010
  • 341661 11/27/10 10:18 efuchey target charts
  • 341660 11/27/10 10:14 efuchey beam update
  • 341659 11/27/10 10:05 a-onl End_of_Run_9033
  • 341658 11/27/10 09:53 a-onl Start_Run_9033,Run_type=Pedestal,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS pedestal run
  • 341657 11/27/10 09:52 efuchey Update on beam status
  • 341656 11/27/10 09:34 efuchey Beam down
  • 341655 11/27/10 09:33 a-onl End_of_Run_9032
  • 341654 11/27/10 08:42 Carlos Run 9031: ROC17 not responding
  • 341653 11/27/10 08:41 Carlos Run 9030: TS autorebooted
  • 341652 11/27/10 08:40 a-onl Start_Run_9032,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341651 11/27/10 08:35 a-onl Start_Run_9031,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341650 11/27/10 08:31 a-onl Start_Run_9030,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341649 11/27/10 08:31 a-onl End_of_Run_9029
  • 341648 11/27/10 08:30 Carlos DAQ screens run 9029
  • 0 11/27/10 08:08 no_name no_key
  • 341646 11/27/10 08:01 adaq Beam Time Accounting
  • 341645 11/27/10 07:56 jroche Owl shift summary
  • 341644 11/27/10 07:32 jroche How I tame Q1 (with two striptools)
  • 341643 11/27/10 07:30 a-onl Start_Run_9029,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341642 11/27/10 07:29 a-onl End_of_Run_9028
  • 341641 11/27/10 07:29 itard entries end of run 9028
  • 341640 11/27/10 06:48 jroche RF separator is loosing tune
  • 341639 11/27/10 06:42 a-onl Start_Run_9028,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341638 11/27/10 06:28 itard entries end of run 9026
  • 341637 11/27/10 05:40 jroche BCM calibration on November 24 (BCM log file)
  • 341636 11/27/10 05:39 a-onl Start_Run_9027,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341635 11/27/10 05:32 a-onl End_of_Run_9026
  • 341634 11/27/10 05:32 a-onl End_of_Run_9021
  • 341633 11/27/10 05:32 itard entries end of run 9026
  • 341632 11/27/10 04:36 jroche calo dc current
  • 341631 11/27/10 04:30 a-onl Start_Run_9026,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341630 11/27/10 04:29 a-onl End_of_Run_9025
  • 341629 11/27/10 04:29 itard entries end of run 9025
  • 341628 11/27/10 03:30 a-onl Start_Run_9025,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341627 11/27/10 03:30 a-onl End_of_Run_9024
  • 341626 11/27/10 03:29 itard entries end of run 9024
  • 341625 11/27/10 02:29 a-onl Start_Run_9024,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341624 11/27/10 02:19 jroche loosing Q1 again
  • 341623 11/27/10 02:15 a-onl End_of_Run_9023
  • 341622 11/27/10 02:14 itard entries end of run
  • 341621 11/27/10 01:56 jroche rebooted DVCS ioc
  • 341620 11/27/10 01:34 a-onl Start_Run_9023,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341619 11/27/10 01:12 jroche Q1 power supply screen
  • 341618 11/27/10 01:12 jroche Left D and Q1 recovered
  • 341617 11/27/10 00:41 itard cryo and magnet problem
  • 341616 11/27/10 00:40 jroche sorry, Q1 and the dipole (not Q3) tripped
  • 341615 11/27/10 00:29 adaq loosing Q1 and then Q3
  • 341614 11/27/10 00:27 a-onl End_of_Run_9022
  • 341613 11/27/10 00:27 itard entries end of run
  • 341612 11/27/10 00:01 a-onl Start_Run_9022,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341611 11/27/10 00:01 a-onl End_of_Run_9016
  • 341609 11/27/10 00:00 Carlos DAQ screens run 9021
  • Friday

  • 341610 11/26/10 23:59 adaq Beam Time Accounting
  • 341608 11/26/10 23:47 efuchey target charts screenshot
  • 341607 11/26/10 23:01 a-onl Start_Run_9021,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341606 11/26/10 23:01 a-onl End_of_Run_9020
  • 341605 11/26/10 22:59 Carlos DAQ screens run 9020
  • 341604 11/26/10 22:07 efuchey target charts screenshot
  • 341603 11/26/10 22:01 a-onl Start_Run_9020,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341602 11/26/10 22:00 a-onl End_of_Run_9019
  • 341601 11/26/10 21:59 Carlos DAQ screens run 9019
  • 341600 11/26/10 21:12 efuchey erratum on compton logbook
  • 341599 11/26/10 21:00 a-onl Start_Run_9019,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341598 11/26/10 20:59 a-onl End_of_Run_9018
  • 341597 11/26/10 20:59 Carlos DAQ screens run 9018
  • 341596 11/26/10 19:59 a-onl Start_Run_9018,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341595 11/26/10 19:59 a-onl End_of_Run_9017
  • 341594 11/26/10 19:59 Carlos DAQ screens run 9017
  • 341593 11/26/10 19:56 efuchey Target chart screenshots
  • 341592 11/26/10 18:59 a-onl Start_Run_9017,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341591 11/26/10 18:40 efuchey Y position in the Compton chicane
  • 341590 11/26/10 18:03 efuchey Target charts
  • 341589 11/26/10 17:58 a-onl Start_Run_9016,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341588 11/26/10 17:58 a-onl End_of_Run_9015
  • 341587 11/26/10 16:58 a-onl Start_Run_9015,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production: Kin1-low
  • 341586 11/26/10 16:56 a-onl End_of_Run_9014
  • 341585 11/26/10 16:36 efuchey Calo base currents
  • 341584 11/26/10 16:34 efuchey Hall A screeshots
  • 341583 11/26/10 16:32 efuchey efficiency run: 9014
  • 341582 11/26/10 16:30 a-onl Start_Run_9014,Run_type=Other,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger efficiency run
  • 341581 11/26/10 16:29 efuchey target charts screenshots
  • 341580 11/26/10 16:27 efuchey Target screenshots
  • 341579 11/26/10 16:18 efuchey trigger supervisor crash
  • 341578 11/26/10 16:01 adaq Beam Time Accounting
  • 341577 11/26/10 15:59 itard Day shift summary
  • 341576 11/26/10 15:38 a-onl Start_Run_9013,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341575 11/26/10 15:34 a-onl End_of_Run_9012
  • 341574 11/26/10 15:33 buddhini coda windows before end of run 9012
  • 341573 11/26/10 14:31 a-onl Start_Run_9012,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341572 11/26/10 14:31 buddhini coda windows before the end of run 9011
  • 341571 11/26/10 14:30 a-onl End_of_Run_9011
  • 341570 11/26/10 13:33 A. Deur closed again JT valve by 2%
  • 341569 11/26/10 13:27 buddhini coda windows before end of run 2010
  • 341568 11/26/10 13:24 a-onl Start_Run_9011,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341567 11/26/10 13:23 a-onl End_of_Run_9010
  • 341566 11/26/10 13:19 A. Deur target snapshot
  • 341565 11/26/10 13:18 A. Deur Closed JT valve by 2%
  • 341564 11/26/10 12:21 adaq coda screens at the end of the run.
  • 341563 11/26/10 12:21 a-onl Start_Run_9010,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341562 11/26/10 12:19 a-onl End_of_Run_9009
  • 341561 11/26/10 11:18 A. Deur target snapshot
  • 341560 11/26/10 11:15 a-onl Start_Run_9009,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341559 11/26/10 11:14 a-onl End_of_Run_9008
  • 341558 11/26/10 11:12 A. Camsonne Possible double pulsing
  • 341557 11/26/10 10:56 itard parameters hall A
  • 341556 11/26/10 10:15 buddhini Comments for run 9007
  • 341555 11/26/10 10:15 a-onl Start_Run_9008,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kinematic 1 low energy.
  • 341554 11/26/10 10:12 A. Camsonne Hall A Compton lock turned on
  • 341553 11/26/10 10:11 a-onl End_of_Run_9007
  • 341552 11/26/10 10:02 A. Deur target action for beam recovery
  • 341551 11/26/10 09:59 a-onl Start_Run_9007,Run_type=Production,target_type=LH2,beam_e0=3.3,beam_raster=off,beam_current=2,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=At the begining of the run MCC was getting ready to send 2uA beam.
  • 341550 11/26/10 09:59 A. Deur cernox 1R is bad
  • 341549 11/26/10 09:56 a-onl Start_Run_9005,Run_type=Cosmics,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ restored previous trigger software
  • 341548 11/26/10 09:46 A. Camsonne Compton background ok - Lowered HV to 1750 V
  • 341547 11/26/10 09:43 A. Deur target snapshot
  • 341546 11/26/10 09:34 a-onl End_of_Run_9006
  • 341545 11/26/10 09:28 a-onl Start_Run_9006,Run_type=Cosmics,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ restored previous trigger software
  • 341544 11/26/10 09:22 buddhini Actually target was empty target.
  • 341543 11/26/10 09:15 a-onl End_of_Run_9005
  • 341542 11/26/10 08:56 adaq Run 9004 : cosmic run, Tantalum target, No beam
  • 341541 11/26/10 08:01 A. Camsonne Restored running conditions for DVCS electronics
  • Thursday

  • 341540 11/25/10 22:41 a-onl End_of_Run_9002
  • 341539 11/25/10 22:40 a-onl Start_Run_9002,Run_type=Cosmics,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ after adaqfs rebuild
  • 341538 11/25/10 22:22 a-onl End_of_Run_9001
  • 341537 11/25/10 22:21 A. Camsonne LeftHRS CODA run successfully started
  • 341536 11/25/10 22:19 a-onl Start_Run_9001,Run_type=Cosmics,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ after adaqfs rebuild
  • 341535 11/25/10 21:36 Carlos Target screenshot
  • 341534 11/25/10 18:54 Ole Hansen adaq and a-onl home dirs fairly up to date
  • 341533 11/25/10 18:41 Ole Hansen Computers rebooted, up and running
  • 341532 11/25/10 18:38 R. Michaels restored diskhost
  • 341531 11/25/10 17:46 A. Camsonne adaqfs partially restored - DAQ almost running - error generating run file name
  • 341530 11/25/10 17:37 a-onl End_of_Run_8999
  • 341529 11/25/10 17:34 a-onl Start_Run_8999,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ after adaqfs rebuild
  • 341528 11/25/10 17:32 a-onl End_of_Run_8998
  • 341527 11/25/10 10:55 O.Glamazdin Moller target encoder
  • 341526 11/25/10 08:10 R. Michaels adaqfs is being rebuilt -- will be ready today
  • 341525 11/25/10 07:57 Carlos Owl shift summary
  • 341524 11/25/10 04:56 A. Camsonne Power cycled adaqfs
  • 341523 11/25/10 02:39 A. Camsonne adaqfs computer in bad shape
  • 341521 11/25/10 00:00 S. Golge hall a target snapshot swing
  • 341522 11/25/10 00:00 adaq Beam Time Accounting
  • Wednesday

  • 341520 11/24/10 19:33 S. Golge fan speed temperature normal
  • 341519 11/24/10 16:58 S. Golge fan readout problem
  • 341518 11/24/10 16:48 A. Camsonne NFS issues on adaq computers
  • 341517 11/24/10 16:29 A. Camsonne DVCS calorimeter cosmics paddles HV on
  • 341516 11/24/10 16:26 A. Camsonne Restarting HV GUI for calorimeter
  • 341514 11/24/10 16:08 A. Deur Day shift summary
  • 341513 11/24/10 16:07 A. Deur 15K temperature keeps increasing and cryoflow is decreasing
  • 341515 11/24/10 16:07 adaq Beam Time Accounting
  • 341512 11/24/10 15:04 A. Deur Target L1 temperature increase. JT valves adjusted
  • 341511 11/24/10 14:49 Kalyan Moller will remain warm
  • 341510 11/24/10 14:25 fansler RQ2 N2
  • 341509 11/24/10 13:46 a-onl End_of_Run_8996
  • 341508 11/24/10 13:06 a-onl Start_Run_8996,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCS
  • 341507 11/24/10 12:56 hyde Deadtime and Livetime
  • 341506 11/24/10 12:38 a-onl End_of_Run_8995
  • 341505 11/24/10 12:38 a-onl End_of_Run_8993
  • 341504 11/24/10 12:32 a-onl Start_Run_8995,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pockells cell at 960 Hz
  • 341503 11/24/10 12:30 a-onl End_of_Run_8994
  • 341502 11/24/10 12:24 a-onl Start_Run_8994,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Helicity flip at 900 Hz
  • 341501 11/24/10 12:20 a-onl End_of_Run_8992
  • 341500 11/24/10 12:19 a-onl Start_Run_8993,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Poeckle cell off
  • 341499 11/24/10 12:07 a-onl Start_Run_8992,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Poeckle cell on and octoquads with 64 windows delay.
  • 341498 11/24/10 12:06 a-onl End_of_Run_8991
  • 341497 11/24/10 11:45 a-onl End_of_Run_8989
  • 341496 11/24/10 11:26 a-onl Start_Run_8991,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Poeckle cell off at the start of this run.
  • 341495 11/24/10 11:13 auto_snap Compton Polarimeter snapshot
  • 341494 11/24/10 11:13 auto_snap Beam Position and Current snapshot
  • 341493 11/24/10 11:08 a-onl End_of_Run_8990
  • 341492 11/24/10 10:47 a-onl Start_Run_8990,Run_type=Calibration,target_type=No Target, Empty,beam_e0=3.355,beam_raster=off,beam_current=---,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=BCM callibration run with different trigger file and PS factors
  • 341491 11/24/10 10:28 a-onl Start_Run_8989,Run_type=Calibration,target_type=No Target, Empty,beam_e0=3.355,beam_raster=off,beam_current=---,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=BCM callibration run with different trigger file and PS factors
  • 341490 11/24/10 10:28 a-onl End_of_Run_8987
  • 341489 11/24/10 10:19 auto_snap Compton Polarimeter snapshot
  • 341488 11/24/10 10:19 auto_snap Beam Position and Current snapshot
  • 341487 11/24/10 09:20 a-onl End_of_Run_8988
  • 341486 11/24/10 08:52 a-onl Start_Run_8988,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run Helicity reversal 900 Hz
  • 341485 11/24/10 08:34 a-onl End_of_Run_8986
  • 341484 11/24/10 08:27 a-onl Start_Run_8987,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341483 11/24/10 08:09 a-onl End_of_Run_8982
  • 341482 11/24/10 08:06 Lamiaa Coda & Rocs snapshot for a run 8986
  • 341481 11/24/10 08:05 Lamiaa Scalers snapshot for a run 8986
  • 341480 11/24/10 08:03 miham Owl Shift Summary (11/24/10)
  • 341479 11/24/10 08:00 adaq Beam Time Accounting
  • 341478 11/24/10 07:26 a-onl Start_Run_8986,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341477 11/24/10 07:23 a-onl End_of_Run_8985
  • 341476 11/24/10 07:06 Lamiaa Scalers snapshot for a run 8985
  • 341475 11/24/10 07:06 Lamiaa Coda & Rocs snapshot for a run 8985
  • 341474 11/24/10 07:03 Lamiaa Target's screens
  • 341473 11/24/10 06:49 miham Another beam current change
  • 341472 11/24/10 06:10 a-onl Start_Run_8985,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341471 11/24/10 06:09 a-onl End_of_Run_8984
  • 341470 11/24/10 06:07 Lamiaa Scalers snapshot for a run 8984
  • 341469 11/24/10 06:06 Lamiaa Coda & Rocs Snapshot for a run 8984
  • 341468 11/24/10 05:08 a-onl Start_Run_8984,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341467 11/24/10 05:08 a-onl End_of_Run_8983
  • 341466 11/24/10 05:07 Lamiaa Scalers snapshot for a run 8983
  • 341465 11/24/10 05:07 Lamiaa Coda & Rocs Snapshot for a run 8983
  • 341464 11/24/10 04:32 miham Compton rate as function of HallC beam current
  • 341463 11/24/10 04:07 a-onl Start_Run_8983,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341462 11/24/10 04:05 Lamiaa Coda & Rocs Snapshot for a run 8982
  • 341461 11/24/10 04:04 Lamiaa Scalers snapshot for a run 8982
  • 341460 11/24/10 03:13 auto_snap Compton Polarimeter snapshot
  • 341459 11/24/10 03:13 auto_snap Beam Position and Current snapshot
  • 341458 11/24/10 03:12 miham Compton signal correlation to HallC
  • 341457 11/24/10 03:08 Lamiaa Scalers snapshot for a run 8981
  • 341456 11/24/10 03:08 Lamiaa Coda & Rocs Snapshot for a run 8981
  • 341455 11/24/10 03:07 a-onl Start_Run_8982,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341454 11/24/10 03:07 a-onl End_of_Run_8981
  • 341453 11/24/10 02:44 miham Beam Problems
  • 341452 11/24/10 02:19 auto_snap Compton Polarimeter snapshot
  • 341451 11/24/10 02:19 auto_snap Beam Position and Current snapshot
  • 341450 11/24/10 02:05 a-onl Start_Run_8981,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341449 11/24/10 02:04 a-onl End_of_Run_8980
  • 341448 11/24/10 02:02 LamiaaLamiaa Coda & Rocs for a run 8980
  • 341447 11/24/10 02:01 Lamiaa Scalers snapshot for a run 8980
  • 341446 11/24/10 01:39 miham Beam current slightly lower than 3.0uA
  • 341445 11/24/10 01:38 Lamiaa Target's gui and charts.
  • 341444 11/24/10 01:15 miham Compton not working properly?
  • 341443 11/24/10 01:08 miham DVCS HV Gui Screenshot - All Green
  • 341442 11/24/10 01:07 a-onl Start_Run_8980,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341441 11/24/10 01:07 miham DVCS Calorimeter ADC readback is Back
  • 341440 11/24/10 01:05 a-onl End_of_Run_8978
  • 341439 11/24/10 01:04 a-onl End_of_Run_8979
  • 341438 11/24/10 00:53 Lamiaa Scalers for a run 8979
  • 341437 11/24/10 00:52 Lamiaa CODA & Rocs for a run 8979
  • 341436 11/24/10 00:36 miham DVCS Calorimeter ADC readback is frozen - need to reboot
  • 341435 11/24/10 00:33 miham HallA tools at the beginning of owl shift
  • 341434 11/24/10 00:25 Sirca Swing Shift Summary (11/23/10)
  • 341433 11/24/10 00:23 adaq Beam Time Accounting
  • 341432 11/24/10 00:13 a-onl Start_Run_8979,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341431 11/24/10 00:01 a-onl End_of_Run_8976
  • 341430 11/24/10 00:00 S. Sirca Coda reboot for Compton
  • Tuesday

  • 341429 11/23/10 23:54 A. Tobias RunControl and ROCs for run 8978
  • 341428 11/23/10 23:52 A. Tobias Scalars for run 8978
  • 341427 11/23/10 23:39 S. Sirca Better zoom of Compton signal
  • 341426 11/23/10 23:31 S. Sirca Excessive G-rates in Compton
  • 341425 11/23/10 23:19 S. Sirca BCM temperature is higher than usual
  • 341424 11/23/10 23:00 A. Tobias Target: Loop2 LH2 settings
  • 341423 11/23/10 22:52 a-onl Start_Run_8978,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341422 11/23/10 22:51 A. Tobias Moved target to Loop2 LH2
  • 341421 11/23/10 22:47 S. Sirca Harp results acceptable now
  • 341420 11/23/10 22:20 adaq Beam Time Accounting
  • 341419 11/23/10 21:24 Jack (Techs) Gas Shed flammable gas sensor
  • 341418 11/23/10 21:18 S. Sirca Harp scans show too large sigma_y
  • 341417 11/23/10 21:17 S. Sirca Restarted GUI
  • 341416 11/23/10 20:52 S. Sirca Switch cal HV on: no V monitor on Slot 0?
  • 341415 11/23/10 20:37 O.Glamazdin Moller measurement summary
  • 341414 11/23/10 19:41 A. Tobias Target: Loop 1 & 2 JT Valves adjusted closed a liitle in the past 3.5 hrs
  • 341413 11/23/10 19:13 auto_snap Compton Polarimeter snapshot
  • 341412 11/23/10 19:13 auto_snap Beam Position and Current snapshot
  • 341411 11/23/10 18:18 auto_snap Compton Polarimeter snapshot
  • 341410 11/23/10 18:18 auto_snap Beam Position and Current snapshot
  • 341409 11/23/10 16:35 efuchey rates are too big.
  • 341408 11/23/10 16:05 S. Sirca Methane alarm
  • 341407 11/23/10 15:40 efuchey Compton rates for spin dance
  • 341406 11/23/10 13:35 J. Butler(te Low Pressure Gas Alarm
  • 341405 11/23/10 13:18 jroche ratio u3/d3 (update)
  • 341404 11/23/10 11:13 auto_snap Compton Polarimeter snapshot
  • 341403 11/23/10 11:13 auto_snap Beam Position and Current snapshot
  • 341402 11/23/10 10:49 Carlos Calorimeter HV off
  • 341401 11/23/10 10:48 Carlos Compton HV off
  • 341400 11/23/10 10:47 Carlos Day shift summary (before spin dance)
  • 341399 11/23/10 10:45 adaq Target temperature increase
  • 341398 11/23/10 10:30 adaq Problems in run 8977
  • 341397 11/23/10 10:18 auto_snap Compton Polarimeter snapshot
  • 341396 11/23/10 10:18 auto_snap Beam Position and Current snapshot
  • 341395 11/23/10 09:44 Carlos Beam down
  • 341394 11/23/10 09:42 a-onl End_of_Run_8977
  • 341393 11/23/10 09:31 Carlos Calorimeter ADCs
  • 341392 11/23/10 09:27 Carlos DVCS IOC reboot
  • 341391 11/23/10 09:23 a-onl Start_Run_8977,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341390 11/23/10 09:18 Carlos Rates RUN 8976
  • 341389 11/23/10 09:18 Carlos Coda screen RUN 8976
  • 341388 11/23/10 09:10 Carlos BCM temperature
  • 341387 11/23/10 08:16 a-onl Start_Run_8976,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341386 11/23/10 08:15 a-onl End_of_Run_8975
  • 341385 11/23/10 08:11 Lamiaa Coda & Rocs Snapshot for a run 8975
  • 341384 11/23/10 08:10 Lamiaa Scalers snapshot for a run 8975
  • 341383 11/23/10 08:10 miham Owl Shift Summary (11/23/10)
  • 341382 11/23/10 08:03 adaq Beam Time Accounting
  • 341381 11/23/10 07:15 miham We are back in production!
  • 341380 11/23/10 07:14 a-onl Start_Run_8975,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341379 11/23/10 07:13 Lamiaa Target gui & charts @ LH2 position
  • 341378 11/23/10 07:13 miham BPMs moved after Harp scan
  • 341377 11/23/10 07:11 miham Harp Scan results
  • 341376 11/23/10 06:32 miham BPMs seem to be OK!
  • 341375 11/23/10 06:31 miham Beam is back !
  • 341374 11/23/10 06:18 miham Still waiting for beam to come back!
  • 341373 11/23/10 05:32 Lamiaa Move target to Empty for beam recovery
  • 341372 11/23/10 04:30 a-onl End_of_Run_8974
  • 341371 11/23/10 04:30 miham No beam!
  • 341370 11/23/10 04:29 Lamiaa Scalers snapshot for a run 8974
  • 341369 11/23/10 04:29 Lamiaa CODA & Rocs Snapshot
  • 341368 11/23/10 04:11 miham Another jump in beam current
  • 341367 11/23/10 03:35 a-onl Start_Run_8974,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341366 11/23/10 03:34 a-onl End_of_Run_8973
  • 341365 11/23/10 03:33 Lamiaa Scalers snapshot for a run 8973
  • 341364 11/23/10 03:32 Lamiaa Snapshot od DOCA & Rocs of a run 8973
  • 341363 11/23/10 03:23 miham Sudden Current change
  • 341362 11/23/10 03:13 auto_snap Compton Polarimeter snapshot
  • 341361 11/23/10 03:13 auto_snap Beam Position and Current snapshot
  • 341360 11/23/10 02:35 a-onl Start_Run_8973,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341359 11/23/10 02:34 a-onl End_of_Run_8972
  • 341358 11/23/10 02:33 Lamiaa Scalers snapshot of a run 8972
  • 341357 11/23/10 02:33 Lamiaa Snapshot od DOCA & Rocs of a run 8972
  • 341356 11/23/10 02:18 auto_snap Compton Polarimeter snapshot
  • 341355 11/23/10 02:18 auto_snap Beam Position and Current snapshot
  • 341354 11/23/10 01:50 miham ROC 17 crashed during run #8971
  • 341353 11/23/10 01:45 miham Compton run #24203 End-of-run script not filled properly
  • 341352 11/23/10 01:39 miham BPMs are better now
  • 341351 11/23/10 01:39 Lamiaa Snapshot of CODA & Rocs of a run 8971
  • 341350 11/23/10 01:35 a-onl Start_Run_8972,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341349 11/23/10 00:56 Lamiaa Target gui and charts snapshot
  • 341348 11/23/10 00:46 a-onl Start_Run_8971,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LH2 Production run
  • 341347 11/23/10 00:45 a-onl End_of_Run_8970
  • 341346 11/23/10 00:44 Lamiaa Scalers snapshot for a run 8970
  • 341345 11/23/10 00:41 Lamiaa Coda & Rocs Snapshot for a run 8970
  • 341344 11/23/10 00:41 miham Calorimeter ADC Readback monitor screenshot
  • 341343 11/23/10 00:38 miham Screenshot of Calorimeter HV monitor - All Green
  • 341342 11/23/10 00:37 miham HallA tools at the beginning of OWL shift
  • 341341 11/23/10 00:20 S. Sirca Swing shift SUMMARY
  • 341340 11/23/10 00:17 adaq Beam Time Accounting
  • Monday

  • 341339 11/22/10 23:44 a-onl Start_Run_8970,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341338 11/22/10 23:43 a-onl End_of_Run_8969
  • 341337 11/22/10 23:42 A. Tobias RunControl & ROCs for run 8969
  • 341336 11/22/10 23:41 A. Tobias Scalars for run 8969
  • 341335 11/22/10 23:10 S. Sirca Rebooted DVCS calo IOC
  • 341334 11/22/10 23:10 A. Tobias Target snapshot
  • 341333 11/22/10 23:01 a-onl Start_Run_8969,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341332 11/22/10 22:41 A. Tobias RunControl & ROCs for run 8968
  • 341331 11/22/10 22:40 A. Tobias Scalars for run 8968
  • 341330 11/22/10 22:39 A. Tobias Compton Snapshot
  • 341329 11/22/10 21:45 a-onl Start_Run_8968,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341328 11/22/10 21:44 a-onl End_of_Run_8967
  • 341327 11/22/10 21:40 A. Tobias RunControl & ROCs for run 8967
  • 341326 11/22/10 21:35 A. Tobias Scalars for run 8967
  • 341325 11/22/10 21:00 A. Tobias Target Snapshot
  • 341324 11/22/10 20:44 a-onl Start_Run_8967,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341323 11/22/10 20:43 a-onl End_of_Run_8966
  • 341322 11/22/10 20:40 A. Tobias Target: Loop 1 Cernox Sensor Alarm Turned OFF
  • 341321 11/22/10 20:38 S. Sirca Beam current fluctuations
  • 341320 11/22/10 20:20 A. Tobias Compton Rates Changed
  • 341319 11/22/10 19:43 A. Tobias Target: Loop 2 LH2 - JT Valve Settings
  • 341318 11/22/10 19:40 Carlos Golden run updated
  • 341317 11/22/10 19:40 a-onl Start_Run_8966,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341316 11/22/10 19:39 a-onl End_of_Run_8965
  • 341315 11/22/10 19:36 a-onl Start_Run_8965,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341314 11/22/10 19:35 a-onl End_of_Run_8964
  • 341313 11/22/10 19:12 auto_snap Compton Polarimeter snapshot
  • 341312 11/22/10 19:12 auto_snap Beam Position and Current snapshot
  • 341311 11/22/10 18:34 a-onl Start_Run_8964,Run_type=Production,target_type=LH2,beam_e0=3.355,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341310 11/22/10 18:24 A. Camsonne Compton signal over background about 9
  • 341309 11/22/10 18:21 A. Tobias Target: Loop 1 Cernox Temp Sensor gone BAD?
  • 341308 11/22/10 18:18 auto_snap Compton Polarimeter snapshot
  • 341307 11/22/10 18:18 auto_snap Beam Position and Current snapshot
  • 341306 11/22/10 16:39 A. Tobias Adjusted Loop Fans and JT Valves
  • 341305 11/22/10 16:33 A. Tobias Target moved to EMPTY
  • 341304 11/22/10 16:06 adaq Beam Time Accounting
  • 341303 11/22/10 16:05 Lindgren Shift sumary
  • 341302 11/22/10 16:04 Lindgren Calo DC levels
  • 341301 11/22/10 16:04 lindgren Calo HV
  • 341300 11/22/10 15:22 Lindgren Problem with electron gun
  • 341299 11/22/10 14:29 Lamiaa Move target to LH2.
  • 341298 11/22/10 14:00 LIndgren Harp scan
  • 341297 11/22/10 13:45 A. Camsonne Compton background rates look reasonable for now
  • 341296 11/22/10 13:17 a-onl End_of_Run_8963
  • 341295 11/22/10 13:13 a-onl Start_Run_8963,Run_type=Other,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.136,leftarm_theta=36.56,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341294 11/22/10 11:16 Lindgren Hall A Tools
  • 341293 11/22/10 11:14 Lindgren Beam Permit
  • 341292 11/22/10 11:12 auto_snap Compton Polarimeter snapshot
  • 341291 11/22/10 11:12 auto_snap Beam Position and Current snapshot
  • 341290 11/22/10 10:24 jroche, muno LHRS to 36.56 deg
  • 341289 11/22/10 10:18 auto_snap Compton Polarimeter snapshot
  • 341288 11/22/10 10:18 auto_snap Beam Position and Current snapshot
  • 341287 11/22/10 09:35 O.Glamazdin Moller magnet has been filled
  • 341286 11/22/10 09:28 Lamiaa Target screens shots
  • 341285 11/22/10 08:50 Lindgren Hall A access
  • 341284 11/22/10 08:34 a-onl End_of_Run_8962
  • 341283 11/22/10 08:22 Lindgren Pass change
  • 341282 11/22/10 08:19 Lindgren day shift starting
  • 341281 11/22/10 08:04 a-onl Start_Run_8962,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341279 11/22/10 08:00 a-onl End_of_Run_8961
  • 341280 11/22/10 08:00 adaq Beam Time Accounting
  • 341278 11/22/10 07:59 Markowitz scalers run 8961
  • 341277 11/22/10 07:59 Markowitz rocs run 8961
  • 341276 11/22/10 07:54 A. Puckett shift summary
  • 341275 11/22/10 06:56 a-onl Start_Run_8961,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341274 11/22/10 06:56 a-onl End_of_Run_8960
  • 341273 11/22/10 06:55 Markowitz scalers run8960
  • 341272 11/22/10 06:54 Markowitz roc run 8960
  • 341271 11/22/10 05:48 a-onl Start_Run_8960,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341270 11/22/10 05:47 a-onl End_of_Run_8959
  • 341269 11/22/10 05:31 Markowitz target screen shots
  • 341268 11/22/10 04:40 a-onl Start_Run_8959,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341267 11/22/10 04:39 a-onl End_of_Run_8958
  • 341266 11/22/10 04:39 Markowitz run 8958 scalers
  • 341265 11/22/10 04:38 Markowitz Run 8958 daq windows
  • 341264 11/22/10 03:55 puckett correction: we are experiencing deadtime comparable to previous shifts on LD2
  • 341263 11/22/10 03:41 puckett High deadtime
  • 341262 11/22/10 03:35 a-onl Start_Run_8958,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341261 11/22/10 03:34 a-onl End_of_Run_8957
  • 341260 11/22/10 03:12 auto_snap Compton Polarimeter snapshot
  • 341259 11/22/10 03:12 auto_snap Beam Position and Current snapshot
  • 341258 11/22/10 02:37 puckett ended CODA run 8956 and did startcoda
  • 341257 11/22/10 02:31 a-onl Start_Run_8957,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341256 11/22/10 02:18 auto_snap Compton Polarimeter snapshot
  • 341255 11/22/10 02:18 auto_snap Beam Position and Current snapshot
  • 341254 11/22/10 02:13 a-onl End_of_Run_8955
  • 341253 11/22/10 02:12 a-onl Start_Run_8956,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341252 11/22/10 02:10 Markowitz scaler screens 8955
  • 341251 11/22/10 02:10 Markowitz screens for run 8955
  • 341250 11/22/10 01:55 puckett rebooted DVCS calo ADC crate
  • 341249 11/22/10 01:05 a-onl Start_Run_8955,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341248 11/22/10 01:01 Markowitz moved to LD2
  • 341247 11/22/10 00:55 puckett Harp scan results
  • 341246 11/22/10 00:01 miham Swing Shift Summary (11/22/10)
  • Sunday

  • 341245 11/21/10 23:59 adaq Beam Time Accounting
  • 341244 11/21/10 22:46 LeRose Target Screens
  • 341243 11/21/10 22:11 miham No beam yet! Problems with IOC persist
  • 341242 11/21/10 21:40 miham Still waiting for beam
  • 341241 11/21/10 20:20 LeRos Target Empty
  • 341240 11/21/10 20:19 LeRose Target Screens
  • 341239 11/21/10 20:07 A. Camsonne could communicate with hlal00
  • 341238 11/21/10 19:46 miham HRS-L slow controls unavailable
  • 341237 11/21/10 19:16 miham Slow control Communication Problems
  • 341236 11/21/10 19:12 auto_snap Radiation and Ozone snapshot
  • 341235 11/21/10 19:12 auto_snap Miscellaneous snapshot
  • 341234 11/21/10 19:12 auto_snap Hall A Tools snapshot
  • 341233 11/21/10 19:12 auto_snap Compton Polarimeter snapshot
  • 341232 11/21/10 19:12 auto_snap Beam Position and Current snapshot
  • 341231 11/21/10 19:10 a-onl End_of_Run_8954
  • 341230 11/21/10 19:09 a-onl Start_Run_8954,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341229 11/21/10 19:05 S. Sirca Scalers at end of run 8953
  • 341228 11/21/10 19:05 S. Sirca ROCs, coda at end of run 8953
  • 341227 11/21/10 18:41 LeRose Target Loops 1 & 2
  • 341226 11/21/10 18:25 LeRose Target Loop 2
  • 341225 11/21/10 18:18 auto_snap Radiation and Ozone snapshot
  • 341224 11/21/10 18:18 auto_snap Miscellaneous snapshot
  • 341223 11/21/10 18:18 auto_snap Hall A Tools snapshot
  • 341222 11/21/10 18:18 auto_snap Compton Polarimeter snapshot
  • 341221 11/21/10 18:17 auto_snap Beam Position and Current snapshot
  • 341220 11/21/10 18:13 LeRose Target Loop 1
  • 341219 11/21/10 18:06 a-onl Start_Run_8953,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341218 11/21/10 18:05 S. Sirca Scalers at end of run 8952
  • 341217 11/21/10 18:05 S. Sirca ROCs, coda at end of run 8952
  • 341216 11/21/10 18:05 a-onl End_of_Run_8952
  • 341215 11/21/10 17:06 LeRose Target Screens
  • 341214 11/21/10 17:05 a-onl Start_Run_8952,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341213 11/21/10 17:05 S. Sirca Scalers at end of run 8951
  • 341212 11/21/10 17:04 S. Sirca ROCs, coda at end of run 8951
  • 341211 11/21/10 17:04 a-onl End_of_Run_8951
  • 341210 11/21/10 16:34 miham Beam stable at the beginning of a shift
  • 341209 11/21/10 16:33 miham Calorimeter HVs (all green)
  • 341208 11/21/10 16:31 miham Calorimeter ADC Readback monitor screenshot
  • 341206 11/21/10 16:08 Lindgren Shift Summary
  • 341207 11/21/10 16:07 adaq Beam Time Accounting
  • 341205 11/21/10 16:07 a-onl Start_Run_8951,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341204 11/21/10 16:05 S. Sirca Scalers at end of run 8950
  • 341203 11/21/10 16:05 S. Sirca ROCs, coda at end of run 8950
  • 341202 11/21/10 16:05 a-onl End_of_Run_8950
  • 341200 11/21/10 15:02 a-onl Start_Run_8950,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341199 11/21/10 15:01 a-onl End_of_Run_8949
  • 341198 11/21/10 15:01 eschulte End of Run 8949 scalers
  • 341197 11/21/10 15:00 eschulte End of Run 8949 Run Control Just before eor
  • 341196 11/21/10 15:00 eschulte End of Run 8949 ROCS
  • 341195 11/21/10 14:45 eschulte Wein Angle Finder
  • 341194 11/21/10 14:29 eschulte Note on Run 8947
  • 341193 11/21/10 14:21 eschulte Run 8949 201 back!
  • 341192 11/21/10 14:03 a-onl Start_Run_8949,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341191 11/21/10 14:03 eschulte End of Run 8948 Run control at EOR
  • 341190 11/21/10 14:02 a-onl End_of_Run_8948
  • 341189 11/21/10 14:02 eschulte End of Run 8948
  • 341188 11/21/10 14:01 eschulte End of Run 8948 Various Screens
  • 341187 11/21/10 13:52 eschulte Run 8948 missing calorimeter segment 201
  • 341186 11/21/10 13:49 eschulte Run 8948 missing/improperly scaled ADC signal
  • 341185 11/21/10 13:35 eschulte End of Run 8947 run control at eor
  • 341184 11/21/10 13:35 a-onl Start_Run_8948,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341183 11/21/10 13:34 a-onl End_of_Run_8947
  • 341182 11/21/10 13:33 eschulte End of Run 8947 scalers
  • 341181 11/21/10 13:33 eschulte End of Run 8947 Run Control just before eor
  • 341180 11/21/10 13:33 eschulte End of Run 8947 ROCs
  • 341179 11/21/10 13:30 Lindgren BPM
  • 341178 11/21/10 13:29 Lindgren Calo ADCS
  • 341177 11/21/10 13:28 Lindgren Tools
  • 341176 11/21/10 13:26 lindgren calor
  • 341175 11/21/10 13:17 eschulte Shift Checklist Wein Angle
  • 341174 11/21/10 12:35 eschulte CODA Crashed Run 8946
  • 341173 11/21/10 12:35 a-onl Start_Run_8947,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341172 11/21/10 12:30 Markowitz Move tp LD2
  • 341171 11/21/10 12:29 a-onl Start_Run_8946,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341170 11/21/10 12:25 Lindgren Mid shift summary
  • 341169 11/21/10 12:21 eschulte End of Run 8945 Run control at EOR
  • 341168 11/21/10 12:20 a-onl End_of_Run_8945
  • 341167 11/21/10 12:19 eschulte End of Run 8945 scalers
  • 341166 11/21/10 12:19 eschulte End of Run 8945 Run Control Just Before EOR
  • 341165 11/21/10 12:18 eschulte End of Run 8945 ROCS
  • 341164 11/21/10 11:49 a-onl Start_Run_8945,Run_type=Optics,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=ps3=ps4=1 ps5=0
  • 341163 11/21/10 11:49 a-onl End_of_Run_8937
  • 341162 11/21/10 11:42 eschulte End of Run 8944 at EOR
  • 341161 11/21/10 11:41 a-onl End_of_Run_8944
  • 341160 11/21/10 11:41 eschulte End of Run 8944 scalers
  • 341159 11/21/10 11:40 eschulte End of Run 8944 Run Control Just befor EOR
  • 341158 11/21/10 11:40 eschulte End of Run 8944 ROCS
  • 341157 11/21/10 11:15 a-onl Start_Run_8944,Run_type=Other,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger Efficiency ps3=ps4=1 ps5=0
  • 341156 11/21/10 11:12 auto_snap Radiation and Ozone snapshot
  • 341155 11/21/10 11:12 auto_snap Miscellaneous snapshot
  • 341154 11/21/10 11:12 auto_snap Hall A Tools snapshot
  • 341153 11/21/10 11:12 auto_snap Compton Polarimeter snapshot
  • 341152 11/21/10 11:12 auto_snap Beam Position and Current snapshot
  • 341151 11/21/10 11:11 eschulte End of Run 8943 Run Control at EOR
  • 341150 11/21/10 11:11 a-onl End_of_Run_8943
  • 341149 11/21/10 11:11 eschulte End of Run 8943 scalers
  • 341148 11/21/10 11:10 eschulte End of Run 8943 Run Control Just before EOR
  • 341147 11/21/10 11:10 eschulte End of Run 8943 ROCs
  • 341146 11/21/10 10:17 auto_snap Radiation and Ozone snapshot
  • 341145 11/21/10 10:17 auto_snap Miscellaneous snapshot
  • 341144 11/21/10 10:17 auto_snap Hall A Tools snapshot
  • 341143 11/21/10 10:17 auto_snap Compton Polarimeter snapshot
  • 341142 11/21/10 10:17 auto_snap Beam Position and Current snapshot
  • 341141 11/21/10 10:12 eschulte End of Run 8942 Run Control just after EOR
  • 341140 11/21/10 10:12 a-onl Start_Run_8943,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341139 11/21/10 10:11 a-onl End_of_Run_8942
  • 341138 11/21/10 10:11 eschulte End of Run 8942 Run Control just before EOR
  • 341137 11/21/10 10:10 eschulte End of Run 8942 scalers
  • 341136 11/21/10 10:10 eschulte End of Run 8942 ROCS
  • 341135 11/21/10 09:47 Markowitz Target screenshots
  • 341134 11/21/10 09:11 a-onl Start_Run_8942,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341133 11/21/10 09:11 eschulte End of run 8941 Run Control just after EOR
  • 341132 11/21/10 09:10 a-onl End_of_Run_8941
  • 341131 11/21/10 09:09 eschulte End of Run 8941 Scalers
  • 341130 11/21/10 09:09 eschulte End of Run 8941 Run Control Just befor EOR
  • 341129 11/21/10 09:08 eschulte End of Run 8941 ROCS
  • 341127 11/21/10 08:18 puckett detailed procedure for DVCS calo IOC reboot
  • 341128 11/21/10 08:18 adaq Beam Time Accounting
  • 341126 11/21/10 08:09 a-onl Start_Run_8941,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341125 11/21/10 08:08 eschulte End of Run 8940 Run Control just at after stopping run
  • 341124 11/21/10 08:08 Markowitz Target computer update issue
  • 341123 11/21/10 08:08 a-onl End_of_Run_8940
  • 341122 11/21/10 08:07 eschulte End of Run 8940 scalers
  • 341121 11/21/10 08:07 eschulte End of Run 8940 Run Control just befor EOR
  • 341120 11/21/10 08:05 eschulte End of Run 8940 ROC screen Shot
  • 341119 11/21/10 08:04 A. Puckett Owl shift summary
  • 341118 11/21/10 07:06 a-onl Start_Run_8940,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341117 11/21/10 07:05 a-onl End_of_Run_8939
  • 341116 11/21/10 07:05 meziani SCALERS
  • 341115 11/21/10 07:04 meziani CODA, ROCS
  • 341114 11/21/10 07:02 meziani Calorimeter
  • 341113 11/21/10 06:12 puckett DVCS calo IOC reboot and HV
  • 341112 11/21/10 06:00 a-onl Start_Run_8939,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341111 11/21/10 05:56 a-onl End_of_Run_8938
  • 341110 11/21/10 05:55 meziani Target screens snapshot
  • 341109 11/21/10 05:49 meziani SCALERS
  • 341108 11/21/10 05:48 meziani CODA, ROCS
  • 341107 11/21/10 05:20 a-onl Start_Run_8938,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341106 11/21/10 05:05 a-onl End_of_Run_8935
  • 341105 11/21/10 04:57 meziani Scalers
  • 341104 11/21/10 04:56 meziani CODA and ROCS
  • 341103 11/21/10 04:20 a-onl Start_Run_8937,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341102 11/21/10 04:17 a-onl End_of_Run_8936
  • 341101 11/21/10 04:04 meziani Target screens
  • 341100 11/21/10 04:01 meziani CODA, ROCS
  • 341099 11/21/10 03:59 meziani Scalers
  • 341098 11/21/10 03:36 puckett Compton: performed CODA reboot
  • 341097 11/21/10 03:27 a-onl Start_Run_8936,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341096 11/21/10 03:21 puckett problem with Compton rates?
  • 341095 11/21/10 03:12 auto_snap Radiation and Ozone snapshot
  • 341094 11/21/10 03:11 auto_snap Miscellaneous snapshot
  • 341093 11/21/10 03:11 auto_snap Hall A Tools snapshot
  • 341092 11/21/10 03:11 auto_snap Compton Polarimeter snapshot
  • 341091 11/21/10 03:11 auto_snap Beam Position and Current snapshot
  • 341090 11/21/10 03:02 a-onl End_of_Run_8934
  • 341089 11/21/10 03:00 meziani Scalers
  • 341088 11/21/10 02:59 meziani CODA, ROCS
  • 341087 11/21/10 02:49 A. Puckett Compton stripcharts
  • 341086 11/21/10 02:21 a-onl Start_Run_8935,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341085 11/21/10 02:17 auto_snap Radiation and Ozone snapshot
  • 341084 11/21/10 02:17 auto_snap Miscellaneous snapshot
  • 341083 11/21/10 02:17 auto_snap Hall A Tools snapshot
  • 341082 11/21/10 02:17 auto_snap Compton Polarimeter snapshot
  • 341081 11/21/10 02:17 auto_snap Beam Position and Current snapshot
  • 341080 11/21/10 02:03 a-onl End_of_Run_8929
  • 341079 11/21/10 02:02 meziani Scalers
  • 341078 11/21/10 02:01 meziani CODA, ROCS
  • 341077 11/21/10 01:01 a-onl Start_Run_8934,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341076 11/21/10 00:59 a-onl End_of_Run_8933
  • 341075 11/21/10 00:58 meziani Scalers Snapshot
  • 341074 11/21/10 00:56 Meziani Snapshots of CODA,ROCS, SCALERS
  • 341073 11/21/10 00:33 A. Puckett DVCS calo ADC readback
  • 341072 11/21/10 00:04 adaq Beam Time Accounting
  • 341071 11/21/10 00:02 miham Swing shift summary (11/20/10)
  • Saturday

  • 341070 11/20/10 23:58 a-onl Start_Run_8933,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341069 11/20/10 23:58 a-onl End_of_Run_8932
  • 341068 11/20/10 23:52 miham Calorimeter ADC Readback Monitor - Still Alive
  • 341067 11/20/10 23:21 A. Tobias Target Screenshots
  • 341066 11/20/10 23:19 A. Tobias Snapshot of Compton Stripchart
  • 341065 11/20/10 23:00 a-onl Start_Run_8932,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341064 11/20/10 23:00 S. Sirca Scalers at end of run 8931
  • 341063 11/20/10 22:59 S. Sirca ROCs, coda at end of run 8931
  • 341062 11/20/10 22:59 a-onl End_of_Run_8931
  • 341061 11/20/10 22:06 A. Camsonne iochacs down - rebooting it
  • 341060 11/20/10 22:01 a-onl Start_Run_8931,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341059 11/20/10 21:58 a-onl End_of_Run_8930
  • 341058 11/20/10 21:58 S. Sirca Scalers at end of run 8930
  • 341057 11/20/10 21:57 S. Sirca ROCs, coda at end of run 8930
  • 341056 11/20/10 21:51 a-onl Start_Run_8930,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341055 11/20/10 21:33 a-onl End_of_Run_8926
  • 341054 11/20/10 21:33 S. Sirca Scalers at end of run 8929
  • 341053 11/20/10 21:32 Sirca ROCs, coda at end of run 8929
  • 341052 11/20/10 21:19 A. Tobias Target Screenshots
  • 341051 11/20/10 21:09 miham Small change in beam current
  • 341050 11/20/10 20:51 a-onl Start_Run_8929,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341049 11/20/10 20:34 S. Sirca Scalers at end of run 8928
  • 341048 11/20/10 20:33 S. Sirca ROCs, coda at end of run 8928
  • 341047 11/20/10 20:33 a-onl End_of_Run_8928
  • 341046 11/20/10 19:20 a-onl Start_Run_8928,Run_type=Production,target_type=LH2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=first LH2 run after Moller refill
  • 341045 11/20/10 19:18 A. Tobias Target: Tweaked JT Valves after target change from LD2 to LH2
  • 341044 11/20/10 19:11 auto_snap Radiation and Ozone snapshot
  • 341043 11/20/10 19:11 auto_snap Miscellaneous snapshot
  • 341042 11/20/10 19:11 auto_snap Hall A Tools snapshot
  • 341041 11/20/10 19:11 auto_snap Compton Polarimeter snapshot
  • 341040 11/20/10 19:11 auto_snap Beam Position and Current snapshot
  • 341039 11/20/10 19:06 miham Beam is Back!!!
  • 341038 11/20/10 18:56 A. Tobias Target moved to Loop2 LH2
  • 341037 11/20/10 18:21 a-onl End_of_Run_8927
  • 341036 11/20/10 18:17 auto_snap Radiation and Ozone snapshot
  • 341035 11/20/10 18:17 auto_snap Miscellaneous snapshot
  • 341034 11/20/10 18:17 auto_snap Hall A Tools snapshot
  • 341033 11/20/10 18:17 auto_snap Compton Polarimeter snapshot
  • 341032 11/20/10 18:17 auto_snap Beam Position and Current snapshot
  • 341031 11/20/10 18:10 A. Tobias Compton Rates decreased
  • 341030 11/20/10 18:08 miham Contolled access to fill Moller
  • 341029 11/20/10 18:04 S. Sirca ROCs, coda at end of run 8927
  • 341028 11/20/10 18:04 S. Sirca Scalers at end of run 8927
  • 341027 11/20/10 17:51 a-onl Start_Run_8927,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341026 11/20/10 17:44 A. Tobias Target: Vacuum Readout getting more noisey
  • 341025 11/20/10 17:33 a-onl End_of_Run_8924
  • 341024 11/20/10 17:33 jroche Difference between upstream and downstream BCM readings at 3 and 1.5 uA
  • 341023 11/20/10 17:33 Carlos Sync plots in online GUI modified
  • 341022 11/20/10 17:32 a-onl Start_Run_8926,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341021 11/20/10 17:20 a-onl End_of_Run_8925
  • 341020 11/20/10 17:19 S. Sirca Scalers at end of run 8925
  • 341019 11/20/10 17:18 S. Sirca ROCs, coda at end of run 8925
  • 341018 11/20/10 16:51 miham Calorimeter ADC Readback Monitor Screenshot
  • 341017 11/20/10 16:46 A. Tobias Target: Opened JT Valve on Loop 2 LH2 one step
  • 341016 11/20/10 16:37 a-onl Start_Run_8925,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341015 11/20/10 16:19 a-onl End_of_Run_8922
  • 341014 11/20/10 16:18 S. Sirca Scalers at end of run 8924
  • 341013 11/20/10 16:17 S. Sirca ROCs+coda at end of run 8924
  • 341012 11/20/10 16:10 adaq Beam Time Accounting
  • 341011 11/20/10 16:06 A. Tobias Target Screen shots at beginning of Swing Shift
  • 341010 11/20/10 16:02 Lindgren Shift Summary
  • 341009 11/20/10 15:39 a-onl Start_Run_8924,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341008 11/20/10 15:34 eschulte Cerenkof shift returned to no-shift
  • 341007 11/20/10 15:17 a-onl End_of_Run_8923
  • 341006 11/20/10 15:16 eschulte End of Run 8923 Run Control just after EOR
  • 341005 11/20/10 15:15 eschulte End of Run 8923 scalers
  • 341004 11/20/10 15:14 eschulte End of Run 8923 Run control just before EOR
  • 341003 11/20/10 15:14 eschulte End of Run 8923 ROCs 14
  • 341002 11/20/10 15:13 eschulte End of Run 8923 ROCS 17 & 18
  • 341001 11/20/10 14:31 a-onl Start_Run_8923,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 341000 11/20/10 14:13 eschulte End of Run 8922 Run Control Just after EOR
  • 340999 11/20/10 14:13 a-onl End_of_Run_8921
  • 340998 11/20/10 14:12 eschulte End of Run 8922 scalers
  • 340997 11/20/10 14:12 eschulte End of Run 8922 Run Control Just befor EOR
  • 340996 11/20/10 14:11 eschulte End of Run 8922 ROC 14
  • 340995 11/20/10 14:11 eschulte End of Run 8922 ROCS 17 & 18
  • 340994 11/20/10 13:38 eschulte slight change in cerenkov spectrum run 8922
  • 340993 11/20/10 13:31 a-onl Start_Run_8922,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340992 11/20/10 13:13 a-onl End_of_Run_8920
  • 340991 11/20/10 13:12 eschulte End of Run 8921 Run Control After EOR
  • 340990 11/20/10 13:11 eschulte End of Run 8921 scalers
  • 340989 11/20/10 13:10 eschulte End of Run 8921 Run Control Before EOR
  • 340988 11/20/10 13:10 eschulte End of run 8921 ROC 14
  • 340987 11/20/10 13:09 eschulte End of Run 8921 ROCS 17 & 18
  • 340986 11/20/10 13:03 Lindgren Logging DVCS Calor
  • 340985 11/20/10 13:00 Lindgren Logging Compton
  • 340984 11/20/10 12:59 Lindgren Logging tools
  • 340983 11/20/10 12:31 a-onl Start_Run_8921,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340982 11/20/10 12:13 a-onl End_of_Run_8919
  • 340981 11/20/10 12:12 eschulte End of Run 8920 Run Control After run Stop
  • 340980 11/20/10 12:11 eschulte End of Run 8920 Scalers
  • 340979 11/20/10 12:11 eschulte End of Run 8920 Run Control Befor Run Stop
  • 340978 11/20/10 12:10 eschulte End of Run 8920 ROC 14
  • 340977 11/20/10 12:10 eschulte End of Run 8920 ROCS 17 & 18
  • 340976 11/20/10 11:30 a-onl Start_Run_8920,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340975 11/20/10 11:12 a-onl End_of_Run_8918
  • 340974 11/20/10 11:11 auto_snap Radiation and Ozone snapshot
  • 340973 11/20/10 11:11 auto_snap Miscellaneous snapshot
  • 340972 11/20/10 11:11 auto_snap Hall A Tools snapshot
  • 340971 11/20/10 11:11 auto_snap Compton Polarimeter snapshot
  • 340970 11/20/10 11:11 auto_snap Beam Position and Current snapshot
  • 340969 11/20/10 11:11 eschulte end of run 8919 run control just after run stop
  • 340968 11/20/10 11:10 eschulte end of run 8919 run control and pre-scales
  • 340967 11/20/10 11:09 eschulte end of run 8919 scalers
  • 340966 11/20/10 11:09 eschulte end of run 8919 ROC 14
  • 340965 11/20/10 11:08 eschulte End of Run 8919 Rocs 17 & 18
  • 340964 11/20/10 10:50 Carlos bcm_readings
  • 340963 11/20/10 10:31 eschulte Run 8919 Online Analysis
  • 340962 11/20/10 10:28 a-onl Start_Run_8919,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340961 11/20/10 10:17 auto_snap Radiation and Ozone snapshot
  • 340960 11/20/10 10:17 auto_snap Miscellaneous snapshot
  • 340959 11/20/10 10:16 auto_snap Hall A Tools snapshot
  • 340958 11/20/10 10:16 auto_snap Compton Polarimeter snapshot
  • 340957 11/20/10 10:16 auto_snap Beam Position and Current snapshot
  • 340956 11/20/10 10:16 A. Camsonne Most like due to asynchronous insertion of scalers
  • 340955 11/20/10 10:09 a-onl End_of_Run_8917
  • 340954 11/20/10 10:09 eschulte End of run 8918 run control (after stop)
  • 340953 11/20/10 10:08 eschulte End of Run 8918 Run Control
  • 340952 11/20/10 10:06 eschulte End of Run 8918 Scalers
  • 340951 11/20/10 10:05 eschulte End of Run 8918 ROCS 14
  • 340950 11/20/10 10:05 eschulte End of Run 8918 ROCS 17 & 18
  • 340949 11/20/10 09:35 eschulte Run 8918 prescales & run control screen
  • 340948 11/20/10 09:34 eschulte ROC 14's again Run 8918
  • 340947 11/20/10 09:34 eschulte Run 8918 Scalar Rates
  • 340946 11/20/10 09:33 eschulte Run 8918 ROC Screens Cont.
  • 340945 11/20/10 09:33 eschulte Run 8918 ROC Screens
  • 340944 11/20/10 09:29 a-onl Start_Run_8918,Run_type=Production,target_type=LD2,beam_e0=4.455,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340943 11/20/10 09:27 eschulte Run 8918 Online Analysis: Lost part of Lumis?
  • 340942 11/20/10 09:22 eschulte L-arm S0 ADC cutoff?
  • 340941 11/20/10 09:21 eschulte Run 8918 Online Notes
  • 340940 11/20/10 09:10 a-onl End_of_Run_8916
  • 340939 11/20/10 09:09 eschulte RunControl
  • 340938 11/20/10 09:08 eschulte ROC screen 14
  • 340937 11/20/10 09:07 eschulte Roc screens
  • 340936 11/20/10 09:07 eschulte Run 8917 scalers
  • 340935 11/20/10 08:48 eschulte Run 8917 End of Online Analysis Plots!
  • 340934 11/20/10 08:42 eschulte Pedestal shifting warnings run 8917 2
  • 340933 11/20/10 08:37 eschulte online analysis 8917 pedestal warning 1
  • 340932 11/20/10 08:35 eschulte online for run 8917
  • 340931 11/20/10 08:25 a-onl Start_Run_8917,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340930 11/20/10 08:10 eschulte Run 8916 Screen Shots
  • 340929 11/20/10 08:07 a-onl End_of_Run_8914
  • 340927 11/20/10 08:03 efuchey 10-20-10 Owl shift summary
  • 340928 11/20/10 08:03 adaq Beam Time Accounting
  • 340926 11/20/10 08:00 efuchey Compton CODA crash
  • 340925 11/20/10 07:39 efuchey Re: Inconsistencies for S0 L ADC spectra
  • 340924 11/20/10 07:30 efuchey Beam position distribution became normal
  • 340923 11/20/10 07:24 a-onl Start_Run_8916,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340922 11/20/10 07:20 a-onl End_of_Run_8915
  • 340921 11/20/10 07:05 efuchey end of run 8915 part 2
  • 340920 11/20/10 07:05 efuchey end of run 8915
  • 340919 11/20/10 06:40 efuchey beam spread.
  • 340918 11/20/10 06:25 a-onl Start_Run_8915,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340917 11/20/10 06:08 meziani Target screens
  • 340916 11/20/10 06:08 a-onl End_of_Run_8913
  • 340915 11/20/10 06:06 efuchey end of run 8914 part 2
  • 340914 11/20/10 06:06 efuchey end of run 8914
  • 340913 11/20/10 05:39 efuchey Hall A tools
  • 340912 11/20/10 05:31 efuchey Calo base currents
  • 340911 11/20/10 05:22 a-onl Start_Run_8914,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340910 11/20/10 05:04 a-onl End_of_Run_8912
  • 340909 11/20/10 05:03 efuchey end of run 8913 part 2
  • 340908 11/20/10 05:02 efuchey end of run 8913
  • 340907 11/20/10 04:20 a-onl Start_Run_8913,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340906 11/20/10 04:02 a-onl End_of_Run_8911
  • 340905 11/20/10 04:01 efuchey end of run 8912 part 2
  • 340904 11/20/10 04:00 efuchey End of run 8912
  • 340903 11/20/10 03:57 meziani Target screens
  • 340902 11/20/10 03:20 a-onl Start_Run_8912,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340901 11/20/10 03:11 auto_snap Radiation and Ozone snapshot
  • 340900 11/20/10 03:11 auto_snap Miscellaneous snapshot
  • 340899 11/20/10 03:11 auto_snap Hall A Tools snapshot
  • 340898 11/20/10 03:10 auto_snap Compton Polarimeter snapshot
  • 340897 11/20/10 03:10 auto_snap Beam Position and Current snapshot
  • 340896 11/20/10 03:03 a-onl End_of_Run_8910
  • 340895 11/20/10 03:02 efuchey end of run 8911 part 2
  • 340894 11/20/10 03:01 efuchey end of run 8911
  • 340893 11/20/10 03:00 efuchey Compton rates
  • 340892 11/20/10 02:21 a-onl Start_Run_8911,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340891 11/20/10 02:16 auto_snap Radiation and Ozone snapshot
  • 340890 11/20/10 02:16 auto_snap Miscellaneous snapshot
  • 340889 11/20/10 02:16 auto_snap Hall A Tools snapshot
  • 340888 11/20/10 02:16 auto_snap Compton Polarimeter snapshot
  • 340887 11/20/10 02:16 auto_snap Beam Position and Current snapshot
  • 340886 11/20/10 02:03 a-onl End_of_Run_8909
  • 340885 11/20/10 02:01 efuchey end of run 8910 part 2
  • 340884 11/20/10 02:01 efuchey end of run 8910
  • 340883 11/20/10 01:24 a-onl Start_Run_8910,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340882 11/20/10 01:24 a-onl End_of_Run_8908
  • 340881 11/20/10 01:05 efuchey end of run 8909 part 2
  • 340880 11/20/10 01:04 efuchey End of run 8909
  • 340879 11/20/10 00:38 efuchey Calo ACDs for run at 1.5 uA
  • 340878 11/20/10 00:31 efuchey Inconsistencies for S0 L ADC spectra
  • 340877 11/20/10 00:21 a-onl Start_Run_8909,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3.0,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340876 11/20/10 00:10 a-onl End_of_Run_8907
  • 340875 11/20/10 00:02 miham Run #8908 ROC Screenshots
  • 340873 11/20/10 00:00 S. Sirca Swing shift SUMMARY
  • Friday

  • 340874 11/19/10 23:59 adaq Beam Time Accounting
  • 340872 11/19/10 23:59 vahe Target GUI snapshot
  • 340871 11/19/10 23:57 a-onl Start_Run_8908,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=1.5,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Special run at lower current (1.5uA) than usual
  • 340870 11/19/10 23:38 a-onl End_of_Run_8906
  • 340869 11/19/10 23:36 miham Run #8907 ROC Screenshots
  • 340868 11/19/10 22:59 a-onl Start_Run_8907,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340867 11/19/10 22:39 a-onl End_of_Run_8905
  • 340866 11/19/10 22:37 miham Run #8906 ROC screenshots
  • 340865 11/19/10 21:59 vahe Target GUI snapshot
  • 340864 11/19/10 21:54 a-onl Start_Run_8906,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340863 11/19/10 21:35 a-onl End_of_Run_8904
  • 340862 11/19/10 21:34 miham Run #8905 ROC Screenshots
  • 340861 11/19/10 20:59 a-onl Start_Run_8905,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340860 11/19/10 20:37 a-onl End_of_Run_8903
  • 340859 11/19/10 20:35 miham Run #8904 ROC Screenshots
  • 340858 11/19/10 20:06 vahe Target GUI snapshot
  • 340857 11/19/10 19:54 a-onl Start_Run_8904,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340856 11/19/10 19:37 a-onl End_of_Run_8901
  • 340855 11/19/10 19:35 miham Run #8903 ROC Screenshots
  • 340854 11/19/10 19:21 vahe 15K supply temperature fluctuations
  • 340853 11/19/10 19:10 auto_snap Radiation and Ozone snapshot
  • 340852 11/19/10 19:10 auto_snap Miscellaneous snapshot
  • 340851 11/19/10 19:10 auto_snap Hall A Tools snapshot
  • 340850 11/19/10 19:10 auto_snap Compton Polarimeter snapshot
  • 340849 11/19/10 19:10 auto_snap Beam Position and Current snapshot
  • 340848 11/19/10 18:54 a-onl Start_Run_8903,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340847 11/19/10 18:49 a-onl End_of_Run_8902
  • 340846 11/19/10 18:38 Carlos Hall C high current
  • 340845 11/19/10 18:35 miham Run #8902 ROC ScreenShots
  • 340844 11/19/10 18:30 S. Sirca Sudden increase in Compton rates
  • 340843 11/19/10 18:28 S. Sirca Crazy BCM readouts
  • 340842 11/19/10 18:16 auto_snap Radiation and Ozone snapshot
  • 340841 11/19/10 18:16 auto_snap Miscellaneous snapshot
  • 340840 11/19/10 18:16 auto_snap Hall A Tools snapshot
  • 340839 11/19/10 18:16 auto_snap Compton Polarimeter snapshot
  • 340838 11/19/10 18:16 auto_snap Beam Position and Current snapshot
  • 340837 11/19/10 17:57 vahe Target GUI snapshot
  • 340836 11/19/10 17:52 a-onl Start_Run_8902,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340835 11/19/10 17:34 miham Run #8901 ROC screenshots
  • 340834 11/19/10 16:58 jroche BCMs gain jump??
  • 340833 11/19/10 16:48 a-onl Start_Run_8901,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340832 11/19/10 16:29 a-onl End_of_Run_8898
  • 340831 11/19/10 16:29 a-onl End_of_Run_8900
  • 340830 11/19/10 16:27 miham Run #8900 ROC Screenshots
  • 340829 11/19/10 16:04 vahe Target GUI snapshot
  • 340828 11/19/10 16:03 hyde Day shift summary
  • 340827 11/19/10 16:00 adaq Beam Time Accounting
  • 340826 11/19/10 15:43 a-onl Start_Run_8900,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340825 11/19/10 15:40 hyde optics reconstruction
  • 340824 11/19/10 15:27 a-onl End_of_Run_8899
  • 340823 11/19/10 15:24 Markowitz back to LD2
  • 340822 11/19/10 15:11 hyde ABU / BANU
  • 340821 11/19/10 14:57 markowitz Loop 2 temperature spike
  • 340820 11/19/10 14:34 a-onl Start_Run_8899,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Optics Check ps3=ps4=1
  • 340819 11/19/10 14:16 a-onl Start_Run_8898,Run_type=Optics,target_type=Optics: multi 1mm C foils,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Optics Check ps3=ps4=1
  • 340818 11/19/10 14:15 a-onl End_of_Run_8896
  • 340817 11/19/10 14:08 a-onl End_of_Run_8897
  • 340816 11/19/10 14:07 Markowitz moved to optics target
  • 340815 11/19/10 14:01 a-onl Start_Run_8897,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Efficiency Check ps3=ps4=1
  • 340814 11/19/10 13:54 hyde Beam with Hall C off
  • 340813 11/19/10 13:46 a-onl End_of_Run_8895
  • 340812 11/19/10 13:21 hyde Beam steps correlated to Hall C?
  • 340811 11/19/10 13:17 hyde Compton
  • 340810 11/19/10 13:07 hyde noise on bpm
  • 340809 11/19/10 13:04 hyde Hall A Tools Snapshot
  • 340808 11/19/10 13:02 a-onl Start_Run_8896,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II low
  • 340807 11/19/10 13:02 hyde T5 rates
  • 340806 11/19/10 12:44 markowitz LD2 target inserted
  • 340805 11/19/10 12:43 hyde harp
  • 340804 11/19/10 12:08 hyde calo cosmic hv off
  • 340803 11/19/10 12:08 Fansler Moller power supply
  • 340802 11/19/10 12:07 a-onl Start_Run_8895,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test with cosmics paddles 775 V
  • 340801 11/19/10 12:03 a-onl End_of_Run_8893
  • 340800 11/19/10 11:59 a-onl End_of_Run_8892
  • 340799 11/19/10 11:59 a-onl Start_Run_8893,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test with cosmics paddles 775 V
  • 340798 11/19/10 11:59 hyde ready for beam
  • 340797 11/19/10 11:56 A. Camsone Test with cosmics paddles 3KHz random input rates
  • 340796 11/19/10 11:10 auto_snap Radiation and Ozone snapshot
  • 340795 11/19/10 11:10 auto_snap Miscellaneous snapshot
  • 340794 11/19/10 11:10 auto_snap Hall A Tools snapshot
  • 340793 11/19/10 11:10 auto_snap Compton Polarimeter snapshot
  • 340792 11/19/10 11:10 auto_snap Beam Position and Current snapshot
  • 340791 11/19/10 10:47 O.Glamazdin Moller magnet has been filled
  • 340790 11/19/10 10:46 roblin prebuncher hiccups and resteering past it
  • 340789 11/19/10 10:18 hyde target gas alarm
  • 340788 11/19/10 10:16 auto_snap Radiation and Ozone snapshot
  • 340787 11/19/10 10:16 auto_snap Miscellaneous snapshot
  • 340786 11/19/10 10:16 auto_snap Hall A Tools snapshot
  • 340785 11/19/10 10:15 auto_snap Compton Polarimeter snapshot
  • 340784 11/19/10 10:15 auto_snap Beam Position and Current snapshot
  • 340783 11/19/10 10:09 Markowitz target to empty position
  • 340782 11/19/10 10:07 A. Camsonne Fan and heater back on
  • 340781 11/19/10 09:33 hyde target fans
  • 340780 11/19/10 09:15 Markowitz More target temperatures
  • 340779 11/19/10 09:12 A. Camsonne Target fan and heater turned manually off
  • 340778 11/19/10 09:09 hyde Controlled Access
  • 340777 11/19/10 08:58 ellie Run 8892 - Lumi4 Rate vs. Event - Not like the others
  • 340776 11/19/10 08:42 Markowitz Hall A target
  • 340775 11/19/10 08:41 hyde target
  • 340774 11/19/10 08:33 hyde target strip charts
  • 340773 11/19/10 08:18 a-onl End_of_Run_8891
  • 340772 11/19/10 08:17 ellie end of run screen shot
  • 340771 11/19/10 08:17 hyde Beam off for target
  • 340770 11/19/10 08:00 adaq Beam Time Accounting
  • 340769 11/19/10 07:56 Jin Huang Shift Summary : Nov 19 Owl
  • 340768 11/19/10 07:54 a-onl Start_Run_8892,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production deuterium 4 pass
  • 340767 11/19/10 07:44 R. Michaels re: cryo, it's due to hall C
  • 340766 11/19/10 07:34 itard entries end of run
  • 340765 11/19/10 07:29 R. Michaels unstable cryo (getting worse)
  • 340764 11/19/10 06:59 A. Camsonne Compton background dropped to 50 Hz/uA
  • 340763 11/19/10 06:57 R. Michaels cryo instabilities
  • 340762 11/19/10 06:51 a-onl Start_Run_8891,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production deuterium 4 pass
  • 340761 11/19/10 06:46 R. Michaels ET params adjusted, try to smoothen deadtime
  • 340760 11/19/10 06:32 A. Camsonne Pion rejector bad block ?
  • 340759 11/19/10 06:29 itard entries end of run
  • 340758 11/19/10 06:12 A. Camsonne Redownloaded the trigger
  • 340757 11/19/10 06:01 A. Camsonne Coincidence peaks for blocks at small angle
  • 340756 11/19/10 05:59 Jin Huang PbF2 ADC Readback, Ch182 is low
  • 340755 11/19/10 05:52 A. Camsonne Run 8890 production test writing to DVCSDAQ2
  • 340754 11/19/10 05:40 a-onl End_of_Run_8889
  • 340753 11/19/10 05:29 atrig Trigger-Download
  • 340752 11/19/10 05:27 itard entries end of run
  • 340751 11/19/10 05:22 A. Camsonne Calorimeter golden run updated to run 8889
  • 340750 11/19/10 04:45 a-onl Start_Run_8889,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production deuterium 4 pass
  • 340749 11/19/10 04:26 a-onl End_of_Run_8887
  • 340748 11/19/10 04:26 a-onl Start_Run_8888,Run_type=Cosmics,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340747 11/19/10 04:11 a-onl Start_Run_8887,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production 4 pass deuterium
  • 340746 11/19/10 04:01 a-onl End_of_Run_8886
  • 340745 11/19/10 04:01 a-onl End_of_Run_8885
  • 340744 11/19/10 03:58 itard entries end of run
  • 340743 11/19/10 03:23 a-onl Start_Run_8886,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production 4 pass deuterium
  • 340742 11/19/10 03:10 auto_snap Radiation and Ozone snapshot
  • 340741 11/19/10 03:10 auto_snap Miscellaneous snapshot
  • 340740 11/19/10 03:10 auto_snap Hall A Tools snapshot
  • 340739 11/19/10 03:10 auto_snap Compton Polarimeter snapshot
  • 340738 11/19/10 03:10 auto_snap Beam Position and Current snapshot
  • 340737 11/19/10 03:05 itard entries end of run
  • 340736 11/19/10 02:15 auto_snap Radiation and Ozone snapshot
  • 340735 11/19/10 02:15 auto_snap Miscellaneous snapshot
  • 340734 11/19/10 02:15 auto_snap Hall A Tools snapshot
  • 340733 11/19/10 02:15 auto_snap Compton Polarimeter snapshot
  • 340732 11/19/10 02:15 auto_snap Beam Position and Current snapshot
  • 340731 11/19/10 01:59 a-onl Start_Run_8885,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production 4 pass deuterium
  • 340730 11/19/10 01:58 a-onl End_of_Run_8884
  • 340729 11/19/10 01:57 itard entries end of run
  • 340728 11/19/10 01:35 A. Camsonne Fluctuations in L1A rates still here with Hall C off
  • 340727 11/19/10 01:26 Jin Huang LeCroy HV alarm
  • 340726 11/19/10 01:09 A. Camsonne Large deadtime fluctuations back
  • 340725 11/19/10 01:06 a-onl Start_Run_8884,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production 4 pass deuterium
  • 340724 11/19/10 00:58 R. Michaels move target to LD2
  • 340723 11/19/10 00:56 a-onl End_of_Run_8883
  • 340722 11/19/10 00:55 Jin Huang Hall A Tools; LHRS Q2 L% reached 100%?
  • 340721 11/19/10 00:38 A. Camsonne Hall C at high current starting 22:55
  • 340719 11/19/10 00:00 S. Sirca Swing shift SUMMARY
  • Thursday

  • 340720 11/18/10 23:59 adaq Beam Time Accounting
  • 340718 11/18/10 23:55 a-onl Start_Run_8883,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340717 11/18/10 23:54 a-onl End_of_Run_8882
  • 340716 11/18/10 23:53 miham Run #8882 ROC Screenshots
  • 340715 11/18/10 23:42 puckett target snaps
  • 340714 11/18/10 23:01 A. Camsonne Compton rates after beam studies look OK
  • 340713 11/18/10 22:54 a-onl Start_Run_8882,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340712 11/18/10 22:42 A. Camsonne Testing end of run scripts
  • 340711 11/18/10 22:40 a-onl End_of_Run_8881
  • 340710 11/18/10 22:39 a-onl End_of_Run_8881
  • 340709 11/18/10 22:36 miham Run #8881 ROC Screenshots
  • 340708 11/18/10 22:35 a-onl End_of_Run_8881
  • 340707 11/18/10 22:01 a-onl Start_Run_8881,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340706 11/18/10 22:00 miham Run #8880 ROC Screenshots
  • 340705 11/18/10 21:25 puckett target snapshots
  • 340704 11/18/10 21:02 a-onl Start_Run_8880,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340703 11/18/10 21:01 a-onl End_of_Run_8879
  • 340702 11/18/10 21:00 miham Run #8879 ROC Screenshots
  • 340701 11/18/10 20:02 a-onl Start_Run_8879,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340700 11/18/10 20:01 a-onl End_of_Run_8878
  • 340699 11/18/10 20:01 miham Run #8878 ROC Screenshots
  • 340698 11/18/10 19:09 auto_snap Radiation and Ozone snapshot
  • 340697 11/18/10 19:09 auto_snap Miscellaneous snapshot
  • 340696 11/18/10 19:09 auto_snap Hall A Tools snapshot
  • 340695 11/18/10 19:09 auto_snap Compton Polarimeter snapshot
  • 340694 11/18/10 19:09 auto_snap Beam Position and Current snapshot
  • 340693 11/18/10 19:04 jroche Tiefenbach energies at 4 passes
  • 340692 11/18/10 19:01 a-onl Start_Run_8878,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=
  • 340691 11/18/10 19:00 a-onl End_of_Run_8877
  • 340690 11/18/10 18:44 Carlos Calibration block 89
  • 340689 11/18/10 18:28 miham Run #8877 is at 26.55
  • 340688 11/18/10 18:15 auto_snap Radiation and Ozone snapshot
  • 340687 11/18/10 18:15 auto_snap Miscellaneous snapshot
  • 340686 11/18/10 18:15 auto_snap Hall A Tools snapshot
  • 340685 11/18/10 18:15 auto_snap Compton Polarimeter snapshot
  • 340684 11/18/10 18:15 auto_snap Beam Position and Current snapshot
  • 340683 11/18/10 17:59 a-onl Start_Run_8877,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=First production after beam recovery
  • 340682 11/18/10 17:58 puckett Moved target to LH2 15 cm loop 2
  • 340681 11/18/10 17:57 S. Sirca Harp scans
  • 340680 11/18/10 17:26 a-onl End_of_Run_8876
  • 340679 11/18/10 17:05 A. Camsonne Calorimeter cosmics paddle HV on at 650 V for cosmics data taking
  • 340678 11/18/10 17:04 hyde Deadtime from ROC17 scalers
  • 340677 11/18/10 16:30 A. Camsonne Hall status switched to sweep in progress
  • 340676 11/18/10 16:29 a-onl Start_Run_8876,Run_type=Cosmics,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS Calorimeter cosmics run - collimator in
  • 340674 11/18/10 16:09 A. Deur Day Shift summary
  • 340675 11/18/10 16:08 adaq Beam Time Accounting
  • 340673 11/18/10 15:35 fansler Moeller
  • 340672 11/18/10 15:04 hyde DVCS clock to ROC17 scaler
  • 340671 11/18/10 15:01 a-onl End_of_Run_8875
  • 340670 11/18/10 14:13 A. Camsonne DVCS sync signal replugged
  • 340669 11/18/10 14:08 a-onl Start_Run_8875,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=test sync ARS
  • 340668 11/18/10 14:06 a-onl End_of_Run_8874
  • 340667 11/18/10 13:53 a-onl Start_Run_8874,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=test sync ARS
  • 340666 11/18/10 13:52 A. Camsonne ARS pedestal run 8873
  • 340665 11/18/10 13:50 a-onl Start_Run_8873,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Pedestal ARS - clock 183 Hz
  • 340664 11/18/10 13:49 A. Camsonne DVCS LED pulser in ROC17 scaler channel 30 ( starting at 1 )
  • 340663 11/18/10 13:48 a-onl End_of_Run_8872
  • 340662 11/18/10 13:46 a-onl Start_Run_8872,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test after firmware change
  • 340661 11/18/10 13:44 a-onl End_of_Run_8871
  • 340660 11/18/10 13:42 a-onl Start_Run_8871,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test after firmware change
  • 340659 11/18/10 13:41 hyde high clock rate
  • 340658 11/18/10 13:14 hyde 6msr Collimator in
  • 340657 11/18/10 12:25 Carlos New setting
  • 340656 11/18/10 12:23 Carlos Cable changed
  • 340655 11/18/10 12:23 hyde intermittant deadtime / daq interuptions
  • 340654 11/18/10 12:21 a-onl Start_Run_8870,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test after firmware change
  • 340653 11/18/10 12:19 A. Camsonne Logic module firmware loaded for S2 cer trigger
  • 340652 11/18/10 11:09 auto_snap Radiation and Ozone snapshot
  • 340651 11/18/10 11:09 auto_snap Miscellaneous snapshot
  • 340650 11/18/10 11:09 auto_snap Hall A Tools snapshot
  • 340649 11/18/10 11:09 auto_snap Compton Polarimeter snapshot
  • 340648 11/18/10 11:09 auto_snap Beam Position and Current snapshot
  • 340647 11/18/10 10:57 Kalyan Hall in Controlled access
  • 340646 11/18/10 10:47 a-onl End_of_Run_8869
  • 340645 11/18/10 10:43 a-onl Start_Run_8869,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test DVCSDAQ2
  • 340644 11/18/10 10:33 a-onl End_of_Run_8868
  • 340643 11/18/10 10:15 auto_snap Radiation and Ozone snapshot
  • 340642 11/18/10 10:15 auto_snap Miscellaneous snapshot
  • 340641 11/18/10 10:15 auto_snap Hall A Tools snapshot
  • 340640 11/18/10 10:15 auto_snap Compton Polarimeter snapshot
  • 340639 11/18/10 10:15 auto_snap Beam Position and Current snapshot
  • 340638 11/18/10 09:59 hyde more ROC17 scalers
  • 340637 11/18/10 09:54 a-onl Start_Run_8868,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test DVCSDAQ2
  • 340636 11/18/10 09:52 a-onl End_of_Run_8867
  • 340635 11/18/10 09:52 a-onl Start_Run_8867,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test DVCSDAQ2
  • 340634 11/18/10 09:50 a-onl End_of_Run_8866
  • 340633 11/18/10 09:34 a-onl End_of_Run_8864
  • 340632 11/18/10 09:17 a-onl Start_Run_8863,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Test DVCSDAQ2
  • 340631 11/18/10 09:14 a-onl End_of_Run_8862
  • 340630 11/18/10 08:42 a-onl End_of_Run_8859
  • 340629 11/18/10 08:22 A. Camsonne Scalers rate 8859
  • 340628 11/18/10 08:21 A. Camsonne Test run with higher pulser rate for dead time check 8859
  • 340627 11/18/10 08:20 a-onl Start_Run_8859,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Elastic setting - Higher pulse rate - Pulser 183 Hz
  • 340626 11/18/10 08:20 Hisham Owl shift summary
  • 340625 11/18/10 08:19 a-onl End_of_Run_8858
  • 340624 11/18/10 08:19 a-onl Start_Run_8858,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Elastic setting - Higher pulse rate - Pulser tweak
  • 340623 11/18/10 08:16 a-onl End_of_Run_8857
  • 340622 11/18/10 08:00 adaq Beam Time Accounting
  • 340621 11/18/10 07:57 vahe Target GUI snapshot
  • 340620 11/18/10 07:19 a-onl Start_Run_8857,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.981,leftarm_theta=40.,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Elastic setting
  • 340619 11/18/10 07:01 itard entries end of run
  • 340618 11/18/10 06:59 a-onl End_of_Run_8856
  • 340617 11/18/10 06:15 a-onl Start_Run_8856,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Elastic setting
  • 340616 11/18/10 06:14 a-onl End_of_Run_8855
  • 340615 11/18/10 06:14 Itard Scalers during run# 8855
  • 340614 11/18/10 06:13 itard entries end of run
  • 340613 11/18/10 06:03 Carlos Block 89
  • 340612 11/18/10 06:03 vahe Target GUI snapshot
  • 340611 11/18/10 05:25 a-onl Start_Run_8855,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=5.5,comment_text=Elastic setting
  • 340610 11/18/10 05:18 a-onl End_of_Run_8854
  • 340609 11/18/10 05:17 itard run#8854 scalers
  • 340608 11/18/10 05:16 itard end of run #8854
  • 340607 11/18/10 05:03 Hisham DVCS DC levels
  • 340606 11/18/10 04:24 itard scaler
  • 340605 11/18/10 04:23 a-onl Start_Run_8854,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting
  • 340604 11/18/10 04:22 a-onl End_of_Run_8853
  • 340603 11/18/10 04:21 itard entries end of run
  • 340602 11/18/10 04:00 vahe Target GUI snapshot
  • 340601 11/18/10 03:46 Carlos Calibration change
  • 340600 11/18/10 03:37 Carlos Calorimeter HV change
  • 340599 11/18/10 03:27 a-onl Start_Run_8853,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting
  • 340598 11/18/10 03:26 a-onl End_of_Run_8852
  • 340597 11/18/10 03:20 a-onl Start_Run_8852,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting
  • 340596 11/18/10 03:18 a-onl Start_Run_8851,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340595 11/18/10 03:17 a-onl End_of_Run_8851
  • 340594 11/18/10 03:10 Hisham Theta_LHRS=36.3 deg, p=2.245 GeV
  • 340593 11/18/10 03:09 auto_snap Radiation and Ozone snapshot
  • 340592 11/18/10 03:09 auto_snap Miscellaneous snapshot
  • 340591 11/18/10 03:09 auto_snap Hall A Tools snapshot
  • 340590 11/18/10 03:09 auto_snap Compton Polarimeter snapshot
  • 340589 11/18/10 03:09 auto_snap Beam Position and Current snapshot
  • 340588 11/18/10 02:48 Hisham Controlled access
  • 340587 11/18/10 02:47 itard scaler
  • 340586 11/18/10 02:46 a-onl End_of_Run_8850
  • 340585 11/18/10 02:46 itard entries end of run
  • 340584 11/18/10 02:14 auto_snap Radiation and Ozone snapshot
  • 340583 11/18/10 02:14 auto_snap Miscellaneous snapshot
  • 340582 11/18/10 02:14 auto_snap Hall A Tools snapshot
  • 340581 11/18/10 02:14 auto_snap Compton Polarimeter snapshot
  • 340580 11/18/10 02:14 auto_snap Beam Position and Current snapshot
  • 340579 11/18/10 02:05 a-onl Start_Run_8850,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340578 11/18/10 02:03 Hisham Run 8849 scalers snapshot
  • 340577 11/18/10 02:03 a-onl End_of_Run_8849
  • 340576 11/18/10 02:02 itard entries end of run
  • 340575 11/18/10 02:01 vahe Target GUI snapshot
  • 340574 11/18/10 01:17 a-onl Start_Run_8849,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340573 11/18/10 00:39 Hisham Theta_LHRS=37.86 deg
  • 340572 11/18/10 00:15 vahe Target GUI snapshot
  • 340571 11/18/10 00:03 a-onl End_of_Run_8848
  • 340569 11/18/10 00:01 miham Swing Shift Summary (11/17/10)
  • 340570 11/18/10 00:00 adaq Beam Time Accounting
  • Wednesday

  • 340568 11/17/10 23:06 a-onl Start_Run_8848,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 3
  • 340567 11/17/10 23:05 a-onl End_of_Run_8847
  • 340566 11/17/10 23:05 miham One of N2 levels in HRS-L is low!
  • 340565 11/17/10 22:08 a-onl Start_Run_8847,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.245,leftarm_theta=36.30,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 3
  • 340564 11/17/10 22:08 miham HRS-L set to new momentum of 2.245GeV/c
  • 340563 11/17/10 22:07 miham HRS-L moved to new angle of 36.3deg
  • 340562 11/17/10 21:58 miham Q2 and Q3 are being cycled
  • 340561 11/17/10 21:55 a-onl End_of_Run_8846
  • 340560 11/17/10 21:47 miham Run #8854 is missing End-of-Run file
  • 340559 11/17/10 21:42 puckett target snapshots
  • 340558 11/17/10 21:38 miham Two Cable got detached
  • 340557 11/17/10 21:17 a-onl Start_Run_8846,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340556 11/17/10 21:16 S. Sirca ROC snapshots for run 8845
  • 340555 11/17/10 20:46 miham Calorimeter ADC Readback Monitor Screenshot
  • 340554 11/17/10 20:26 a-onl Start_Run_8845,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340553 11/17/10 20:24 a-onl Start_Run_8844,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=2.131,leftarm_theta=37.86,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 2
  • 340552 11/17/10 20:21 miham HRS-L set to new momentum p=2.131GeV/c
  • 340551 11/17/10 20:17 miham HRS-L moved to new angle
  • 340550 11/17/10 20:01 miham Cycling Q2 and Q3 for new momentum setting
  • 340549 11/17/10 19:59 a-onl End_of_Run_8843
  • 340548 11/17/10 19:59 S. Sirca ROC snapshots for run 8843
  • 340547 11/17/10 19:59 a-onl Start_Run_8841,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=check standard DAQ
  • 340546 11/17/10 19:14 miham Run #8842 is Junk
  • 340545 11/17/10 19:09 A. Camsonne Logic module firmware loaded for S2 S0 trigger
  • 340544 11/17/10 19:09 auto_snap Radiation and Ozone snapshot
  • 340543 11/17/10 19:09 auto_snap Miscellaneous snapshot
  • 340542 11/17/10 19:08 auto_snap Hall A Tools snapshot
  • 340541 11/17/10 19:08 auto_snap Compton Polarimeter snapshot
  • 340540 11/17/10 19:08 auto_snap Beam Position and Current snapshot
  • 340539 11/17/10 18:58 a-onl Start_Run_8843,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Elastic setting 1
  • 340538 11/17/10 18:48 efuchey Compton rates for 11-17-10
  • 340537 11/17/10 18:37 a-onl End_of_Run_8842
  • 340536 11/17/10 18:14 auto_snap Radiation and Ozone snapshot
  • 340535 11/17/10 18:14 auto_snap Miscellaneous snapshot
  • 340534 11/17/10 18:14 auto_snap Hall A Tools snapshot
  • 340533 11/17/10 18:14 auto_snap Compton Polarimeter snapshot
  • 340532 11/17/10 18:14 auto_snap Beam Position and Current snapshot
  • 340531 11/17/10 18:06 cshloka more info on compton plots
  • 340530 11/17/10 17:53 puckett increased J/T valve by two steps
  • 340529 11/17/10 17:23 miham Harp scan results (during beam recovery procedure)
  • 340528 11/17/10 17:18 miham Beam is recovered and we are taking data!
  • 340527 11/17/10 17:15 puckett moved target to loop 2, LH2 15 cm
  • 340526 11/17/10 17:15 a-onl Start_Run_8842,Run_type=Elastic Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.981,leftarm_theta=40.00,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=first run after beam back
  • 340525 11/17/10 17:06 a-onl End_of_Run_8841
  • 340524 11/17/10 17:03 a-onl End_of_Run_8840
  • 340523 11/17/10 17:02 a-onl Start_Run_8840,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DVCSDAQ2
  • 340522 11/17/10 16:58 a-onl End_of_Run_8839
  • 340521 11/17/10 16:57 a-onl Start_Run_8839,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test DVCSDAQ2
  • 340520 11/17/10 16:56 miham HRS-L Magnets are set.
  • 340519 11/17/10 16:53 miham BPMs are set!
  • 340518 11/17/10 16:33 miham Returning to Beam Permit
  • 340517 11/17/10 16:30 miham Cycling of Q2 and Q3 magnets
  • 340516 11/17/10 16:28 jroche, cshl Left HRS at 40 deg
  • 340515 11/17/10 16:08 A. Camsonne Small current correlation of Hall A vs Hall C
  • 340514 11/17/10 16:03 hsmith Left Dipole polarity change
  • 340513 11/17/10 15:52 adaq Beam Time Accounting
  • 340512 11/17/10 15:43 Ole Hansen adaq1/adaql2 disk mounts
  • 340511 11/17/10 15:43 Saha Shift Summary
  • 340510 11/17/10 15:32 a-onl Start_Run_8831,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Switching polarity, cosmics
  • 340509 11/17/10 15:32 a-onl Start_Run_8832,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Switching polarity, cosmics
  • 340508 11/17/10 15:32 a-onl End_of_Run_8830
  • 340507 11/17/10 15:32 a-onl End_of_Run_8831
  • 340506 11/17/10 15:32 a-onl End_of_Run_8832
  • 340505 11/17/10 15:28 cshloka, jro compton asymmetry plots
  • 340504 11/17/10 15:10 a-onl Start_Run_8830,Run_type=Cosmics,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Switching polarity, cosmics
  • 340503 11/17/10 15:07 hyde ABU / BANU
  • 340502 11/17/10 15:07 hyde HRS-L Polarity Switch
  • 340501 11/17/10 15:03 hyde Target to Empty
  • 340500 11/17/10 15:02 hyde HRS-L Collimator
  • 340499 11/17/10 14:53 a-onl End_of_Run_8829
  • 340498 11/17/10 14:41 a-onl Start_Run_8829,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Efficiency Kin II 4 pass LD2 LeftHRS ps3=ps4=1
  • 340497 11/17/10 14:40 hyde end of run 8827 ROCs
  • 340496 11/17/10 14:38 a-onl End_of_Run_8828
  • 340495 11/17/10 14:10 a-onl Start_Run_8828,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340494 11/17/10 14:09 a-onl End_of_Run_8825
  • 340493 11/17/10 14:08 n.muangma runcontrol, rocs 8826
  • 340492 11/17/10 13:31 a-onl Start_Run_8827,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340491 11/17/10 13:29 a-onl End_of_Run_8826
  • 340490 11/17/10 13:29 n.muangma 8826m runcontrol, rocs
  • 340489 11/17/10 13:18 N. Manugma Traget snapshot
  • 340488 11/17/10 12:30 a-onl Start_Run_8826,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340487 11/17/10 12:29 n.muangma 8825, runcontrol, rocs
  • 340486 11/17/10 12:10 hyde BPM, BCM strip chart
  • 340485 11/17/10 12:09 hyde DVCS DC Levels
  • 340484 11/17/10 11:54 A. Camsonne Flag for enabling adaql1 put to 1
  • 340483 11/17/10 11:32 a-onl Start_Run_8825,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340482 11/17/10 11:32 a-onl End_of_Run_8824
  • 340481 11/17/10 11:31 n.muangma 8824 runcontrol rocs
  • 340480 11/17/10 11:08 auto_snap Radiation and Ozone snapshot
  • 340479 11/17/10 11:08 auto_snap Miscellaneous snapshot
  • 340478 11/17/10 11:08 auto_snap Hall A Tools snapshot
  • 340477 11/17/10 11:08 auto_snap Compton Polarimeter snapshot
  • 340476 11/17/10 11:08 auto_snap Beam Position and Current snapshot
  • 340475 11/17/10 10:35 a-onl Start_Run_8824,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340474 11/17/10 10:34 a-onl End_of_Run_8823
  • 340473 11/17/10 10:17 n.muangma runcontrol, rocs, 8823
  • 340472 11/17/10 10:16 J. Zhang target pressure increasing
  • 340471 11/17/10 10:14 auto_snap Radiation and Ozone snapshot
  • 340470 11/17/10 10:14 auto_snap Miscellaneous snapshot
  • 340469 11/17/10 10:14 auto_snap Hall A Tools snapshot
  • 340468 11/17/10 10:14 auto_snap Compton Polarimeter snapshot
  • 340467 11/17/10 10:14 auto_snap Beam Position and Current snapshot
  • 340466 11/17/10 09:51 R. Michaels adaql2 disks practically full
  • 340465 11/17/10 09:23 a-onl Start_Run_8823,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340464 11/17/10 09:22 a-onl End_of_Run_8822
  • 340463 11/17/10 09:18 n.muangma runcontrol, rocs 9922
  • 340462 11/17/10 08:26 a-onl Start_Run_8822,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340461 11/17/10 08:25 a-onl End_of_Run_8821
  • 340460 11/17/10 08:02 vahe Owl Shift Summar
  • 340459 11/17/10 08:00 adaq Beam Time Accounting
  • 340458 11/17/10 07:19 a-onl Start_Run_8821,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340457 11/17/10 07:18 a-onl End_of_Run_8820
  • 340456 11/17/10 07:16 sutera rates run 8820
  • 340455 11/17/10 07:15 sutera roc windows run 8820
  • 340454 11/17/10 06:19 a-onl Start_Run_8820,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340453 11/17/10 06:18 a-onl End_of_Run_8819
  • 340452 11/17/10 06:13 sutera rates run 8819
  • 340451 11/17/10 06:12 sutera roc windows run 8819
  • 340450 11/17/10 05:19 a-onl Start_Run_8819,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340449 11/17/10 05:18 a-onl End_of_Run_8818
  • 340448 11/17/10 05:15 sutera rates run 8818
  • 340447 11/17/10 05:14 sutera roc windows run 8818
  • 340446 11/17/10 04:21 a-onl Start_Run_8818,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340445 11/17/10 04:20 a-onl End_of_Run_8817
  • 340444 11/17/10 03:27 a-onl Start_Run_8817,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340443 11/17/10 03:27 a-onl End_of_Run_8809
  • 340442 11/17/10 03:26 a-onl End_of_Run_8816
  • 340441 11/17/10 03:25 sutera rates run 8816
  • 340440 11/17/10 03:24 sutera roc window run 8816
  • 340439 11/17/10 03:08 auto_snap Radiation and Ozone snapshot
  • 340438 11/17/10 03:08 auto_snap Miscellaneous snapshot
  • 340437 11/17/10 03:08 auto_snap Hall A Tools snapshot
  • 340436 11/17/10 03:08 vahe Beam position drift starting 3:00 AM
  • 340435 11/17/10 03:08 auto_snap Compton Polarimeter snapshot
  • 340434 11/17/10 03:08 auto_snap Beam Position and Current snapshot
  • 340433 11/17/10 02:29 a-onl Start_Run_8816,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340432 11/17/10 02:28 a-onl End_of_Run_8815
  • 340431 11/17/10 02:27 sutera rates run 8815
  • 340430 11/17/10 02:26 sutera roc windows run 8815
  • 340429 11/17/10 02:14 auto_snap Radiation and Ozone snapshot
  • 340428 11/17/10 02:14 auto_snap Miscellaneous snapshot
  • 340427 11/17/10 02:13 auto_snap Hall A Tools snapshot
  • 340426 11/17/10 02:13 auto_snap Compton Polarimeter snapshot
  • 340425 11/17/10 02:13 auto_snap Beam Position and Current snapshot
  • 340424 11/17/10 01:28 a-onl Start_Run_8815,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340423 11/17/10 01:27 a-onl End_of_Run_8814
  • 340422 11/17/10 01:23 sutera rates run 8814
  • 340421 11/17/10 01:22 sutera roc windows run 8814
  • 340420 11/17/10 00:29 a-onl Start_Run_8814,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340419 11/17/10 00:28 a-onl End_of_Run_8813
  • 340418 11/17/10 00:24 sutera rates run 8813
  • 340417 11/17/10 00:23 sutera roc windows run 8813
  • 340416 11/17/10 00:03 miham Swing Shift Summary (11/16/10)
  • 340415 11/17/10 00:00 adaq Beam Time Accounting
  • Tuesday

  • 340414 11/16/10 23:59 A. Puckett target snapshot
  • 340413 11/16/10 23:47 S. Sirca ROC and scaler snapshots
  • 340412 11/16/10 23:32 miham Calorimeter ADC Readback Monitor Screenshot
  • 340411 11/16/10 23:29 a-onl Start_Run_8813,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340410 11/16/10 23:28 a-onl End_of_Run_8812
  • 340409 11/16/10 22:36 miham Q3 magnet trip (screenshots)
  • 340408 11/16/10 22:32 miham Q3 magnet trip
  • 340407 11/16/10 22:28 a-onl Start_Run_8812,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340406 11/16/10 22:17 A. Puckett target snapshot
  • 340405 11/16/10 22:08 a-onl End_of_Run_8811
  • 340404 11/16/10 21:58 miham Beam is back
  • 340403 11/16/10 21:57 a-onl Start_Run_8811,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340402 11/16/10 21:35 miham No beam!!!
  • 340401 11/16/10 21:32 a-onl End_of_Run_8810
  • 340400 11/16/10 21:22 miham End-of-run file for run #8809 is missing
  • 340399 11/16/10 20:46 a-onl Start_Run_8810,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340398 11/16/10 20:25 A. Puckett target stripcharts
  • 340397 11/16/10 19:49 a-onl Start_Run_8809,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340396 11/16/10 19:48 a-onl End_of_Run_8808
  • 340395 11/16/10 19:08 auto_snap Radiation and Ozone snapshot
  • 340394 11/16/10 19:08 auto_snap Miscellaneous snapshot
  • 340393 11/16/10 19:08 auto_snap Hall A Tools snapshot
  • 340392 11/16/10 19:08 auto_snap Compton Polarimeter snapshot
  • 340391 11/16/10 19:07 auto_snap Beam Position and Current snapshot
  • 340390 11/16/10 18:47 miham Beam is back
  • 340389 11/16/10 18:44 a-onl Start_Run_8808,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340388 11/16/10 18:22 miham Problem with a Gun
  • 340387 11/16/10 18:18 a-onl End_of_Run_8807
  • 340386 11/16/10 18:13 auto_snap Radiation and Ozone snapshot
  • 340385 11/16/10 18:13 auto_snap Miscellaneous snapshot
  • 340384 11/16/10 18:13 auto_snap Hall A Tools snapshot
  • 340383 11/16/10 18:13 auto_snap Compton Polarimeter snapshot
  • 340382 11/16/10 18:13 auto_snap Beam Position and Current snapshot
  • 340381 11/16/10 17:57 puckett target stripcharts and GUI snapshot
  • 340380 11/16/10 17:30 S. Sirca BPM histograms after prebuncher fix
  • 340379 11/16/10 17:18 miham Two peaks in BPM-y
  • 340378 11/16/10 17:15 a-onl Start_Run_8807,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340377 11/16/10 17:15 a-onl End_of_Run_8806
  • 340376 11/16/10 16:55 S. Sirca ROC snapshots
  • 340375 11/16/10 16:31 miham Calorimeter ADC Readback monitor (Screenshot)
  • 340374 11/16/10 16:26 miham HallA Tools (screenshot)
  • 340373 11/16/10 16:08 a-onl Start_Run_8806,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340372 11/16/10 16:07 a-onl End_of_Run_8805
  • 340371 11/16/10 16:05 adaq Beam Time Accounting
  • 340370 11/16/10 15:59 efuchey 11-16-10 day shift summary
  • 340369 11/16/10 15:59 A. Camsonne Prebuncher adjustement
  • 340368 11/16/10 15:45 Hisham Target snapshots
  • 340367 11/16/10 15:45 efuchey Hall A tools
  • 340366 11/16/10 15:43 efuchey Calo DC currents
  • 340365 11/16/10 15:40 efuchey Compton rates check retry.
  • 340364 11/16/10 15:37 A. Camsonne BPMs in Hall A
  • 340363 11/16/10 15:25 A. Camsonne Run 8805 looks good after 6 splits
  • 340362 11/16/10 15:12 A. Camsonne adaql2 run 8805
  • 340361 11/16/10 15:09 a-onl Start_Run_8805,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340360 11/16/10 14:54 a-onl End_of_Run_8803
  • 340359 11/16/10 14:53 efuchey Compton rates for today 11-16
  • 340358 11/16/10 14:50 A. Camsonne L1A rate on adaql2
  • 340357 11/16/10 14:39 a-onl Start_Run_8803,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2
  • 340356 11/16/10 14:37 A. Camsonne Deadtime with NoCheckDAQ2
  • 340355 11/16/10 14:36 A. Camsonne L1A rate on DVCSDAQ2
  • 340354 11/16/10 14:34 Hisham Target moved to LD2
  • 340353 11/16/10 14:29 a-onl Start_Run_8801,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DVCSDAQ2 with EB2 in config ( seems to work ... )
  • 340352 11/16/10 14:23 efuchey Beam is back
  • 340351 11/16/10 13:06 efuchey cryo flow came back
  • 340350 11/16/10 12:26 efuchey N2 at low level for Q3
  • 340349 11/16/10 11:48 Hisham Target: D2Pressures
  • 340348 11/16/10 11:42 a-onl End_of_Run_8800
  • 340347 11/16/10 11:41 a-onl Start_Run_8800,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340346 11/16/10 11:07 auto_snap Radiation and Ozone snapshot
  • 340345 11/16/10 11:07 auto_snap Miscellaneous snapshot
  • 340344 11/16/10 11:07 auto_snap Hall A Tools snapshot
  • 340343 11/16/10 11:07 auto_snap Compton Polarimeter snapshot
  • 340342 11/16/10 11:07 auto_snap Beam Position and Current snapshot
  • 340341 11/16/10 10:57 Fansler/Ewin Moeller/RQ2/N2
  • 340340 11/16/10 10:54 a-onl End_of_Run_8799
  • 340339 11/16/10 10:52 a-onl Start_Run_8799,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340338 11/16/10 10:48 a-onl End_of_Run_8798
  • 340337 11/16/10 10:47 a-onl Start_Run_8798,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340336 11/16/10 10:13 auto_snap Radiation and Ozone snapshot
  • 340335 11/16/10 10:13 auto_snap Miscellaneous snapshot
  • 340334 11/16/10 10:13 auto_snap Hall A Tools snapshot
  • 340333 11/16/10 10:13 auto_snap Compton Polarimeter snapshot
  • 340332 11/16/10 10:13 auto_snap Beam Position and Current snapshot
  • 340331 11/16/10 10:06 efuchey QWP scan for 11-16-10
  • 340330 11/16/10 09:58 a-onl End_of_Run_8796
  • 340329 11/16/10 09:56 R. Michaels further comments re: deadtime
  • 340328 11/16/10 09:52 R. Michaels switch to only adaql2 disks
  • 340327 11/16/10 09:34 a-onl Start_Run_8796,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340326 11/16/10 09:32 Bob/Alex/Dav DAQ deadtime studies
  • 340325 11/16/10 09:26 a-onl End_of_Run_8784
  • 340324 11/16/10 09:03 a-onl Start_Run_8794,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340323 11/16/10 09:01 a-onl End_of_Run_8793
  • 340322 11/16/10 08:58 J. Zhang JT Valve adjustment
  • 340321 11/16/10 08:58 a-onl Start_Run_8793,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340320 11/16/10 08:56 a-onl End_of_Run_8792
  • 340319 11/16/10 08:55 a-onl Start_Run_8791,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test -
  • 340318 11/16/10 08:55 a-onl Start_Run_8792,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test
  • 340317 11/16/10 08:54 a-onl End_of_Run_8791
  • 340316 11/16/10 08:51 Carlos Controlled access
  • 340315 11/16/10 08:47 Hisham Target moved to empty
  • 340314 11/16/10 08:44 a-onl End_of_Run_8790
  • 340313 11/16/10 08:36 a-onl Start_Run_8790,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test
  • 340312 11/16/10 08:35 a-onl End_of_Run_8789
  • 340311 11/16/10 08:27 a-onl Start_Run_8789,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test
  • 340310 11/16/10 08:26 a-onl End_of_Run_8788
  • 340309 11/16/10 08:19 a-onl Start_Run_8788,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test
  • 340308 11/16/10 08:18 a-onl End_of_Run_8786
  • 340307 11/16/10 08:15 a-onl End_of_Run_8786
  • 340306 11/16/10 08:15 a-onl Start_Run_8786,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=daq test
  • 340305 11/16/10 08:14 a-onl End_of_Run_8785
  • 340304 11/16/10 08:06 a-onl Start_Run_8785,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340302 11/16/10 08:04 sutera rates run 8784
  • 340303 11/16/10 08:03 adaq Beam Time Accounting
  • 340301 11/16/10 08:02 sutera roc windows run 8784
  • 340300 11/16/10 08:01 vahe Owl Shift Summary
  • 340299 11/16/10 07:45 Carlos Beam position drift starting at 7AM
  • 340298 11/16/10 07:06 a-onl Start_Run_8784,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340297 11/16/10 07:05 a-onl End_of_Run_8783
  • 340296 11/16/10 07:04 sutera rates run 8783
  • 340295 11/16/10 07:03 sutera roc windows run 8783
  • 340294 11/16/10 06:06 a-onl Start_Run_8783,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340293 11/16/10 06:05 a-onl End_of_Run_8782
  • 340292 11/16/10 06:04 sutera rates run 8782
  • 340291 11/16/10 06:03 sutera roc windows run 8782
  • 340290 11/16/10 05:05 a-onl Start_Run_8782,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340289 11/16/10 05:05 a-onl End_of_Run_8779
  • 340288 11/16/10 05:04 a-onl End_of_Run_8781
  • 340287 11/16/10 05:02 sutera rates run 8781
  • 340286 11/16/10 05:01 sutera roc windows run 8781
  • 340285 11/16/10 04:04 a-onl Start_Run_8781,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340284 11/16/10 04:00 a-onl End_of_Run_8780
  • 340283 11/16/10 03:59 sutera rates run 8780
  • 340282 11/16/10 03:58 sutera roc windows run 8780
  • 340281 11/16/10 03:07 auto_snap Radiation and Ozone snapshot
  • 340280 11/16/10 03:07 auto_snap Miscellaneous snapshot
  • 340279 11/16/10 03:07 auto_snap Hall A Tools snapshot
  • 340278 11/16/10 03:07 auto_snap Compton Polarimeter snapshot
  • 340277 11/16/10 03:07 auto_snap Beam Position and Current snapshot
  • 340276 11/16/10 03:00 a-onl Start_Run_8780,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340275 11/16/10 02:59 sutera rates run 8779
  • 340274 11/16/10 02:58 sutera roc windows run 8779
  • 340273 11/16/10 02:13 auto_snap Radiation and Ozone snapshot
  • 340272 11/16/10 02:13 auto_snap Miscellaneous snapshot
  • 340271 11/16/10 02:13 auto_snap Hall A Tools snapshot
  • 340270 11/16/10 02:13 auto_snap Compton Polarimeter snapshot
  • 340269 11/16/10 02:12 auto_snap Beam Position and Current snapshot
  • 340268 11/16/10 02:00 a-onl Start_Run_8779,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340267 11/16/10 01:59 a-onl End_of_Run_8778
  • 340266 11/16/10 01:58 sutera rates run 8778
  • 340265 11/16/10 01:57 sutera roc windows run 8778
  • 340264 11/16/10 01:00 a-onl Start_Run_8778,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340263 11/16/10 00:59 a-onl End_of_Run_8777
  • 340262 11/16/10 00:58 sutera rates run 8777
  • 340261 11/16/10 00:57 sutera roc windows run 8777
  • 340260 11/16/10 00:32 Min Huang Target snapshot
  • 340259 11/16/10 00:04 A. Puckett 11/15 Swing Shift Summary
  • 340258 11/16/10 00:01 adaq Beam Time Accounting
  • Monday

  • 340257 11/15/10 23:57 a-onl Start_Run_8777,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340256 11/15/10 23:57 a-onl End_of_Run_8776
  • 340255 11/15/10 23:56 pzhu roc,rc,scaler windows for run 8776
  • 340254 11/15/10 23:32 Hisham ADC helicity display
  • 340253 11/15/10 23:06 a-onl Start_Run_8776,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340252 11/15/10 23:02 a-onl End_of_Run_8775
  • 340251 11/15/10 23:01 pzhu roc,rc,scaler windows for run 8775
  • 340250 11/15/10 23:01 pzhu roc,rc,scaler windows for run 8775
  • 340249 11/15/10 22:08 Min Huang target status
  • 340248 11/15/10 22:01 a-onl Start_Run_8775,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340247 11/15/10 22:00 a-onl End_of_Run_8774
  • 340246 11/15/10 21:50 pzhu roc,rc,scaler windows for run 8774
  • 340245 11/15/10 20:58 a-onl Start_Run_8774,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340244 11/15/10 20:58 a-onl End_of_Run_8773
  • 340243 11/15/10 20:57 pzhu roc,rc,scaler window for run 8773
  • 340242 11/15/10 20:07 Min Huang Target status
  • 340241 11/15/10 19:52 a-onl Start_Run_8773,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340240 11/15/10 19:51 pzhu roc,rc,scaler screenshot for run 8772
  • 340239 11/15/10 19:51 a-onl End_of_Run_8772
  • 340238 11/15/10 19:33 Carlos Beam positions after beam trip
  • 340237 11/15/10 18:54 Carlos Configuration downloaded
  • 340236 11/15/10 18:51 a-onl Start_Run_8772,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340235 11/15/10 18:50 a-onl End_of_Run_8771
  • 340234 11/15/10 18:49 pzhu roc,rc,scaler screenshot for run 8771
  • 340233 11/15/10 18:17 A. Camsonne Modified ts_scaler.crl
  • 340232 11/15/10 18:12 auto_snap Radiation and Ozone snapshot
  • 340231 11/15/10 18:12 auto_snap Miscellaneous snapshot
  • 340230 11/15/10 18:12 auto_snap Hall A Tools snapshot
  • 340229 11/15/10 18:12 auto_snap Compton Polarimeter snapshot
  • 340228 11/15/10 18:12 auto_snap Beam Position and Current snapshot
  • 340227 11/15/10 17:51 a-onl Start_Run_8771,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340226 11/15/10 17:50 a-onl End_of_Run_8770
  • 340225 11/15/10 17:50 pzhu roc,rc,scaler screen for run 8770
  • 340224 11/15/10 17:49 pzhu roc,rc,scaler screen for run 8770
  • 340223 11/15/10 16:51 a-onl Start_Run_8770,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340222 11/15/10 16:50 a-onl End_of_Run_8769
  • 340221 11/15/10 16:50 pzhu roc,rc,scaler screen for run 8769
  • 340220 11/15/10 16:14 Min Huang Target status
  • 340218 11/15/10 16:01 hyde/Saha shift summary
  • 340219 11/15/10 16:00 adaq Beam Time Accounting
  • 340217 11/15/10 15:50 a-onl Start_Run_8769,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340216 11/15/10 15:46 a-onl End_of_Run_8768
  • 340215 11/15/10 15:20 a-onl Start_Run_8768,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340214 11/15/10 15:19 a-onl End_of_Run_8767
  • 340213 11/15/10 15:17 sutera rates run 8767
  • 340212 11/15/10 15:16 sutera roc windows run 8767
  • 340211 11/15/10 14:20 a-onl Start_Run_8767,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340210 11/15/10 14:14 a-onl End_of_Run_8766
  • 340209 11/15/10 13:45 a-onl Start_Run_8766,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 trigger efficiencies
  • 340208 11/15/10 13:37 a-onl End_of_Run_8765
  • 340207 11/15/10 13:35 sutera rates run 8765
  • 340206 11/15/10 13:34 sutera roc windows run 8765
  • 340205 11/15/10 12:38 a-onl Start_Run_8765,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340204 11/15/10 12:36 a-onl End_of_Run_8764
  • 340203 11/15/10 12:35 sutera rates run 8764
  • 340202 11/15/10 12:34 sutera roc windows run 8764
  • 340201 11/15/10 12:17 hyde gas shed alarm reset
  • 340200 11/15/10 12:07 hyde gas shed alarm
  • 340199 11/15/10 11:37 a-onl Start_Run_8764,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340198 11/15/10 11:36 a-onl End_of_Run_8763
  • 340197 11/15/10 11:07 auto_snap Radiation and Ozone snapshot
  • 340196 11/15/10 11:07 auto_snap Miscellaneous snapshot
  • 340195 11/15/10 11:07 auto_snap Hall A Tools snapshot
  • 340194 11/15/10 11:07 auto_snap Compton Polarimeter snapshot
  • 340193 11/15/10 11:07 auto_snap Beam Position and Current snapshot
  • 340192 11/15/10 11:06 hyde BPM Time history
  • 340191 11/15/10 10:51 jmammei scattering chamber vacuum high
  • 340190 11/15/10 10:42 a-onl Start_Run_8763,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340189 11/15/10 10:41 a-onl End_of_Run_8762
  • 340188 11/15/10 10:39 sutera rates run 8762
  • 340187 11/15/10 10:39 sutera roc windows run 8762
  • 340186 11/15/10 10:28 hyde DVCS DC levels
  • 340185 11/15/10 10:27 hyde Hall A Tools
  • 340184 11/15/10 10:22 hyde beam position
  • 340183 11/15/10 10:21 hyde beam trip run 8762
  • 340182 11/15/10 10:15 J. Zhang Update 'lastrun' and 'spot++' to adopt disk in adaql1
  • 340181 11/15/10 10:12 auto_snap Radiation and Ozone snapshot
  • 340180 11/15/10 10:12 auto_snap Miscellaneous snapshot
  • 340179 11/15/10 10:12 auto_snap Hall A Tools snapshot
  • 340178 11/15/10 10:12 auto_snap Compton Polarimeter snapshot
  • 340177 11/15/10 10:12 auto_snap Beam Position and Current snapshot
  • 340176 11/15/10 09:46 jmammei 4K cryo flow crazy
  • 340175 11/15/10 09:45 hyde Y position study
  • 340174 11/15/10 09:42 a-onl Start_Run_8762,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340173 11/15/10 09:41 a-onl End_of_Run_8761
  • 340172 11/15/10 09:34 a-onl Start_Run_8761,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340171 11/15/10 09:33 a-onl End_of_Run_8760
  • 340170 11/15/10 09:29 a-onl Start_Run_8760,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340169 11/15/10 09:28 a-onl End_of_Run_8759
  • 340168 11/15/10 09:26 a-onl Start_Run_8759,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340167 11/15/10 09:26 a-onl End_of_Run_8758
  • 340166 11/15/10 09:20 a-onl Start_Run_8758,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340165 11/15/10 09:19 a-onl End_of_Run_8757
  • 340164 11/15/10 09:15 a-onl Start_Run_8757,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340163 11/15/10 09:14 a-onl End_of_Run_8756
  • 340162 11/15/10 09:10 a-onl Start_Run_8756,Run_type=Spot++,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2 Testing
  • 340161 11/15/10 09:06 a-onl End_of_Run_8755
  • 340160 11/15/10 08:45 efuchey QWP scan results for 11-14
  • 340159 11/15/10 08:42 efuchey QWP scan results for 11/13
  • 340158 11/15/10 08:29 a-onl Start_Run_8755,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340157 11/15/10 08:28 a-onl End_of_Run_8754
  • 340156 11/15/10 08:28 sutera rates run 8754
  • 340154 11/15/10 08:27 sutera Roc windows run 8754
  • 340153 11/15/10 08:18 jmammei closed Update Manager
  • 340152 11/15/10 08:14 R. Michaels re: how to turn off pedsup
  • 340151 11/15/10 08:12 jmammei target alarms
  • 340150 11/15/10 08:08 adaq Beam Time Accounting
  • 340149 11/15/10 08:07 jroche Owl shift summary
  • 340148 11/15/10 07:43 jroche, hish Helicity ADC pedestal fixed.
  • 340147 11/15/10 07:29 a-onl Start_Run_8754,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340146 11/15/10 07:28 a-onl End_of_Run_8753
  • 340145 11/15/10 07:26 Fansler Q3 N2 level
  • 340144 11/15/10 07:19 a-onl Start_Run_8753,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340143 11/15/10 07:18 a-onl End_of_Run_8752
  • 340142 11/15/10 07:17 Hisham RC, scalers, ROCs: run # 8752
  • 340141 11/15/10 07:16 Hisham ADC helicity and pedestal subtraction
  • 340140 11/15/10 06:17 a-onl Start_Run_8752,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340139 11/15/10 06:16 a-onl End_of_Run_8751
  • 340138 11/15/10 06:09 Hisham RC, scalers, ROCs: run#8751
  • 340137 11/15/10 05:50 Hisham Target and charts snapshots
  • 340136 11/15/10 05:22 jroche Updated epics target readout list
  • 340135 11/15/10 05:15 a-onl Start_Run_8751,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340134 11/15/10 05:14 a-onl End_of_Run_8750
  • 340133 11/15/10 05:00 Hisham RC, scalers, ROCs: run # 8750
  • 340132 11/15/10 04:29 jroche Calorimeter DC currents
  • 340131 11/15/10 04:27 jroche [TECHS]: Left Q3 low in nitrogen
  • 340130 11/15/10 04:13 a-onl Start_Run_8750,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340129 11/15/10 04:13 a-onl End_of_Run_8749
  • 340128 11/15/10 04:06 Hisham RC, scalers, ROCs : run# 8749
  • 340127 11/15/10 03:30 jroche beam position at the time of the vacuum excursion
  • 340126 11/15/10 03:13 a-onl Start_Run_8749,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340125 11/15/10 03:10 a-onl End_of_Run_8748
  • 340124 11/15/10 03:08 a-onl Start_Run_8748,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340123 11/15/10 03:06 auto_snap Radiation and Ozone snapshot
  • 340122 11/15/10 03:06 auto_snap Miscellaneous snapshot
  • 340121 11/15/10 03:06 a-onl End_of_Run_8747
  • 340120 11/15/10 03:06 auto_snap Hall A Tools snapshot
  • 340119 11/15/10 03:06 auto_snap Compton Polarimeter snapshot
  • 340118 11/15/10 03:06 auto_snap Beam Position and Current snapshot
  • 340117 11/15/10 03:05 a-onl Start_Run_8747,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text= KinII 4 pass LH2, only ps9=1
  • 340116 11/15/10 03:03 a-onl End_of_Run_8746
  • 340115 11/15/10 03:00 Hisham RC, scalers, ROCs : run#
  • 340114 11/15/10 02:49 Hisham Target status: valve trip
  • 340113 11/15/10 02:12 auto_snap Radiation and Ozone snapshot
  • 340112 11/15/10 02:12 auto_snap Miscellaneous snapshot
  • 340111 11/15/10 02:12 auto_snap Hall A Tools snapshot
  • 340110 11/15/10 02:12 auto_snap Compton Polarimeter snapshot
  • 340109 11/15/10 02:12 auto_snap Beam Position and Current snapshot
  • 340108 11/15/10 02:04 a-onl Start_Run_8746,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340107 11/15/10 02:03 a-onl End_of_Run_8745
  • 340106 11/15/10 01:52 Hisham RC, scalers, ROCs: run # 8745
  • 340105 11/15/10 00:56 a-onl Start_Run_8745,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340104 11/15/10 00:56 a-onl End_of_Run_8744
  • 340103 11/15/10 00:53 Hisham RunControl, scalers, ROCs : run 8744
  • 340101 11/15/10 00:02 Pawel Swing shift summary
  • 340102 11/15/10 00:01 adaq Beam Time Accounting
  • Sunday

  • 340100 11/14/10 23:56 a-onl Start_Run_8744,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340099 11/14/10 23:55 a-onl End_of_Run_8743
  • 340098 11/14/10 23:54 sutera rates run 8743
  • 340097 11/14/10 23:53 sutera roc windows run 8743
  • 340096 11/14/10 22:59 a-onl Start_Run_8743,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340095 11/14/10 22:59 a-onl End_of_Run_8742
  • 340094 11/14/10 22:58 sutera rates run 8742
  • 340093 11/14/10 22:57 sutera roc windows run 8742
  • 340092 11/14/10 22:48 Pawel Calorimeter HV
  • 340091 11/14/10 21:58 a-onl Start_Run_8742,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340090 11/14/10 21:58 a-onl End_of_Run_8741
  • 340089 11/14/10 21:57 sutera rates run 8741
  • 340088 11/14/10 21:55 sutera roc windows run 8741
  • 340087 11/14/10 21:01 a-onl Start_Run_8741,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340086 11/14/10 21:00 a-onl End_of_Run_8740
  • 340085 11/14/10 20:59 sutera rates run 8740
  • 340084 11/14/10 20:58 sutera roc windows run 8740
  • 340083 11/14/10 20:01 a-onl Start_Run_8740,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340082 11/14/10 20:00 a-onl End_of_Run_8739
  • 340081 11/14/10 19:59 sutera rates run 8739
  • 340080 11/14/10 19:57 sutera roc windows run 8739
  • 340079 11/14/10 19:06 auto_snap Radiation and Ozone snapshot
  • 340078 11/14/10 19:06 auto_snap Miscellaneous snapshot
  • 340077 11/14/10 19:06 auto_snap Hall A Tools snapshot
  • 340076 11/14/10 19:06 auto_snap Compton Polarimeter snapshot
  • 340075 11/14/10 19:06 auto_snap Beam Position and Current snapshot
  • 340074 11/14/10 19:01 Pawel RF problems resolved - beam is back
  • 340073 11/14/10 19:00 a-onl Start_Run_8739,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340072 11/14/10 19:00 a-onl End_of_Run_8726
  • 340071 11/14/10 18:18 Pawel MCC has RF problems
  • 340070 11/14/10 18:17 a-onl End_of_Run_8738
  • 340069 11/14/10 18:11 auto_snap Radiation and Ozone snapshot
  • 340068 11/14/10 18:11 auto_snap Miscellaneous snapshot
  • 340067 11/14/10 18:11 auto_snap Hall A Tools snapshot
  • 340066 11/14/10 18:11 auto_snap Compton Polarimeter snapshot
  • 340065 11/14/10 18:11 auto_snap Beam Position and Current snapshot
  • 340064 11/14/10 17:51 a-onl Start_Run_8738,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340063 11/14/10 17:50 a-onl End_of_Run_8737
  • 340062 11/14/10 17:50 sutera rates run 8737
  • 340061 11/14/10 17:49 sutera roc windows run 8737
  • 340060 11/14/10 16:51 a-onl Start_Run_8737,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340059 11/14/10 16:47 a-onl End_of_Run_8736
  • 340058 11/14/10 16:27 adaq Beam Time Accounting
  • 340057 11/14/10 16:17 a-onl Start_Run_8736,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2, trigger efficiency
  • 340056 11/14/10 16:13 a-onl End_of_Run_8735
  • 340055 11/14/10 16:13 sutera rates run 8735
  • 340054 11/14/10 16:12 sutera roc windows run 8735
  • 340053 11/14/10 16:10 sutera roc windows run 8735
  • 340052 11/14/10 16:08 Boeglin Shift Summary
  • 340051 11/14/10 15:22 a-onl Start_Run_8735,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340050 11/14/10 15:21 a-onl End_of_Run_8734
  • 340049 11/14/10 15:21 efuchey rates run 8734
  • 340048 11/14/10 15:20 efuchey ROC windows run 8734
  • 340046 11/14/10 14:16 efuchey Target switch
  • 340045 11/14/10 14:14 a-onl Start_Run_8734,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LH2
  • 340044 11/14/10 14:07 J. Zhang Switch to LH2 target and 3uA beam
  • 340043 11/14/10 14:06 a-onl End_of_Run_8733
  • 340042 11/14/10 14:06 efuchey Rates run 8733
  • 340041 11/14/10 14:05 efuchey ROC windows 8733
  • 340040 11/14/10 13:28 Boeglin HallA Tools Screen
  • 340039 11/14/10 13:27 Boeglin Q3 LN2 level
  • 340038 11/14/10 13:09 a-onl Start_Run_8733,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340037 11/14/10 13:08 a-onl End_of_Run_8732
  • 340036 11/14/10 13:08 efuchey Rates run 8732
  • 340035 11/14/10 13:06 efuchey ROC windows run 8732
  • 340034 11/14/10 12:16 efuchey edit for entry 340030
  • 340033 11/14/10 12:11 a-onl Start_Run_8732,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340032 11/14/10 12:10 a-onl End_of_Run_8731
  • 340031 11/14/10 12:10 efuchey Rates run 8731
  • 340030 11/14/10 12:09 efuchey ROC windows, run 8732
  • 340029 11/14/10 12:05 Boeglin Q3 LN2 fine again
  • 340028 11/14/10 12:01 Boeglin CALO DC levels
  • 340027 11/14/10 11:57 Boeglin Q3 low in N2 ?!
  • 340026 11/14/10 11:49 efuchey target charts
  • 340025 11/14/10 11:29 efuchey Target GUI
  • 340024 11/14/10 11:07 a-onl Start_Run_8731,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340023 11/14/10 11:06 a-onl End_of_Run_8730
  • 340022 11/14/10 11:06 auto_snap Radiation and Ozone snapshot
  • 340021 11/14/10 11:06 auto_snap Miscellaneous snapshot
  • 340019 11/14/10 11:06 auto_snap Hall A Tools snapshot
  • 340018 11/14/10 11:06 auto_snap Compton Polarimeter snapshot
  • 340017 11/14/10 11:06 auto_snap Beam Position and Current snapshot
  • 340016 11/14/10 11:05 efuchey ROC windows, run 8730
  • 340015 11/14/10 10:11 auto_snap Radiation and Ozone snapshot
  • 340014 11/14/10 10:11 auto_snap Miscellaneous snapshot
  • 340013 11/14/10 10:11 auto_snap Hall A Tools snapshot
  • 340012 11/14/10 10:11 auto_snap Compton Polarimeter snapshot
  • 340011 11/14/10 10:11 auto_snap Beam Position and Current snapshot
  • 340010 11/14/10 10:05 a-onl Start_Run_8730,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340009 11/14/10 10:05 a-onl End_of_Run_8729
  • 340008 11/14/10 10:04 efuchey Rates for run 8729
  • 340007 11/14/10 10:04 efuchey ROC windows run 8729
  • 340006 11/14/10 09:08 a-onl Start_Run_8729,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340005 11/14/10 09:07 a-onl End_of_Run_8728
  • 340004 11/14/10 09:07 efuchey rates for run 8728
  • 340003 11/14/10 09:06 efuchey ROC windows for run 8728
  • 340002 11/14/10 08:13 a-onl Start_Run_8728,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 340001 11/14/10 08:13 a-onl End_of_Run_8727
  • 340000 11/14/10 08:12 efuchey End of run 8727 rates
  • 339999 11/14/10 08:11 efuchey ROC windows
  • 339998 11/14/10 08:02 Hisham Owl shift summary
  • 339997 11/14/10 08:00 adaq Beam Time Accounting
  • 339996 11/14/10 07:14 a-onl Start_Run_8727,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339995 11/14/10 07:12 a-onl End_of_Run_8726
  • 339994 11/14/10 07:10 sutera rates run 8726
  • 339993 11/14/10 07:09 sutera roc windows run 8726
  • 339992 11/14/10 06:11 a-onl Start_Run_8726,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339991 11/14/10 06:11 a-onl End_of_Run_8725
  • 339990 11/14/10 06:10 sutera rates run 8725
  • 339989 11/14/10 06:09 sutera roc windows run 8725
  • 339988 11/14/10 06:07 HIsham Target snapshot
  • 339987 11/14/10 05:11 a-onl Start_Run_8725,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=2,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339986 11/14/10 05:09 a-onl End_of_Run_8724
  • 339985 11/14/10 05:08 sutera rates run 8724
  • 339984 11/14/10 05:07 sutera roc windows run 8724
  • 339983 11/14/10 04:11 a-onl Start_Run_8724,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339982 11/14/10 04:10 a-onl End_of_Run_8723
  • 339981 11/14/10 04:09 HIsham Calo anode current
  • 339980 11/14/10 04:09 sutera rates run 8723
  • 339979 11/14/10 04:08 sutera roc windows run 8723
  • 339978 11/14/10 04:00 Hisham Compton Coda crashed
  • 339977 11/14/10 03:11 a-onl Start_Run_8723,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339976 11/14/10 03:10 a-onl End_of_Run_8722
  • 339975 11/14/10 03:09 sutera rates run 8722
  • 339974 11/14/10 03:08 sutera roc windows run 8722
  • 339973 11/14/10 03:06 auto_snap Radiation and Ozone snapshot
  • 339972 11/14/10 03:06 auto_snap Miscellaneous snapshot
  • 339971 11/14/10 03:05 auto_snap Hall A Tools snapshot
  • 339970 11/14/10 03:05 auto_snap Compton Polarimeter snapshot
  • 339969 11/14/10 03:05 auto_snap Beam Position and Current snapshot
  • 339968 11/14/10 02:11 a-onl Start_Run_8722,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339967 11/14/10 02:11 auto_snap Radiation and Ozone snapshot
  • 339966 11/14/10 02:11 auto_snap Miscellaneous snapshot
  • 339965 11/14/10 02:11 auto_snap Hall A Tools snapshot
  • 339964 11/14/10 02:11 auto_snap Compton Polarimeter snapshot
  • 339963 11/14/10 02:11 auto_snap Beam Position and Current snapshot
  • 339962 11/14/10 02:10 a-onl End_of_Run_8721
  • 339961 11/14/10 02:10 a-onl Start_Run_8715,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ROC14Only T5
  • 339960 11/14/10 02:09 sutera rates run 8721
  • 339959 11/14/10 02:07 sutera roc windows run 8721
  • 339958 11/14/10 01:10 a-onl Start_Run_8721,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339957 11/14/10 01:09 a-onl End_of_Run_8720
  • 339956 11/14/10 01:08 sutera rates run 8720
  • 339955 11/14/10 01:07 sutera roc windows run 8720
  • 339954 11/14/10 00:07 a-onl Start_Run_8720,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339952 11/14/10 00:00 S. Golge Swing Shift Summary
  • 339953 11/14/10 00:00 adaq Beam Time Accounting
  • Saturday

  • 339951 11/13/10 23:54 cshloka HV turned off
  • 339950 11/13/10 23:29 a-onl End_of_Run_8719
  • 339949 11/13/10 23:29 chench screenshot of Roc14, Roc17 and Roc18(run8719)
  • 339948 11/13/10 23:27 chench screenshot of DAQ window and scalers(run8719)
  • 339947 11/13/10 22:39 cshloka target screens and status
  • 339946 11/13/10 22:30 a-onl Start_Run_8719,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339945 11/13/10 22:29 a-onl End_of_Run_8718
  • 339944 11/13/10 22:27 chench screenshot of Roc14, Roc17 and Roc18(run8718)
  • 339943 11/13/10 22:27 chench screenshot of DAQ window and scalers (run8718)
  • 339942 11/13/10 21:29 a-onl Start_Run_8718,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339941 11/13/10 21:28 a-onl End_of_Run_8717
  • 339940 11/13/10 21:26 chench screenshot of Roc14, Roc17 and Roc18(run8717)
  • 339939 11/13/10 21:25 chench screenshot of DAQ window and scalers (run8717)
  • 339938 11/13/10 20:28 a-onl Start_Run_8717,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339937 11/13/10 20:27 a-onl End_of_Run_8716
  • 339936 11/13/10 20:23 chench screenshot of Roc14, Roc17 and Roc18(run8716)
  • 339935 11/13/10 20:22 chench screenshot of DAQ window and scalers (run8716)
  • 339934 11/13/10 19:28 a-onl Start_Run_8716,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production KinII 4 pass LD2
  • 339933 11/13/10 19:26 a-onl End_of_Run_8715
  • 339932 11/13/10 19:23 a-onl Start_Run_8710,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test left HRS with T5
  • 339931 11/13/10 19:22 a-onl Start_Run_8713,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ROC1417Buff T5
  • 339930 11/13/10 19:21 a-onl Start_Run_8712,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test leftHRSBuff with T5
  • 339929 11/13/10 19:21 a-onl End_of_Run_8712
  • 339928 11/13/10 19:19 a-onl End_of_Run_8711
  • 339927 11/13/10 19:18 a-onl Start_Run_8711,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test left HRS with T5 and T1
  • 339926 11/13/10 19:17 a-onl End_of_Run_8710
  • 339925 11/13/10 19:15 a-onl End_of_Run_8709
  • 339924 11/13/10 19:05 auto_snap Radiation and Ozone snapshot
  • 339923 11/13/10 19:05 auto_snap Miscellaneous snapshot
  • 339922 11/13/10 19:05 auto_snap Hall A Tools snapshot
  • 339921 11/13/10 19:05 auto_snap Compton Polarimeter snapshot
  • 339920 11/13/10 19:05 auto_snap Beam Position and Current snapshot
  • 339919 11/13/10 18:51 A. Camsonne High dead time on deuterium
  • 339918 11/13/10 18:38 chench correction: the comment of the run8707 and 8708 is wrong
  • 339917 11/13/10 18:34 a-onl Start_Run_8709,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin II 4 pass LD2 target
  • 339916 11/13/10 18:30 cshloka target screens
  • 339915 11/13/10 18:28 cshloka Target moved to Loop 1, 15 cm (LD2)
  • 339914 11/13/10 18:24 S. Golge ALARM Detector Flammable Gas sensor alarm
  • 339913 11/13/10 18:23 a-onl End_of_Run_8708
  • 339912 11/13/10 18:11 auto_snap Radiation and Ozone snapshot
  • 339911 11/13/10 18:11 auto_snap Miscellaneous snapshot
  • 339910 11/13/10 18:10 auto_snap Hall A Tools snapshot
  • 339909 11/13/10 18:10 auto_snap Compton Polarimeter snapshot
  • 339908 11/13/10 18:10 auto_snap Beam Position and Current snapshot
  • 339907 11/13/10 17:49 a-onl Start_Run_8708,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LeftHRS efficiency run
  • 339906 11/13/10 17:49 a-onl End_of_Run_8701
  • 339905 11/13/10 17:48 a-onl End_of_Run_8707
  • 339904 11/13/10 17:48 chench Screenshot of Roc14, Roc17 and Roc18(run8707)
  • 339903 11/13/10 17:46 chench Screenshot of DAQ window and scalers (run8707)
  • 339902 11/13/10 16:52 chench change the prescale and configuration for run8707
  • 339901 11/13/10 16:49 a-onl Start_Run_8707,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LeftHRS efficiency run
  • 339900 11/13/10 16:45 a-onl End_of_Run_8706
  • 339899 11/13/10 16:30 A. Camsonne Test runs
  • 339898 11/13/10 16:16 A. Camsonne Rates LeftHRS config and L1A rate
  • 339897 11/13/10 16:13 a-onl Start_Run_8706,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=LeftHRS efficiency run
  • 339896 11/13/10 16:09 Boeglin end of run for 8702 probably belongs to 8701
  • 339895 11/13/10 16:08 A. Camsonne L1A rates run 8704 DVCSDAQ2 config
  • 339894 11/13/10 16:08 Boeglin Shift Summary
  • 339893 11/13/10 16:03 a-onl End_of_Run_8703
  • 339891 11/13/10 16:01 a-onl Start_Run_8703,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - Test noROC4 config
  • 339892 11/13/10 16:01 adaq Beam Time Accounting
  • 339890 11/13/10 16:00 a-onl End_of_Run_8702
  • 339889 11/13/10 16:00 a-onl Start_Run_8702,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - Test noROC3 config
  • 339888 11/13/10 15:52 jroche Helicity problem "history"
  • 339887 11/13/10 15:28 A. Camsonne Added roc4 flag to ts_scaler.crl
  • 339886 11/13/10 15:26 A. Camsonne Follow up warning message
  • 339885 11/13/10 15:01 A. Camsonne Removed timers from data stream for run 8701 for ROC3 and ROC4
  • 339884 11/13/10 15:00 a-onl Start_Run_8701,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339883 11/13/10 15:00 a-onl End_of_Run_8700
  • 339882 11/13/10 14:55 A. Camsonne Warning message ROC4 and ROC3
  • 339881 11/13/10 14:39 a-onl Start_Run_8700,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339880 11/13/10 14:37 a-onl End_of_Run_8699
  • 339879 11/13/10 14:33 a-onl Start_Run_8699,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339878 11/13/10 14:32 a-onl End_of_Run_8698
  • 339877 11/13/10 14:31 sutera rates run 8698
  • 339876 11/13/10 14:30 sutera roc windows run 8698
  • 339875 11/13/10 13:33 a-onl Start_Run_8698,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339874 11/13/10 13:31 a-onl End_of_Run_8697
  • 339873 11/13/10 13:31 sutera rates run 8697
  • 339872 11/13/10 13:29 sutera roc windows run 8697
  • 339871 11/13/10 12:29 a-onl Start_Run_8697,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339870 11/13/10 12:27 a-onl End_of_Run_8696
  • 339869 11/13/10 12:23 Boeglin New information in data stream
  • 339868 11/13/10 12:23 a-onl Start_Run_8696,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC - Check DAQ
  • 339867 11/13/10 12:22 A. Camsonne Added timer libraries to ROC3, ROC4 and TS11 boot scripts and rebooted ROCs
  • 339866 11/13/10 12:21 a-onl End_of_Run_8695
  • 339865 11/13/10 12:04 a-onl End_of_Run_8693
  • 339864 11/13/10 12:01 sutera rates run 8693
  • 339863 11/13/10 12:00 sutera roc windows run 8693
  • 339862 11/13/10 11:44 Boeglin spot++ result
  • 339861 11/13/10 11:42 A. Camsonne Added timers for dead time study to ROC3, ROC4 and TS11
  • 339860 11/13/10 11:39 Boeglin CALO currents
  • 339858 11/13/10 11:05 a-onl Start_Run_8693,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339859 11/13/10 11:05 auto_snap Radiation and Ozone snapshot
  • 339857 11/13/10 11:05 auto_snap Miscellaneous snapshot
  • 339856 11/13/10 11:05 auto_snap Hall A Tools snapshot
  • 339855 11/13/10 11:05 auto_snap Compton Polarimeter snapshot
  • 339854 11/13/10 11:05 auto_snap Beam Position and Current snapshot
  • 339853 11/13/10 11:04 a-onl End_of_Run_8692
  • 339852 11/13/10 11:04 sutera rates run 8692
  • 339851 11/13/10 11:02 sutera roc windows run 8692
  • 339850 11/13/10 10:56 efuchey target charts
  • 339849 11/13/10 10:49 efuchey Taget screeshots
  • 339848 11/13/10 10:10 auto_snap Radiation and Ozone snapshot
  • 339847 11/13/10 10:10 auto_snap Miscellaneous snapshot
  • 339846 11/13/10 10:10 auto_snap Hall A Tools snapshot
  • 339845 11/13/10 10:10 auto_snap Compton Polarimeter snapshot
  • 339844 11/13/10 10:10 auto_snap Beam Position and Current snapshot
  • 339843 11/13/10 10:03 a-onl Start_Run_8692,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339842 11/13/10 10:02 a-onl End_of_Run_8691
  • 339841 11/13/10 10:00 sutera rates run 8691
  • 339840 11/13/10 09:59 sutera roc windows run 8691
  • 339839 11/13/10 09:03 a-onl Start_Run_8691,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339838 11/13/10 09:02 a-onl End_of_Run_8690
  • 339837 11/13/10 09:00 sutera Rates run 8690
  • 339836 11/13/10 08:59 sutera Roc windows run 8690
  • 339835 11/13/10 08:03 Hisham Owl shift summary
  • 339834 11/13/10 08:03 a-onl Start_Run_8690,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339833 11/13/10 08:01 a-onl End_of_Run_8689
  • 339832 11/13/10 07:59 adaq Beam Time Accounting
  • 339831 11/13/10 07:32 a-onl Start_Run_8689,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - Trigger Efficiencies ps3=1 ps4=1 else=0
  • 339830 11/13/10 07:21 N. Muangma RunControl, scalers, rocs: Run#8687
  • 339829 11/13/10 06:24 a-onl Start_Run_8687,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339828 11/13/10 06:23 a-onl End_of_Run_8686
  • 339827 11/13/10 06:23 N Muangma RunControl. scaler, rocs for Run 8686
  • 339826 11/13/10 05:42 meziani Target screen
  • 339825 11/13/10 05:20 a-onl Start_Run_8686,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339824 11/13/10 05:19 a-onl End_of_Run_8685
  • 339823 11/13/10 05:19 n. muangma RunCqntrol, scalers, rocs : Run#8685
  • 339822 11/13/10 04:39 Hisham Calo anode currents
  • 339821 11/13/10 04:19 a-onl Start_Run_8685,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339820 11/13/10 04:19 a-onl End_of_Run_8684
  • 339819 11/13/10 04:09 n.muangma RunControl, Scaler, rocs :Run # 8684
  • 339818 11/13/10 03:14 a-onl Start_Run_8684,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339817 11/13/10 03:14 a-onl End_of_Run_8683
  • 339816 11/13/10 03:13 n. Muangma RunControl, Scaler, rocs for run#8683
  • 339815 11/13/10 03:05 auto_snap Radiation and Ozone snapshot
  • 339814 11/13/10 03:05 auto_snap Miscellaneous snapshot
  • 339813 11/13/10 03:05 auto_snap Hall A Tools snapshot
  • 339812 11/13/10 03:05 auto_snap Compton Polarimeter snapshot
  • 339811 11/13/10 03:04 auto_snap Beam Position and Current snapshot
  • 339810 11/13/10 02:20 a-onl Start_Run_8683,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339809 11/13/10 02:19 a-onl End_of_Run_8682
  • 339808 11/13/10 02:19 n.muangma RunControl, scaler, rocs run#8682
  • 339807 11/13/10 02:10 auto_snap Radiation and Ozone snapshot
  • 339806 11/13/10 02:10 auto_snap Miscellaneous snapshot
  • 339805 11/13/10 02:10 auto_snap Hall A Tools snapshot
  • 339804 11/13/10 02:10 auto_snap Compton Polarimeter snapshot
  • 339803 11/13/10 02:10 auto_snap Beam Position and Current snapshot
  • 339802 11/13/10 01:19 a-onl Start_Run_8682,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339801 11/13/10 01:19 a-onl End_of_Run_8681
  • 339800 11/13/10 01:18 n. muangma rc, scaler, rocs, 8681
  • 339799 11/13/10 00:32 A. Camsonne Check clean start - moved EB and ER to adaql2-192
  • 339798 11/13/10 00:18 a-onl Start_Run_8681,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339797 11/13/10 00:17 a-onl End_of_Run_8680
  • 339796 11/13/10 00:09 a-onl Start_Run_8680,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339795 11/13/10 00:03 a-onl End_of_Run_8679
  • 339794 11/13/10 00:02 Punjabi roc,rc,scaler screens
  • 339792 11/13/10 00:01 Boeglin Shift Summary
  • 339793 11/13/10 00:00 adaq Beam Time Accounting
  • Friday

  • 339791 11/12/10 23:30 Boeglin Harp scan at beginning of data taking
  • 339790 11/12/10 23:04 a-onl Start_Run_8679,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339789 11/12/10 23:02 A. Camsonne TS11 rebooted by itself
  • 339788 11/12/10 22:47 a-onl Start_Run_8678,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339787 11/12/10 22:43 a-onl End_of_Run_8677
  • 339786 11/12/10 22:43 Punjabi roc,rc,scaler screens
  • 339785 11/12/10 21:55 A. Camsonne Logged out and killed all ssh to adaql2
  • 339784 11/12/10 21:46 a-onl Start_Run_8677,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339783 11/12/10 21:42 a-onl End_of_Run_8676
  • 339782 11/12/10 21:41 Punjabi roc,rc,scaler screens
  • 339781 11/12/10 21:10 A. Camsonne Network statistics
  • 339780 11/12/10 21:09 A. Camsonne Network statistics
  • 339779 11/12/10 20:46 a-onl Start_Run_8676,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339778 11/12/10 20:42 S. Golge Hall A CryoTarget Snapshot swing
  • 339777 11/12/10 20:26 Boeglin CALO currents
  • 339776 11/12/10 20:02 a-onl Start_Run_8675,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339775 11/12/10 20:00 a-onl End_of_Run_8674
  • 339774 11/12/10 19:31 A. Camsonne Trying TrigDVCSNoC configuration starting run 8674
  • 339773 11/12/10 19:30 a-onl Start_Run_8674,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339772 11/12/10 19:29 a-onl End_of_Run_8673
  • 339771 11/12/10 19:28 a-onl Start_Run_8673,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - TrigDVCSNoC
  • 339770 11/12/10 19:26 a-onl End_of_Run_8672
  • 339769 11/12/10 19:25 a-onl End_of_Run_8671
  • 339768 11/12/10 19:21 a-onl Start_Run_8671,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam
  • 339767 11/12/10 19:15 a-onl End_of_Run_8667
  • 339766 11/12/10 19:10 Punjabi beam position
  • 339765 11/12/10 19:04 auto_snap Radiation and Ozone snapshot
  • 339764 11/12/10 19:04 auto_snap Miscellaneous snapshot
  • 339763 11/12/10 19:04 auto_snap Hall A Tools snapshot
  • 339762 11/12/10 19:04 auto_snap Compton Polarimeter snapshot
  • 339761 11/12/10 19:04 auto_snap Beam Position and Current snapshot
  • 339760 11/12/10 18:46 A. Camsonne Readded scaler read out for run 8667 - no significant change in dead time
  • 339759 11/12/10 18:46 a-onl Start_Run_8667,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam
  • 339758 11/12/10 18:44 a-onl End_of_Run_8666
  • 339757 11/12/10 18:41 adaq Spot++
  • 339756 11/12/10 18:38 A. Camsonne Run 8666 without scaler readout
  • 339755 11/12/10 18:38 a-onl Start_Run_8666,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam - no scaler readout in TS
  • 339754 11/12/10 18:36 a-onl End_of_Run_8665
  • 339753 11/12/10 18:35 Punjabi roc,rc,scaler screens
  • 339752 11/12/10 18:10 auto_snap Radiation and Ozone snapshot
  • 339751 11/12/10 18:10 auto_snap Miscellaneous snapshot
  • 339750 11/12/10 18:10 auto_snap Hall A Tools snapshot
  • 339749 11/12/10 18:10 auto_snap Compton Polarimeter snapshot
  • 339748 11/12/10 18:09 auto_snap Beam Position and Current snapshot
  • 339747 11/12/10 17:41 a-onl Start_Run_8665,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kin 2 4 pass beam
  • 339746 11/12/10 16:43 jroche injector spin setting A-ok (ct'd)
  • 339745 11/12/10 16:32 jroche helcity setting A-ok
  • 339744 11/12/10 16:04 a-onl End_of_Run_8664
  • 339742 11/12/10 16:00 a-onl Start_Run_8664,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test TrigDVCSBuff
  • 339741 11/12/10 16:00 A. Deur Day run summary
  • 339743 11/12/10 15:59 adaq Beam Time Accounting
  • 339740 11/12/10 15:45 a-onl Start_Run_8663,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TSOnly
  • 339739 11/12/10 15:45 a-onl End_of_Run_8661
  • 339738 11/12/10 15:44 A. Deur Wien angle
  • 339737 11/12/10 15:44 a-onl End_of_Run_8662
  • 339736 11/12/10 15:43 a-onl Start_Run_8662,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TSOnly
  • 339735 11/12/10 15:34 a-onl Start_Run_8661,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339734 11/12/10 15:13 a-onl Start_Run_8658,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRS
  • 339733 11/12/10 15:11 a-onl End_of_Run_8657
  • 339732 11/12/10 15:06 a-onl Start_Run_8657,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRS
  • 339731 11/12/10 14:58 a-onl End_of_Run_8656
  • 339730 11/12/10 14:58 a-onl End_of_Run_8653
  • 339729 11/12/10 14:56 a-onl End_of_Run_8655
  • 339728 11/12/10 14:55 a-onl Start_Run_8655,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRSBuff
  • 339727 11/12/10 14:54 a-onl Start_Run_8654,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRSBuff
  • 339726 11/12/10 14:54 a-onl End_of_Run_8654
  • 339725 11/12/10 14:53 a-onl End_of_Run_8652
  • 339724 11/12/10 14:50 a-onl Start_Run_8653,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRSBuff
  • 339723 11/12/10 14:50 a-onl End_of_Run_8650
  • 339722 11/12/10 14:48 a-onl Start_Run_8652,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRSBuff
  • 339721 11/12/10 14:47 a-onl End_of_Run_8651
  • 339720 11/12/10 14:45 a-onl Start_Run_8651,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339719 11/12/10 14:44 a-onl Start_Run_8650,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339718 11/12/10 14:43 a-onl End_of_Run_8649
  • 339717 11/12/10 14:26 a-onl Start_Run_8649,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339716 11/12/10 13:47 Fansler Moeller
  • 339715 11/12/10 12:04 a-onl End_of_Run_8648
  • 339714 11/12/10 12:01 a-onl Start_Run_8648,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339713 11/12/10 11:48 Carlos Time distribution for trigger routine
  • 339712 11/12/10 11:24 a-onl End_of_Run_8647
  • 339711 11/12/10 11:21 a-onl Start_Run_8647,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339710 11/12/10 11:19 A. Camsonne TS Busy LeftHRSBuff
  • 339709 11/12/10 11:11 a-onl End_of_Run_8645
  • 339708 11/12/10 11:07 a-onl Start_Run_8645,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test TrigDVCSBuff
  • 339707 11/12/10 11:06 a-onl End_of_Run_8644
  • 339706 11/12/10 11:04 auto_snap Radiation and Ozone snapshot
  • 339705 11/12/10 11:04 auto_snap Miscellaneous snapshot
  • 339704 11/12/10 11:04 auto_snap Hall A Tools snapshot
  • 339703 11/12/10 11:04 auto_snap Compton Polarimeter snapshot
  • 339702 11/12/10 11:04 auto_snap Beam Position and Current snapshot
  • 339701 11/12/10 10:53 a-onl Start_Run_8644,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test LeftHRSBuff
  • 339700 11/12/10 10:09 auto_snap Radiation and Ozone snapshot
  • 339699 11/12/10 10:09 auto_snap Miscellaneous snapshot
  • 339698 11/12/10 10:09 auto_snap Hall A Tools snapshot
  • 339697 11/12/10 10:09 auto_snap Compton Polarimeter snapshot
  • 339696 11/12/10 10:09 auto_snap Beam Position and Current snapshot
  • 339695 11/12/10 09:06 Fansler Moeller
  • 339694 11/12/10 09:04 a-onl End_of_Run_8643
  • 339693 11/12/10 09:02 a-onl Start_Run_8643,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339692 11/12/10 09:01 a-onl End_of_Run_8642
  • 339691 11/12/10 09:00 Saha Run 8642 Scalar Entry
  • 339690 11/12/10 08:59 Saha Run 8642 Screen entry
  • 339689 11/12/10 08:58 A. Deur Target parameter snapshot and stripcharts
  • 339688 11/12/10 08:57 J. Zhang Hall will be in controlled access from 9 AM, Friday
  • 339687 11/12/10 08:54 A. Deur Snapshot of calorimeter anode currents
  • 339686 11/12/10 08:49 a-onl Start_Run_8642,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339685 11/12/10 08:47 a-onl End_of_Run_8641
  • 339684 11/12/10 08:45 A. Deur snapshot of Hall A tools
  • 339683 11/12/10 08:25 Saha Run 8641 Error in online analysis
  • 339682 11/12/10 08:13 Saha BPM y distribution Run 8640
  • 339681 11/12/10 08:05 a-onl Start_Run_8641,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339680 11/12/10 08:03 a-onl End_of_Run_8640
  • 339679 11/12/10 07:59 adaq Beam Time Accounting
  • 339678 11/12/10 07:57 R. Michaels Owl Shift Summary
  • 339677 11/12/10 07:43 a-onl Start_Run_8640,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339676 11/12/10 07:42 a-onl End_of_Run_8639
  • 339675 11/12/10 07:00 Carlos Re: wide Y beam
  • 339674 11/12/10 06:50 a-onl Start_Run_8639,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339673 11/12/10 06:46 itard entries end of run
  • 339672 11/12/10 06:45 a-onl End_of_Run_8637
  • 339671 11/12/10 06:25 R. Michaels helicity problems gone
  • 339670 11/12/10 05:46 a-onl Start_Run_8637,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339669 11/12/10 05:46 R. Michaels reboot ROC17 twice ?
  • 339668 11/12/10 05:39 a-onl End_of_Run_8635
  • 339667 11/12/10 05:20 a-onl Start_Run_8635,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339666 11/12/10 05:18 itard entries end of run
  • 339665 11/12/10 05:03 R. Michaels dark color on floor near Moller (water ?)
  • 339664 11/12/10 04:16 R. Michaels CODA reboot, ROC17 again
  • 339663 11/12/10 04:14 a-onl Start_Run_8634,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339662 11/12/10 03:43 a-onl Start_Run_8632,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339661 11/12/10 03:25 a-onl End_of_Run_8631
  • 339660 11/12/10 03:24 itard entries end of run
  • 339659 11/12/10 03:04 auto_snap Radiation and Ozone snapshot
  • 339658 11/12/10 03:04 auto_snap Miscellaneous snapshot
  • 339657 11/12/10 03:04 auto_snap Hall A Tools snapshot
  • 339656 11/12/10 03:04 auto_snap Compton Polarimeter snapshot
  • 339655 11/12/10 03:04 auto_snap Beam Position and Current snapshot
  • 339654 11/12/10 02:26 R. Michaels run 8629 died, restart coda
  • 339653 11/12/10 02:26 a-onl Start_Run_8631,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339652 11/12/10 02:18 a-onl End_of_Run_8629
  • 339651 11/12/10 02:09 auto_snap Radiation and Ozone snapshot
  • 339650 11/12/10 02:09 auto_snap Miscellaneous snapshot
  • 339649 11/12/10 02:09 auto_snap Hall A Tools snapshot
  • 339648 11/12/10 02:09 auto_snap Compton Polarimeter snapshot
  • 339647 11/12/10 02:09 auto_snap Beam Position and Current snapshot
  • 339646 11/12/10 02:04 R. Michaels re: wide Y beam
  • 339645 11/12/10 01:54 itard bad BPMA y and BPMB y
  • 339644 11/12/10 01:28 a-onl Start_Run_8629,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339643 11/12/10 01:27 a-onl End_of_Run_8628
  • 339642 11/12/10 01:27 itard entries end of run
  • 339641 11/12/10 00:28 a-onl Start_Run_8628,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339640 11/12/10 00:27 a-onl End_of_Run_8627
  • 339639 11/12/10 00:27 itard entries end of run
  • 339638 11/12/10 00:16 Carlos Re: sync plots
  • 339637 11/12/10 00:10 efuchey Restarting taking compton data
  • 339636 11/12/10 00:08 R. Michaels ROC synch check looks odd
  • 339635 11/12/10 00:06 efuchey Compton cavity scan
  • 339634 11/12/10 00:03 Boeglin Shift Summary
  • 339633 11/12/10 00:01 adaq Beam Time Accounting
  • Thursday

  • 339632 11/11/10 23:48 efuchey Compton table scan for 11-11-10
  • 339631 11/11/10 23:29 a-onl Start_Run_8627,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=TrigDVCSBuff Production
  • 339630 11/11/10 23:26 a-onl End_of_Run_8626
  • 339629 11/11/10 23:26 sutera rates run 8626
  • 339628 11/11/10 23:24 sutera roc windows run 8626
  • 339627 11/11/10 23:00 efuchey Compton Rates for tonight.
  • 339626 11/11/10 22:57 a-onl Start_Run_8626,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering, trig eff.
  • 339625 11/11/10 22:55 a-onl End_of_Run_8625
  • 339624 11/11/10 22:54 a-onl Start_Run_8625,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering, trig eff.
  • 339623 11/11/10 22:51 a-onl End_of_Run_8624
  • 339622 11/11/10 22:50 a-onl Start_Run_8624,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering, trig eff.
  • 339621 11/11/10 22:49 S.Golge Hall A cryo snap swing
  • 339620 11/11/10 22:26 Boeglin CALO currents
  • 339619 11/11/10 22:25 Boeglin Harp Scan Result/ BPM not yet OK
  • 339618 11/11/10 21:39 Boeglin first beam after beam studies
  • 339617 11/11/10 20:40 a-onl Start_Run_8621,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering
  • 339616 11/11/10 20:40 a-onl Start_Run_8622,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering
  • 339615 11/11/10 20:40 a-onl Start_Run_8623,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering
  • 339614 11/11/10 20:40 a-onl End_of_Run_8621
  • 339613 11/11/10 20:40 a-onl End_of_Run_8622
  • 339612 11/11/10 20:40 a-onl End_of_Run_8623
  • 339611 11/11/10 20:36 A. Camsonne Dead time at Trigger Supervisor
  • 339610 11/11/10 20:03 a-onl End_of_Run_8620
  • 339609 11/11/10 19:54 a-onl Start_Run_8620,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering
  • 339608 11/11/10 19:30 a-onl End_of_Run_8619
  • 339607 11/11/10 19:24 a-onl End_of_Run_8618
  • 339606 11/11/10 19:07 a-onl Start_Run_8618,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS without buffering
  • 339605 11/11/10 19:05 a-onl End_of_Run_8617
  • 339604 11/11/10 19:05 a-onl Start_Run_8617,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS with buffering
  • 339603 11/11/10 19:04 auto_snap Radiation and Ozone snapshot
  • 339602 11/11/10 19:03 auto_snap Miscellaneous snapshot
  • 339601 11/11/10 19:03 auto_snap Hall A Tools snapshot
  • 339600 11/11/10 19:03 auto_snap Compton Polarimeter snapshot
  • 339599 11/11/10 19:03 auto_snap Beam Position and Current snapshot
  • 339598 11/11/10 18:45 a-onl Start_Run_8616,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS with buffering
  • 339597 11/11/10 18:43 a-onl End_of_Run_8615
  • 339596 11/11/10 18:41 a-onl Start_Run_8615,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test LeftHRS with buffering
  • 339595 11/11/10 18:39 a-onl End_of_Run_8614
  • 339594 11/11/10 18:37 a-onl Start_Run_8614,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339593 11/11/10 18:36 a-onl Start_Run_8613,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339592 11/11/10 18:32 A. Camsonne Compton Edet webcam plugged on APC port 5
  • 339591 11/11/10 18:28 J. Zhang Run Plan
  • 339590 11/11/10 18:09 auto_snap Radiation and Ozone snapshot
  • 339589 11/11/10 18:09 auto_snap Miscellaneous snapshot
  • 339588 11/11/10 18:09 auto_snap Hall A Tools snapshot
  • 339587 11/11/10 18:09 auto_snap Compton Polarimeter snapshot
  • 339586 11/11/10 18:09 auto_snap Beam Position and Current snapshot
  • 339585 11/11/10 17:52 a-onl Start_Run_8605,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339583 11/11/10 17:52 a-onl End_of_Run_8605
  • 339582 11/11/10 16:53 efuchey 11-11-10 Owl shift summary
  • 339581 11/11/10 16:29 J. Butler(te Flammable Gas Detectors
  • 339580 11/11/10 16:00 adaq Beam Time Accounting
  • 339579 11/11/10 15:59 J. Zhang Hall will be open till beam study finished, may be 8 PM
  • 339578 11/11/10 15:54 A. Camsonne TS Busy and L1A on scope
  • 339577 11/11/10 15:53 A. Camsonne Replaced cable for Norm L1A
  • 339576 11/11/10 15:49 Hisham Day shift summary
  • 339575 11/11/10 15:48 A. Camsonne Put copy of L1A on spare scaler in TS crate channel 15
  • 339574 11/11/10 15:15 Hisham Target snapshot
  • 339573 11/11/10 15:14 a-onl Start_Run_8604,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339572 11/11/10 14:45 a-onl End_of_Run_8599
  • 339571 11/11/10 14:42 a-onl Start_Run_8599,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339570 11/11/10 14:31 a-onl End_of_Run_8597
  • 339569 11/11/10 14:31 A. Camsonne L1A rate fluctuation at 3KHz of trigger rates
  • 339568 11/11/10 14:22 A. Camsonne Dead time checks
  • 339567 11/11/10 14:20 a-onl Start_Run_8597,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - Removed checks and sending directly to CODA buffer
  • 339566 11/11/10 13:50 Fansler Left Dipole
  • 339565 11/11/10 13:28 a-onl Start_Run_8587,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339564 11/11/10 13:28 a-onl Start_Run_8588,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339563 11/11/10 13:28 a-onl End_of_Run_8587
  • 339562 11/11/10 13:28 a-onl End_of_Run_8588
  • 339561 11/11/10 13:06 A. Camsonne DVCS cosmics trigger on channel 31 ( starting at 1 ) of scaler of ROC17
  • 339560 11/11/10 12:44 A. Camsonne ROC3 and ROC4 blaster - 2nd attempt
  • 339559 11/11/10 12:44 A. Camsonne ROC3 and ROC4 blaster - 2nd attempt
  • 339558 11/11/10 12:35 A. Camsonne Blaster test ROCs
  • 339557 11/11/10 12:28 Hisham Left Arm Dipole Alarm
  • 339556 11/11/10 12:19 adaq left arm Dipole system alarm
  • 339555 11/11/10 12:17 A. Camsonne DVCS readout list switched back to polling
  • 339554 11/11/10 12:09 hisham Calorimeter anode currents
  • 339553 11/11/10 12:09 a-onl Start_Run_8579,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339552 11/11/10 12:09 a-onl Start_Run_8580,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339551 11/11/10 12:09 a-onl End_of_Run_8578
  • 339550 11/11/10 12:09 a-onl End_of_Run_8579
  • 339549 11/11/10 12:09 a-onl End_of_Run_8580
  • 339548 11/11/10 11:56 a-onl Start_Run_8577,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339547 11/11/10 11:55 a-onl Start_Run_8578,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339546 11/11/10 11:55 a-onl End_of_Run_8577
  • 339545 11/11/10 11:34 a-onl Start_Run_8576,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339544 11/11/10 11:34 a-onl End_of_Run_8575
  • 339543 11/11/10 11:32 A. Camsonne Scintillator paddle HV on for calorimeter tests
  • 339542 11/11/10 11:29 a-onl Start_Run_8575,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test DAQ - ROC17 ADC sync
  • 339541 11/11/10 11:03 auto_snap Radiation and Ozone snapshot
  • 339540 11/11/10 11:03 auto_snap Miscellaneous snapshot
  • 339539 11/11/10 11:03 auto_snap Hall A Tools snapshot
  • 339538 11/11/10 11:03 auto_snap Compton Polarimeter snapshot
  • 339537 11/11/10 11:03 auto_snap Beam Position and Current snapshot
  • 339536 11/11/10 10:33 Bob / Julie helicity issues, part N
  • 339535 11/11/10 10:09 auto_snap Radiation and Ozone snapshot
  • 339534 11/11/10 10:08 auto_snap Miscellaneous snapshot
  • 339533 11/11/10 10:08 auto_snap Hall A Tools snapshot
  • 339532 11/11/10 10:08 auto_snap Compton Polarimeter snapshot
  • 339531 11/11/10 10:08 auto_snap Beam Position and Current snapshot
  • 339530 11/11/10 09:14 cshloka Cerenkov efficiency
  • 339529 11/11/10 08:13 adaq Beam Time Accounting
  • 339528 11/11/10 08:08 hisham moved to empty target
  • 339527 11/11/10 07:03 efuchey Beam stops, go to controlled access
  • 339526 11/11/10 07:00 a-onl End_of_Run_8574
  • 339525 11/11/10 06:59 sutera rates run 8574
  • 339524 11/11/10 06:58 sutera roc windows run 8574
  • 339523 11/11/10 06:00 a-onl Start_Run_8574,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339522 11/11/10 06:00 a-onl End_of_Run_8573
  • 339521 11/11/10 05:58 sutera rares run 8573
  • 339520 11/11/10 05:57 sutera roc windows run 8573
  • 339519 11/11/10 05:51 efuchey Calo HV crash
  • 339518 11/11/10 05:01 a-onl Start_Run_8573,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339517 11/11/10 05:01 a-onl End_of_Run_8572
  • 339516 11/11/10 04:59 sutera rates run 8572
  • 339515 11/11/10 04:58 sutera roc windows run 8572
  • 339514 11/11/10 04:03 a-onl Start_Run_8572,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339513 11/11/10 04:02 a-onl End_of_Run_8571
  • 339512 11/11/10 04:01 sutera rates run 8571
  • 339511 11/11/10 04:00 sutera roc windows run 8571
  • 339510 11/11/10 03:20 efuchey Hall A tools screenshots
  • 339509 11/11/10 03:17 efuchey PbF2 Calo base currents
  • 339508 11/11/10 03:03 auto_snap Radiation and Ozone snapshot
  • 339507 11/11/10 03:03 auto_snap Miscellaneous snapshot
  • 339506 11/11/10 03:03 auto_snap Hall A Tools snapshot
  • 339505 11/11/10 03:03 auto_snap Compton Polarimeter snapshot
  • 339504 11/11/10 03:03 auto_snap Beam Position and Current snapshot
  • 339503 11/11/10 03:00 a-onl Start_Run_8571,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339502 11/11/10 02:59 a-onl End_of_Run_8570
  • 339501 11/11/10 02:55 sutera rates run 8570
  • 339500 11/11/10 02:54 efuchey Compton rates
  • 339499 11/11/10 02:53 sutera roc windows run 8570
  • 339498 11/11/10 02:08 auto_snap Radiation and Ozone snapshot
  • 339497 11/11/10 02:08 auto_snap Miscellaneous snapshot
  • 339496 11/11/10 02:08 auto_snap Hall A Tools snapshot
  • 339495 11/11/10 02:08 auto_snap Compton Polarimeter snapshot
  • 339494 11/11/10 02:08 auto_snap Beam Position and Current snapshot
  • 339493 11/11/10 02:03 a-onl Start_Run_8570,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339492 11/11/10 02:01 a-onl End_of_Run_8569
  • 339491 11/11/10 02:00 sutera rates for run 8569
  • 339490 11/11/10 01:59 sutera ROC windows for 8569
  • 339489 11/11/10 01:49 adaq Rates for 11-11-10
  • 339488 11/11/10 01:36 jpchen setting for unused target
  • 339487 11/11/10 00:58 a-onl Start_Run_8569,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339486 11/11/10 00:57 a-onl End_of_Run_8568
  • 339485 11/11/10 00:03 Boeglin Shift Summary
  • 339484 11/11/10 00:01 adaq Beam Time Accounting
  • Wednesday

  • 339483 11/10/10 23:57 a-onl Start_Run_8568,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339482 11/10/10 23:56 a-onl Start_Run_8555,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339481 11/10/10 23:55 a-onl End_of_Run_8567
  • 339480 11/10/10 23:30 skoirala 8567
  • 339479 11/10/10 23:03 a-onl Start_Run_8567,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339478 11/10/10 23:02 a-onl End_of_Run_8566
  • 339477 11/10/10 22:21 skoirala Run 8566
  • 339476 11/10/10 22:01 a-onl Start_Run_8566,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339475 11/10/10 22:00 a-onl End_of_Run_8565
  • 339474 11/10/10 21:27 skoirala Run8565
  • 339473 11/10/10 21:00 a-onl Start_Run_8565,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339472 11/10/10 20:59 a-onl End_of_Run_8564
  • 339471 11/10/10 20:29 skoirala Run 8564
  • 339470 11/10/10 20:20 Boeglin CALO DC levels
  • 339469 11/10/10 19:58 a-onl Start_Run_8564,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339468 11/10/10 19:57 a-onl End_of_Run_8563
  • 339467 11/10/10 19:57 skoirala Run 8563
  • 339466 11/10/10 19:03 auto_snap Radiation and Ozone snapshot
  • 339465 11/10/10 19:03 auto_snap Miscellaneous snapshot
  • 339464 11/10/10 19:02 auto_snap Hall A Tools snapshot
  • 339463 11/10/10 19:02 auto_snap Compton Polarimeter snapshot
  • 339462 11/10/10 19:02 auto_snap Beam Position and Current snapshot
  • 339461 11/10/10 19:00 a-onl Start_Run_8563,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339460 11/10/10 18:59 a-onl End_of_Run_8562
  • 339459 11/10/10 18:53 skoirala Run 8562
  • 339458 11/10/10 18:08 auto_snap Radiation and Ozone snapshot
  • 339457 11/10/10 18:08 auto_snap Miscellaneous snapshot
  • 339456 11/10/10 18:08 auto_snap Hall A Tools snapshot
  • 339455 11/10/10 18:08 auto_snap Compton Polarimeter snapshot
  • 339454 11/10/10 18:08 auto_snap Beam Position and Current snapshot
  • 339453 11/10/10 18:00 a-onl Start_Run_8562,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339452 11/10/10 17:53 Boeglin Harp scan
  • 339451 11/10/10 17:23 a-onl End_of_Run_8561
  • 339450 11/10/10 17:22 a-onl Start_Run_8561,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339449 11/10/10 17:21 a-onl End_of_Run_8560
  • 339448 11/10/10 17:21 a-onl Start_Run_8560,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339447 11/10/10 17:18 a-onl End_of_Run_8559
  • 339446 11/10/10 17:17 a-onl Start_Run_8559,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339445 11/10/10 17:17 a-onl End_of_Run_8558
  • 339444 11/10/10 17:14 a-onl Start_Run_8558,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339443 11/10/10 17:13 a-onl End_of_Run_8557
  • 339442 11/10/10 17:12 a-onl Start_Run_8557,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339441 11/10/10 17:11 a-onl End_of_Run_8556
  • 339440 11/10/10 17:11 a-onl Start_Run_8556,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339439 11/10/10 17:05 a-onl End_of_Run_8555
  • 339438 11/10/10 17:05 a-onl End_of_Run_8552
  • 339437 11/10/10 17:04 a-onl End_of_Run_8554
  • 339436 11/10/10 16:55 a-onl Start_Run_8554,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339435 11/10/10 16:54 a-onl End_of_Run_8553
  • 339434 11/10/10 16:52 A. Camsonne Added timers to DVCS crates read out list
  • 339433 11/10/10 16:00 adaq Beam Time Accounting
  • 339432 11/10/10 15:57 Saha PSS Down No Beam
  • 339431 11/10/10 15:57 Saha Shift Summary
  • 339430 11/10/10 15:54 a-onl Start_Run_8553,Run_type=Cosmics,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339429 11/10/10 14:46 a-onl Start_Run_8552,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339428 11/10/10 14:44 a-onl End_of_Run_8551
  • 339427 11/10/10 14:39 a-onl Start_Run_8551,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339426 11/10/10 14:38 a-onl End_of_Run_8550
  • 339425 11/10/10 14:20 a-onl Start_Run_8550,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339424 11/10/10 14:19 a-onl End_of_Run_8549
  • 339423 11/10/10 14:04 hyde rocs 17 18 run 8549
  • 339422 11/10/10 14:04 hyde eb, roc14 run 8549
  • 339421 11/10/10 13:57 hyde Deadtime run 8549
  • 339420 11/10/10 13:44 a-onl Start_Run_8549,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339419 11/10/10 13:27 hyde Coda no roc14
  • 339418 11/10/10 13:25 a-onl Start_Run_8545,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339417 11/10/10 13:21 a-onl End_of_Run_8544
  • 339416 11/10/10 13:21 hyde coda...coda
  • 339415 11/10/10 13:20 a-onl Start_Run_8544,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339414 11/10/10 13:17 hyde roc terminal
  • 339413 11/10/10 13:16 hyde coda crash
  • 339412 11/10/10 13:15 a-onl Start_Run_8543,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339411 11/10/10 13:11 a-onl Start_Run_8542,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339410 11/10/10 13:10 a-onl End_of_Run_8541
  • 339409 11/10/10 12:37 jmammei JT valve/ heater for unused target
  • 339408 11/10/10 12:34 a-onl Start_Run_8541,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 Deuterium
  • 339407 11/10/10 12:32 jmammei target moved to 15cm LD2
  • 339406 11/10/10 12:28 hyde Switch to LD2 15 cm
  • 339405 11/10/10 12:27 a-onl End_of_Run_8540
  • 339404 11/10/10 11:30 a-onl Start_Run_8540,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1
  • 339403 11/10/10 11:29 a-onl End_of_Run_8539
  • 339402 11/10/10 11:29 hyde S2, Cer rates
  • 339401 11/10/10 11:28 hyde scaler screen snapshot
  • 339400 11/10/10 11:27 hyde roc 18
  • 339399 11/10/10 11:26 hyde Run 8539 Roc 14 and 17
  • 339398 11/10/10 11:02 auto_snap Radiation and Ozone snapshot
  • 339397 11/10/10 11:02 auto_snap Miscellaneous snapshot
  • 339396 11/10/10 11:02 auto_snap Hall A Tools snapshot
  • 339395 11/10/10 11:02 auto_snap Compton Polarimeter snapshot
  • 339394 11/10/10 11:02 auto_snap Beam Position and Current snapshot
  • 339393 11/10/10 10:52 hyde Cerenkov Efficiency Run 8537
  • 339392 11/10/10 10:26 a-onl Start_Run_8539,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1
  • 339391 11/10/10 10:25 a-onl End_of_Run_8538
  • 339390 11/10/10 10:08 auto_snap Radiation and Ozone snapshot
  • 339389 11/10/10 10:08 auto_snap Miscellaneous snapshot
  • 339388 11/10/10 10:07 auto_snap Hall A Tools snapshot
  • 339387 11/10/10 10:07 auto_snap Compton Polarimeter snapshot
  • 339386 11/10/10 10:07 auto_snap Beam Position and Current snapshot
  • 339385 11/10/10 09:33 a-onl Start_Run_8538,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1
  • 339384 11/10/10 09:31 a-onl End_of_Run_8537
  • 339383 11/10/10 09:13 a-onl Start_Run_8537,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger Efficiency run LeftHRS ps3=ps4=1
  • 339382 11/10/10 09:09 a-onl End_of_Run_8536
  • 339381 11/10/10 09:06 jmammei Target screens
  • 339380 11/10/10 08:09 a-onl Start_Run_8536,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339378 11/10/10 08:01 Hisham Owl shift summary
  • 339379 11/10/10 08:00 adaq Beam Time Accounting
  • 339377 11/10/10 07:53 Hisham Beam down
  • 339376 11/10/10 07:51 a-onl End_of_Run_8535
  • 339375 11/10/10 07:51 hkkang end-of-run 8535
  • 339374 11/10/10 07:01 a-onl Start_Run_8535,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339373 11/10/10 07:01 a-onl End_of_Run_8534
  • 339372 11/10/10 07:00 hkkang end-of-run 8534
  • 339371 11/10/10 06:02 a-onl Start_Run_8534,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339370 11/10/10 06:02 a-onl End_of_Run_8533
  • 339369 11/10/10 06:02 hoyoung Target screens
  • 339368 11/10/10 06:01 hkkang end-of-run 8533
  • 339367 11/10/10 06:01 Hisham Calorimeter dc level
  • 339366 11/10/10 05:03 a-onl Start_Run_8533,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339365 11/10/10 05:03 a-onl End_of_Run_8532
  • 339364 11/10/10 05:02 hkkang end-of-run 8532
  • 339363 11/10/10 04:01 a-onl Start_Run_8532,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339362 11/10/10 04:00 a-onl End_of_Run_8531
  • 339361 11/10/10 04:00 a-onl End_of_Run_8528
  • 339360 11/10/10 04:00 hkkang end-of-run 8531
  • 339359 11/10/10 03:07 Hisham S1 inefficiency follow up
  • 339358 11/10/10 03:02 auto_snap Radiation and Ozone snapshot
  • 339357 11/10/10 03:02 auto_snap Miscellaneous snapshot
  • 339356 11/10/10 03:02 auto_snap Hall A Tools snapshot
  • 339355 11/10/10 03:02 auto_snap Compton Polarimeter snapshot
  • 339354 11/10/10 03:02 auto_snap Beam Position and Current snapshot
  • 339353 11/10/10 02:59 a-onl Start_Run_8531,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339352 11/10/10 02:58 a-onl End_of_Run_8530
  • 339351 11/10/10 02:57 hkkang end-of-run
  • 339350 11/10/10 02:07 auto_snap Radiation and Ozone snapshot
  • 339349 11/10/10 02:07 auto_snap Miscellaneous snapshot
  • 339348 11/10/10 02:07 auto_snap Hall A Tools snapshot
  • 339347 11/10/10 02:07 auto_snap Compton Polarimeter snapshot
  • 339346 11/10/10 02:07 auto_snap Beam Position and Current snapshot
  • 339345 11/10/10 02:01 hoyoung Target screens
  • 339344 11/10/10 01:58 a-onl Start_Run_8530,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339343 11/10/10 01:58 a-onl End_of_Run_8529
  • 339342 11/10/10 01:57 hkkang end-of-run
  • 339341 11/10/10 00:48 a-onl Start_Run_8529,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339340 11/10/10 00:48 a-onl End_of_Run_8527
  • 339339 11/10/10 00:42 hkkang end-of-run
  • 1 11/10/10 00:08 Beoglin Run Summary
  • 339338 11/10/10 00:07 adaq Beam Time Accounting
  • Tuesday

  • 339337 11/09/10 23:47 a-onl Start_Run_8528,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339336 11/09/10 23:46 Min Huang Target status
  • 339335 11/09/10 23:46 pzhu roc,rc,scaler screenshot for run 8527
  • 339334 11/09/10 22:51 a-onl Start_Run_8527,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339333 11/09/10 22:51 pzhu roc,rc,scaler screen for run 8526
  • 339332 11/09/10 22:50 a-onl End_of_Run_8526
  • 339331 11/09/10 21:41 Boeglin Q3 (left) is back
  • 339330 11/09/10 21:40 a-onl Start_Run_8526,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339329 11/09/10 21:38 Min Huang target status
  • 339328 11/09/10 21:29 Boeglin Q3 Trip
  • 339327 11/09/10 21:25 Boeglin Q3 levels back up
  • 339326 11/09/10 21:06 a-onl End_of_Run_8525
  • 339325 11/09/10 20:54 a-onl Start_Run_8525,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339324 11/09/10 20:53 a-onl End_of_Run_8524
  • 339323 11/09/10 20:53 pzhu roc,rc,scaler screen for run 8524
  • 339322 11/09/10 20:03 adaq cal dc levels
  • 339321 11/09/10 19:52 a-onl Start_Run_8524,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339320 11/09/10 19:50 a-onl End_of_Run_8523
  • 339319 11/09/10 19:36 a-onl End_of_Run_8522
  • 339318 11/09/10 19:36 a-onl Start_Run_8523,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339317 11/09/10 19:35 pzhu roc,rc,scaler screenshot for run 8522
  • 339316 11/09/10 19:02 auto_snap Radiation and Ozone snapshot
  • 339315 11/09/10 19:02 auto_snap Miscellaneous snapshot
  • 339314 11/09/10 19:02 auto_snap Hall A Tools snapshot
  • 339313 11/09/10 19:02 auto_snap Compton Polarimeter snapshot
  • 339312 11/09/10 19:01 auto_snap Beam Position and Current snapshot
  • 339311 11/09/10 18:30 a-onl Start_Run_8522,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339310 11/09/10 18:29 a-onl End_of_Run_8521
  • 339309 11/09/10 18:28 pzhu roc,rc,scaler screenshot for run 8521
  • 339308 11/09/10 18:20 pzhu calo analysis for run 8521
  • 339307 11/09/10 18:07 auto_snap Radiation and Ozone snapshot
  • 339306 11/09/10 18:07 auto_snap Miscellaneous snapshot
  • 339305 11/09/10 18:07 auto_snap Hall A Tools snapshot
  • 339304 11/09/10 18:07 auto_snap Compton Polarimeter snapshot
  • 339303 11/09/10 18:07 auto_snap Beam Position and Current snapshot
  • 339302 11/09/10 17:31 a-onl Start_Run_8521,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339301 11/09/10 17:29 a-onl End_of_Run_8520
  • 339300 11/09/10 17:10 a-onl Start_Run_8520,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339299 11/09/10 17:09 a-onl End_of_Run_8519
  • 339298 11/09/10 17:09 pzhu roc,rc,scaler screenshot for run 8519
  • 339297 11/09/10 16:39 pzhu calo analysis for run 8519
  • 339296 11/09/10 16:15 Min Huang target status
  • 339295 11/09/10 16:11 a-onl Start_Run_8519,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339294 11/09/10 16:06 A. Tobias & Adjusted Loop1 (LD2) JT Valve to lower HP Heater to about 90W
  • 339292 11/09/10 16:01 Holmstrom Day Shift Summary
  • 339293 11/09/10 16:00 adaq Beam Time Accounting
  • 339291 11/09/10 15:58 A. Tobias RunControl for run 8518: short production run (beam OFF for Hall B)
  • 339290 11/09/10 15:45 a-onl End_of_Run_8518
  • 339289 11/09/10 15:41 A. Tobias Target Status: Loop 1 (LD2)
  • 339288 11/09/10 15:33 a-onl Start_Run_8518,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339287 11/09/10 15:33 a-onl End_of_Run_8517
  • 339286 11/09/10 15:13 A. Tobias RunControl and ROCs for run 8517
  • 339285 11/09/10 14:38 a-onl Start_Run_8517,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339284 11/09/10 14:37 a-onl End_of_Run_8516
  • 339283 11/09/10 14:33 A. Tobias RunControl and ROCs for run 8516
  • 339282 11/09/10 13:38 a-onl Start_Run_8516,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339281 11/09/10 13:37 a-onl End_of_Run_8511
  • 339280 11/09/10 13:37 a-onl End_of_Run_8515
  • 339279 11/09/10 13:32 A. Tobias RunControl and ROCs for run 8515
  • 339278 11/09/10 12:38 a-onl Start_Run_8515,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339277 11/09/10 12:38 a-onl End_of_Run_8514
  • 339276 11/09/10 12:37 A. Tobias RunControl and ROCs for run 8514
  • 339275 11/09/10 11:38 a-onl Start_Run_8514,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339274 11/09/10 11:37 a-onl End_of_Run_8513
  • 339273 11/09/10 11:34 A. Tobias RunControl and ROCs for run 8513
  • 339272 11/09/10 11:01 auto_snap Radiation and Ozone snapshot
  • 339271 11/09/10 11:01 auto_snap Miscellaneous snapshot
  • 339270 11/09/10 11:01 auto_snap Hall A Tools snapshot
  • 339269 11/09/10 11:01 auto_snap Compton Polarimeter snapshot
  • 339268 11/09/10 11:01 auto_snap Beam Position and Current snapshot
  • 339267 11/09/10 10:39 a-onl Start_Run_8513,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339266 11/09/10 10:38 a-onl End_of_Run_8512
  • 339265 11/09/10 10:33 A. Tobias RunControl and ROCs for run 8512
  • 339264 11/09/10 10:07 auto_snap Radiation and Ozone snapshot
  • 339263 11/09/10 10:07 auto_snap Miscellaneous snapshot
  • 339262 11/09/10 10:07 auto_snap Hall A Tools snapshot
  • 339261 11/09/10 10:07 auto_snap Compton Polarimeter snapshot
  • 339260 11/09/10 10:06 auto_snap Beam Position and Current snapshot
  • 339259 11/09/10 09:40 a-onl Start_Run_8512,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339258 11/09/10 09:36 A. Tobias RunControl and ROCs for run 8511
  • 339257 11/09/10 09:15 R. Michaels info about helicity problems
  • 339256 11/09/10 08:40 a-onl Start_Run_8511,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339255 11/09/10 08:39 a-onl End_of_Run_8510
  • 339254 11/09/10 08:38 A. Tobias RunControl & ROCs for run 8510
  • 339253 11/09/10 08:01 adaq Beam Time Accounting
  • 339252 11/09/10 07:59 Hisham Owl shift summary &
  • 339251 11/09/10 07:37 a-onl Start_Run_8510,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trigDVCSbuff
  • 339250 11/09/10 07:36 a-onl End_of_Run_8507
  • 339249 11/09/10 07:35 a-onl End_of_Run_8509
  • 339248 11/09/10 07:32 hkkang end-of-run
  • 339247 11/09/10 07:08 a-onl Start_Run_8509,Run_type=Calibration,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=trigger efficiency test ps3=ps4=1, ps5=0
  • 339246 11/09/10 07:06 a-onl End_of_Run_8508
  • 339245 11/09/10 06:45 Hisham ADC readback
  • 339244 11/09/10 06:31 Hisham Run 8507
  • 339243 11/09/10 06:20 hoyoung Target screens
  • 339242 11/09/10 06:08 a-onl Start_Run_8508,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339241 11/09/10 06:03 hkkang end-of-run
  • 339240 11/09/10 05:59 hkkang Luminosity plots
  • 339239 11/09/10 05:06 a-onl Start_Run_8507,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339238 11/09/10 05:05 a-onl End_of_Run_8506
  • 339237 11/09/10 05:02 hkkang end-of-run
  • 339236 11/09/10 04:05 a-onl Start_Run_8506,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339235 11/09/10 04:04 a-onl End_of_Run_8505
  • 339234 11/09/10 04:03 hkkang end-of-run
  • 339233 11/09/10 03:06 Hisham Calotimeter currents
  • 339232 11/09/10 03:03 a-onl Start_Run_8505,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339231 11/09/10 03:03 a-onl End_of_Run_8504
  • 339230 11/09/10 03:01 auto_snap Radiation and Ozone snapshot
  • 339229 11/09/10 03:01 auto_snap Miscellaneous snapshot
  • 339228 11/09/10 03:01 auto_snap Hall A Tools snapshot
  • 339227 11/09/10 03:01 auto_snap Compton Polarimeter snapshot
  • 339226 11/09/10 03:01 auto_snap Beam Position and Current snapshot
  • 339225 11/09/10 02:56 hkkang end-of-run
  • 339224 11/09/10 02:26 hkkang shifted ADC histo.
  • 339223 11/09/10 02:06 auto_snap Radiation and Ozone snapshot
  • 339222 11/09/10 02:06 auto_snap Miscellaneous snapshot
  • 339221 11/09/10 02:06 auto_snap Hall A Tools snapshot
  • 339220 11/09/10 02:06 auto_snap Compton Polarimeter snapshot
  • 339219 11/09/10 02:06 auto_snap Beam Position and Current snapshot
  • 339218 11/09/10 02:06 hoyoung Target screens
  • 339217 11/09/10 01:59 a-onl Start_Run_8504,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339216 11/09/10 01:58 a-onl End_of_Run_8503
  • 339215 11/09/10 01:56 hkkang end-of-run
  • 339214 11/09/10 00:59 a-onl Start_Run_8503,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339213 11/09/10 00:58 a-onl End_of_Run_8502
  • 339212 11/09/10 00:53 hkkang end-of-run
  • 339211 11/09/10 00:07 adaq Beam Time Accounting
  • Monday

  • 339210 11/08/10 23:58 a-onl Start_Run_8502,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339209 11/08/10 23:57 Abdurahim Shift Summary
  • 339208 11/08/10 23:56 a-onl End_of_Run_8501
  • 339207 11/08/10 23:55 pzhu roc,rc,scaler screenshot for run 8501
  • 339206 11/08/10 23:18 Min Huang target status
  • 339205 11/08/10 23:17 pzhu pedestal shift in run 8501
  • 339204 11/08/10 22:55 a-onl Start_Run_8501,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339203 11/08/10 22:54 a-onl End_of_Run_8500
  • 339202 11/08/10 22:54 pzhu roc,rc,scaler screenshot for run 8500
  • 339201 11/08/10 22:14 pzhu a little pedestal shiftin run 8500
  • 339200 11/08/10 22:02 Abdurahim Calorimeter currents
  • 339199 11/08/10 21:57 a-onl Start_Run_8500,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339198 11/08/10 21:56 a-onl End_of_Run_8499
  • 339197 11/08/10 21:55 pzhu roc,rc,scaler screenshot for run 8499
  • 339196 11/08/10 21:15 pzhu pedestal shift in run 8499
  • 339195 11/08/10 21:07 Abdurahim Compton rates
  • 339194 11/08/10 21:07 Min Huang target status
  • 339193 11/08/10 20:56 a-onl Start_Run_8499,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339192 11/08/10 20:54 a-onl End_of_Run_8498
  • 339191 11/08/10 20:53 pzhu roc,rc,scaler screenshot for run 8498
  • 339190 11/08/10 20:37 pzhu roc,rc,scaler snapshot for run 8498
  • 339189 11/08/10 20:32 A. Camsonne Scaler rates
  • 339188 11/08/10 19:58 a-onl Start_Run_8498,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339187 11/08/10 19:57 pzhu roc,rc,scaler screen for run 8497
  • 339186 11/08/10 19:56 a-onl End_of_Run_8497
  • 339185 11/08/10 19:01 auto_snap Radiation and Ozone snapshot
  • 339184 11/08/10 19:01 auto_snap Miscellaneous snapshot
  • 339183 11/08/10 19:01 auto_snap Hall A Tools snapshot
  • 339182 11/08/10 19:01 auto_snap Compton Polarimeter snapshot
  • 339181 11/08/10 19:01 auto_snap Beam Position and Current snapshot
  • 339180 11/08/10 18:57 a-onl Start_Run_8497,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339179 11/08/10 18:56 pzhu roc,rc,scaler screen for run8496, TrigDVCSBuff
  • 339178 11/08/10 18:54 a-onl End_of_Run_8496
  • 339177 11/08/10 18:54 a-onl Start_Run_8492,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test
  • 339176 11/08/10 18:42 pzhu The replay analysis seems to be OK after Carlos fix the pedtals
  • 339175 11/08/10 18:32 Min Huang target status
  • 339174 11/08/10 18:06 auto_snap Radiation and Ozone snapshot
  • 339173 11/08/10 18:06 auto_snap Miscellaneous snapshot
  • 339172 11/08/10 18:06 auto_snap Hall A Tools snapshot
  • 339171 11/08/10 18:06 auto_snap Compton Polarimeter snapshot
  • 339170 11/08/10 18:06 auto_snap Beam Position and Current snapshot
  • 339169 11/08/10 17:53 Abdurahim Beam is back
  • 339168 11/08/10 17:52 a-onl Start_Run_8496,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=production, trig dvcs buff
  • 339167 11/08/10 17:43 A. Camsonne Pedestal run 8495
  • 339166 11/08/10 17:43 a-onl End_of_Run_8495
  • 339164 11/08/10 17:43 a-onl Start_Run_8495,Run_type=Pedestal,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run with new board
  • 339163 11/08/10 17:42 a-onl End_of_Run_8494
  • 339162 11/08/10 17:39 a-onl Start_Run_8494,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test
  • 339161 11/08/10 17:38 a-onl End_of_Run_8493
  • 339160 11/08/10 17:38 a-onl Start_Run_8493,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test
  • 339159 11/08/10 17:37 a-onl End_of_Run_8492
  • 339158 11/08/10 17:29 A. Camsonne Swapped board 0x7 with spare
  • 339157 11/08/10 16:56 pzhu ROC,RC,SCALER screen for run 8490,configuration of LeftHrs
  • 339156 11/08/10 16:55 a-onl End_of_Run_8490
  • 339155 11/08/10 16:51 Abdurahim escorted access
  • 339154 11/08/10 16:46 a-onl Start_Run_8490,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger efficiency run
  • 339153 11/08/10 16:44 pzhu ROC,RC,SCALER screen for run 8489,configuration of LeftHrs
  • 339152 11/08/10 16:41 a-onl End_of_Run_8489
  • 339150 11/08/10 16:05 W.Boeglin Shift Summary
  • 339151 11/08/10 16:04 adaq Beam Time Accounting
  • 339149 11/08/10 15:47 punjabi errors in online analysis
  • 339148 11/08/10 15:46 a-onl Start_Run_8489,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Trigger efficiency run
  • 339147 11/08/10 15:43 a-onl End_of_Run_8488
  • 339146 11/08/10 15:36 cshloka changed L1 JT valve settings
  • 339145 11/08/10 15:34 a-onl Start_Run_8488,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LD2
  • 339144 11/08/10 15:27 a-onl End_of_Run_8487
  • 339143 11/08/10 15:18 a-onl Start_Run_8487,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LD2
  • 339142 11/08/10 15:17 a-onl End_of_Run_8486
  • 339141 11/08/10 15:14 cshloka target screens snapshot
  • 339140 11/08/10 15:02 a-onl Start_Run_8486,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LD2
  • 339139 11/08/10 14:59 cshloka Target moved to Loop 1,15 cm (LD2)
  • 339138 11/08/10 14:55 a-onl End_of_Run_8485
  • 339137 11/08/10 14:55 punjabi roc,scaler,rc screens
  • 339136 11/08/10 14:31 R. Michaels updated deadtime calculation
  • 339135 11/08/10 14:28 cshloka target screens snapshot
  • 339134 11/08/10 13:58 a-onl Start_Run_8485,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339133 11/08/10 13:57 a-onl End_of_Run_8484
  • 339132 11/08/10 13:57 Punjabi roc,scaler,rc screens
  • 339131 11/08/10 12:56 a-onl Start_Run_8484,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339130 11/08/10 12:55 a-onl End_of_Run_8483
  • 339129 11/08/10 12:53 Punjabi roc,rc,scaler screens
  • 339128 11/08/10 11:56 a-onl Start_Run_8483,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339127 11/08/10 11:55 a-onl End_of_Run_8482
  • 339126 11/08/10 11:52 punjabi roc,rc,scaler screens
  • 339125 11/08/10 11:49 Boeglin Calo currents
  • 339123 11/08/10 11:01 auto_snap Radiation and Ozone snapshot
  • 339122 11/08/10 11:00 auto_snap Miscellaneous snapshot
  • 339121 11/08/10 11:00 auto_snap Hall A Tools snapshot
  • 339120 11/08/10 11:00 auto_snap Compton Polarimeter snapshot
  • 339119 11/08/10 11:00 auto_snap Beam Position and Current snapshot
  • 339118 11/08/10 10:57 a-onl Start_Run_8482,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339117 11/08/10 10:56 a-onl End_of_Run_8481
  • 339116 11/08/10 10:56 punjabi roc,rc,scaler screens
  • 339115 11/08/10 10:55 Hisham s2 efficiency
  • 339114 11/08/10 10:50 Hisham s2 efficiency
  • 339113 11/08/10 10:44 jpchen Fan response very slow, alarms ok
  • 339112 11/08/10 10:36 Hisham s1 efficiency
  • 339111 11/08/10 10:09 cshloka target screens snapshot
  • 339110 11/08/10 10:06 auto_snap Radiation and Ozone snapshot
  • 339109 11/08/10 10:06 auto_snap Miscellaneous snapshot
  • 339108 11/08/10 10:06 auto_snap Hall A Tools snapshot
  • 339107 11/08/10 10:06 auto_snap Compton Polarimeter snapshot
  • 339106 11/08/10 10:06 auto_snap Beam Position and Current snapshot
  • 339105 11/08/10 10:05 hyde S1 right[4] HV
  • 339104 11/08/10 10:01 a-onl Start_Run_8481,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339103 11/08/10 10:00 a-onl End_of_Run_8480
  • 339102 11/08/10 09:59 Punjabi roc,rc, scaler screens
  • 339101 11/08/10 08:58 a-onl Start_Run_8480,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339100 11/08/10 08:57 a-onl End_of_Run_8479
  • 339099 11/08/10 08:55 Punjabi roc screens
  • 339097 11/08/10 08:00 efuchey Shift summary for 11-08 owl
  • 339098 11/08/10 08:00 adaq Beam Time Accounting
  • 339096 11/08/10 07:59 a-onl Start_Run_8479,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339095 11/08/10 07:58 a-onl End_of_Run_8478
  • 339094 11/08/10 07:51 hkkang end-of-run
  • 339093 11/08/10 06:57 a-onl Start_Run_8478,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339092 11/08/10 06:56 a-onl End_of_Run_8477
  • 339091 11/08/10 06:53 hkkang end-of-run
  • 339090 11/08/10 06:19 hoyoung Target screens
  • 339089 11/08/10 05:57 a-onl Start_Run_8477,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339088 11/08/10 05:57 a-onl End_of_Run_8476
  • 339087 11/08/10 05:51 hkkang end-of-run
  • 339086 11/08/10 05:48 Carlos Replay can start as soon as the run starts
  • 339085 11/08/10 05:04 efuchey replay showed error messages
  • 339084 11/08/10 04:57 a-onl Start_Run_8476,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339083 11/08/10 04:56 a-onl End_of_Run_8475
  • 339082 11/08/10 04:50 hkkang end-of-run
  • 339081 11/08/10 04:09 efuchey Compton rates for 11-08-10
  • 339080 11/08/10 03:57 a-onl Start_Run_8475,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339079 11/08/10 03:57 a-onl End_of_Run_8474
  • 339078 11/08/10 03:53 hkkang end-of -run
  • 339077 11/08/10 03:00 auto_snap Radiation and Ozone snapshot
  • 339076 11/08/10 03:00 auto_snap Miscellaneous snapshot
  • 339075 11/08/10 03:00 auto_snap Hall A Tools snapshot
  • 339074 11/08/10 03:00 auto_snap Compton Polarimeter snapshot
  • 339073 11/08/10 03:00 auto_snap Beam Position and Current snapshot
  • 339072 11/08/10 02:58 efuchey PbF2 calo base currents
  • 339071 11/08/10 02:57 a-onl Start_Run_8474,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339070 11/08/10 02:56 a-onl End_of_Run_8473
  • 339069 11/08/10 02:55 efuchey Hall A tools
  • 339068 11/08/10 02:52 hkkang end-of-run
  • 339067 11/08/10 02:30 hoyoung Target screens
  • 339066 11/08/10 02:05 auto_snap Radiation and Ozone snapshot
  • 339065 11/08/10 02:05 auto_snap Miscellaneous snapshot
  • 339064 11/08/10 02:05 auto_snap Hall A Tools snapshot
  • 339063 11/08/10 02:05 auto_snap Compton Polarimeter snapshot
  • 339062 11/08/10 02:05 auto_snap Beam Position and Current snapshot
  • 339061 11/08/10 01:56 a-onl Start_Run_8473,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339060 11/08/10 01:56 efuchey beam back
  • 339059 11/08/10 01:53 a-onl End_of_Run_8472
  • 339058 11/08/10 01:45 a-onl Start_Run_8472,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2 pedestals
  • 339057 11/08/10 01:44 a-onl End_of_Run_8471
  • 339056 11/08/10 01:22 a-onl Start_Run_8471,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339055 11/08/10 01:21 a-onl End_of_Run_8470
  • 339054 11/08/10 01:13 hkkang end-of-run screenshots
  • 339053 11/08/10 00:47 efuchey QWP scan for 11-08-10
  • 339052 11/08/10 00:18 a-onl Start_Run_8470,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339051 11/08/10 00:16 a-onl End_of_Run_8469
  • 339050 11/08/10 00:00 adaq Beam Time Accounting
  • Sunday

  • 339049 11/07/10 23:48 Ramesh Subed Shift Summary
  • 339048 11/07/10 23:46 efuchey compton rates
  • 339047 11/07/10 23:46 a-onl Start_Run_8469,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339046 11/07/10 23:45 a-onl End_of_Run_8468
  • 339045 11/07/10 23:43 A. Shahinyan Roc screen
  • 339044 11/07/10 23:39 Yi Qiang Compton DAQ works normal now
  • 339043 11/07/10 23:26 Yi Qiang Cryo Target Screen Shots
  • 339042 11/07/10 23:12 Yi Qiang Compton DAQ Fails to Write Events
  • 339041 11/07/10 22:47 a-onl Start_Run_8468,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339040 11/07/10 22:46 a-onl End_of_Run_8467
  • 339039 11/07/10 22:45 A. Shahinyan Roc screen
  • 339038 11/07/10 21:53 Ramesh Trig. Eff. run 8465
  • 339037 11/07/10 21:49 a-onl Start_Run_8467,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339036 11/07/10 21:47 a-onl End_of_Run_8466
  • 339035 11/07/10 21:43 A. Shahinyan Roc screen
  • 339034 11/07/10 20:49 a-onl Start_Run_8466,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339033 11/07/10 20:47 a-onl End_of_Run_8465
  • 339032 11/07/10 20:47 A Shahinyan Roc screen
  • 339031 11/07/10 20:18 a-onl Start_Run_8465,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2, Trig eff run for half hour with LHRS config
  • 339030 11/07/10 20:16 a-onl End_of_Run_8464
  • 339029 11/07/10 20:16 A. Shahinyan Roc screen
  • 339028 11/07/10 20:02 Ramesh Calorimeter Current
  • 339027 11/07/10 19:19 a-onl Start_Run_8464,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339026 11/07/10 19:19 A. Shahinyan Roc screen
  • 339025 11/07/10 19:18 a-onl End_of_Run_8463
  • 339024 11/07/10 19:00 auto_snap Radiation and Ozone snapshot
  • 339023 11/07/10 19:00 auto_snap Miscellaneous snapshot
  • 339022 11/07/10 19:00 auto_snap Hall A Tools snapshot
  • 339021 11/07/10 19:00 auto_snap Compton Polarimeter snapshot
  • 339020 11/07/10 19:00 auto_snap Beam Position and Current snapshot
  • 339019 11/07/10 18:13 a-onl Start_Run_8463,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339018 11/07/10 18:10 a-onl End_of_Run_8462
  • 339017 11/07/10 18:10 A. Shahinyan Roc screen
  • 339016 11/07/10 18:05 auto_snap Radiation and Ozone snapshot
  • 339015 11/07/10 18:05 auto_snap Miscellaneous snapshot
  • 339014 11/07/10 18:05 auto_snap Hall A Tools snapshot
  • 339013 11/07/10 18:05 auto_snap Compton Polarimeter snapshot
  • 339012 11/07/10 18:05 auto_snap Beam Position and Current snapshot
  • 339011 11/07/10 17:16 a-onl Start_Run_8462,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339010 11/07/10 17:16 A. Shahinyna Roc
  • 339009 11/07/10 17:11 a-onl End_of_Run_8461
  • 339008 11/07/10 17:10 A. Shahinyan Roc screen
  • 339007 11/07/10 16:08 a-onl Start_Run_8461,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339006 11/07/10 16:07 a-onl End_of_Run_8460
  • 339005 11/07/10 16:07 Boeglin Shift Summary
  • 339004 11/07/10 16:04 skoirala Roc 14, 17, 18
  • 339003 11/07/10 15:20 efuchey Compton rates 11-07-10 part 3
  • 339002 11/07/10 15:15 efuchey IOC reboot
  • 339001 11/07/10 14:57 a-onl Start_Run_8460,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Kin2 4 pass LH2
  • 339000 11/07/10 14:56 a-onl End_of_Run_8459
  • 338999 11/07/10 14:56 a-onl Start_Run_8459,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS pedestal run
  • 338998 11/07/10 14:17 efuchey Target motion
  • 338997 11/07/10 14:05 a-onl End_of_Run_8456
  • 338996 11/07/10 13:41 efuchey Target motion
  • 338995 11/07/10 13:27 a-onl Start_Run_8456,Run_type=Pedestal,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS pedestal run
  • 338994 11/07/10 13:26 a-onl Start_Run_8451,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338993 11/07/10 13:24 R. Michaels fixed script to choose disk, DAQ hopefully stable now
  • 338992 11/07/10 13:06 a-onl Start_Run_8455,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338991 11/07/10 13:05 a-onl End_of_Run_8454
  • 338990 11/07/10 12:49 a-onl Start_Run_8454,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338989 11/07/10 12:46 a-onl End_of_Run_8453
  • 338988 11/07/10 12:45 Boeglin DAQ rate drop 8453
  • 338987 11/07/10 12:44 skoirala Roc 14, 17, 18
  • 338986 11/07/10 12:16 Boeglin DAQ rate dropped suddenly for run 8452
  • 338985 11/07/10 12:12 a-onl Start_Run_8453,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338984 11/07/10 12:11 a-onl End_of_Run_8452
  • 338983 11/07/10 12:06 efuchey LD2 charts
  • 338982 11/07/10 12:02 Boeglin Calo anode currents
  • 338981 11/07/10 11:48 R. Michaels implemented writing to adaql1
  • 338980 11/07/10 11:45 a-onl Start_Run_8452,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338979 11/07/10 11:45 a-onl End_of_Run_8451
  • 338978 11/07/10 11:45 a-onl End_of_Run_8444
  • 338977 11/07/10 11:44 a-onl End_of_Run_8450
  • 338976 11/07/10 11:44 a-onl Start_Run_8450,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338975 11/07/10 11:43 a-onl End_of_Run_8449
  • 338974 11/07/10 11:43 a-onl Start_Run_8449,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338973 11/07/10 11:42 a-onl End_of_Run_8448
  • 338972 11/07/10 11:41 a-onl Start_Run_8448,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338971 11/07/10 11:41 a-onl End_of_Run_8447
  • 338970 11/07/10 11:41 a-onl Start_Run_8447,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338969 11/07/10 11:40 a-onl End_of_Run_8446
  • 338968 11/07/10 11:39 efuchey Compton rates 11-07-10 part 2
  • 338967 11/07/10 11:34 a-onl Start_Run_8446,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338966 11/07/10 11:33 a-onl End_of_Run_8445
  • 338965 11/07/10 11:30 a-onl Start_Run_8445,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338964 11/07/10 11:14 a-onl Start_Run_8444,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338963 11/07/10 11:13 a-onl End_of_Run_8443
  • 338962 11/07/10 11:11 skoirala Roc 14, 17, 18
  • 338961 11/07/10 11:00 auto_snap Radiation and Ozone snapshot
  • 338960 11/07/10 11:00 auto_snap Miscellaneous snapshot
  • 338959 11/07/10 11:00 auto_snap Hall A Tools snapshot
  • 338958 11/07/10 10:59 auto_snap Compton Polarimeter snapshot
  • 338957 11/07/10 10:59 auto_snap Beam Position and Current snapshot
  • 338956 11/07/10 10:15 a-onl Start_Run_8443,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338955 11/07/10 10:14 a-onl End_of_Run_8442
  • 338954 11/07/10 10:13 skoirala End of Run Screen Shots
  • 338953 11/07/10 10:08 A. Camsonne If you need a bit of time to let files to go to silo switch to LH2 to reduce rate
  • 338952 11/07/10 10:05 auto_snap Radiation and Ozone snapshot
  • 338951 11/07/10 10:05 auto_snap Miscellaneous snapshot
  • 338950 11/07/10 10:05 auto_snap Hall A Tools snapshot
  • 338949 11/07/10 10:05 auto_snap Compton Polarimeter snapshot
  • 338948 11/07/10 10:05 auto_snap Beam Position and Current snapshot
  • 338947 11/07/10 09:12 a-onl Start_Run_8442,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338946 11/07/10 09:10 a-onl End_of_Run_8441
  • 338945 11/07/10 09:06 skoirala End of Run's Screen Shots
  • 338944 11/07/10 08:52 efuchey rates for 11-07
  • 338943 11/07/10 08:26 efuchey LD2 charts
  • 338942 11/07/10 08:25 efuchey target tools screenshot
  • 338941 11/07/10 08:17 R. Michaels re: disk usage by DAQ
  • 338940 11/07/10 08:12 adaq rates too high
  • 338939 11/07/10 08:01 Ole Hansen Owl Shift Summary
  • 338938 11/07/10 07:58 a-onl Start_Run_8441,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338937 11/07/10 07:55 a-onl End_of_Run_8440
  • 338936 11/07/10 07:53 hkkang end-of-run
  • 338935 11/07/10 07:18 Ole Hansen ABU/BANU numbers reported
  • 338934 11/07/10 07:15 Ole Hansen data disk copying
  • 338933 11/07/10 06:56 a-onl Start_Run_8440,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338932 11/07/10 06:55 a-onl End_of_Run_8439
  • 338931 11/07/10 06:53 Ole Hansen Calo ADCs
  • 338930 11/07/10 06:41 benesch Pockel Cell +HV Alarm
  • 338929 11/07/10 05:56 a-onl Start_Run_8439,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338928 11/07/10 05:55 a-onl End_of_Run_8438
  • 338927 11/07/10 05:52 hkkang end-of-run screenshots
  • 338926 11/07/10 04:54 a-onl Start_Run_8438,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338925 11/07/10 04:53 a-onl End_of_Run_8437
  • 338924 11/07/10 04:49 hkkang end-of-run
  • 338923 11/07/10 03:53 a-onl Start_Run_8437,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338922 11/07/10 03:53 a-onl End_of_Run_8436
  • 338921 11/07/10 03:47 hkkang end-of-run
  • 338920 11/07/10 02:59 auto_snap Radiation and Ozone snapshot
  • 338919 11/07/10 02:59 auto_snap Miscellaneous snapshot
  • 338918 11/07/10 02:59 auto_snap Hall A Tools snapshot
  • 338917 11/07/10 02:59 auto_snap Compton Polarimeter snapshot
  • 338916 11/07/10 02:59 auto_snap Beam Position and Current snapshot
  • 338915 11/07/10 02:52 a-onl Start_Run_8436,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338914 11/07/10 02:52 a-onl End_of_Run_8435
  • 338913 11/07/10 02:50 hkkang end-of-run screenshots
  • 338912 11/07/10 02:49 Ole Hansen Run 8434 is 55 minutes long
  • 338911 11/07/10 02:05 auto_snap Radiation and Ozone snapshot
  • 338910 11/07/10 02:05 auto_snap Miscellaneous snapshot
  • 338909 11/07/10 02:04 auto_snap Hall A Tools snapshot
  • 338908 11/07/10 02:04 auto_snap Compton Polarimeter snapshot
  • 338907 11/07/10 02:04 auto_snap Beam Position and Current snapshot
  • 338906 11/07/10 01:52 a-onl Start_Run_8435,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338902 11/07/10 01:49 a-onl Start_Run_8434,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338901 11/07/10 01:48 a-onl End_of_Run_8433
  • 338905 11/07/10 01:45 a-onl End_of_Run_8434
  • 338904 11/07/10 01:45 Ole Hansen CODA crash
  • 338900 11/07/10 01:38 hkkang end-of-run screenshots
  • 338903 11/07/10 01:07 Ole Hansen Disks almost full
  • 338899 11/07/10 00:47 a-onl Start_Run_8433,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338898 11/07/10 00:45 a-onl End_of_Run_8432
  • 338897 11/07/10 00:45 adaq end-of-run screenshots
  • 338896 11/07/10 00:00 adaq Beam Time Accounting
  • Saturday

  • 338895 11/06/10 23:53 Ramesh Shift Summary
  • 338894 11/06/10 23:45 a-onl Start_Run_8432,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338893 11/06/10 23:45 a-onl End_of_Run_8431
  • 338892 11/06/10 23:44 A. Shahinyan roc.run control
  • 338891 11/06/10 23:41 Carlos "ROCs Sync check" replay plots
  • 338890 11/06/10 22:50 A. S Roc,runcotrol
  • 338889 11/06/10 22:50 a-onl Start_Run_8431,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338888 11/06/10 22:48 a-onl End_of_Run_8430
  • 338887 11/06/10 22:17 Ramesh Calorimeter current
  • 338886 11/06/10 22:10 Abdurahim Target screens
  • 338885 11/06/10 21:57 a-onl Start_Run_8430,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338884 11/06/10 21:56 a-onl End_of_Run_8429
  • 338883 11/06/10 21:55 a-onl Start_Run_8429,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338882 11/06/10 21:53 a-onl End_of_Run_8428
  • 338881 11/06/10 21:50 a-onl Start_Run_8428,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338880 11/06/10 21:48 Ramesh 100% deadtime, aborting and restarting
  • 338879 11/06/10 21:47 a-onl Start_Run_8427,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338878 11/06/10 21:32 A, Camsonne Dropped VME adc from ROC17
  • 338877 11/06/10 21:29 A. Shahinyan Roc17 problem
  • 338876 11/06/10 21:27 a-onl Start_Run_8426,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338875 11/06/10 20:52 a-onl Start_Run_8425,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338874 11/06/10 20:51 a-onl End_of_Run_8424
  • 338873 11/06/10 20:51 A. Shahinyan Roc screen
  • 338872 11/06/10 20:06 Abdurahim Target screen shots
  • 338871 11/06/10 19:59 auto_snap Radiation and Ozone snapshot
  • 338870 11/06/10 19:59 auto_snap Miscellaneous snapshot
  • 338869 11/06/10 19:59 auto_snap Hall A Tools snapshot
  • 338868 11/06/10 19:59 auto_snap Compton Polarimeter snapshot
  • 338867 11/06/10 19:59 auto_snap Beam Position and Current snapshot
  • 338866 11/06/10 19:56 Ramesh Coda reboot
  • 338865 11/06/10 19:52 a-onl Start_Run_8424,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338864 11/06/10 19:46 A. Shahinyan 100% deadtime
  • 338863 11/06/10 19:44 a-onl End_of_Run_8422
  • 338862 11/06/10 19:15 Ole Hansen Re: error from extract_runend_info.sh (fixed)
  • 338861 11/06/10 19:13 a-onl Start_Run_8422,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338860 11/06/10 19:12 a-onl End_of_Run_8421
  • 338859 11/06/10 19:12 A. Shahinyan Roc screen
  • 338858 11/06/10 19:04 auto_snap Radiation and Ozone snapshot
  • 338857 11/06/10 19:04 auto_snap Miscellaneous snapshot
  • 338856 11/06/10 19:04 auto_snap Hall A Tools snapshot
  • 338855 11/06/10 19:04 auto_snap Compton Polarimeter snapshot
  • 338854 11/06/10 19:04 auto_snap Beam Position and Current snapshot
  • 338853 11/06/10 18:10 a-onl Start_Run_8421,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338852 11/06/10 18:09 a-onl End_of_Run_8420
  • 338851 11/06/10 18:05 A. Shahinyan Roc scrren
  • 338850 11/06/10 17:52 Abdurahim Target screen shots
  • 338849 11/06/10 17:48 efuchey rates check 11-06-10
  • 338848 11/06/10 17:11 A. Shahinyan Roc, coda,scaler
  • 338847 11/06/10 17:09 a-onl Start_Run_8420,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338846 11/06/10 16:51 Abdurahim Compton EPICS has been rebooted
  • 338845 11/06/10 16:47 Carlos LD2 target rates
  • 338844 11/06/10 16:30 a-onl Start_Run_8419,Run_type=Other,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target, Trig eff check run
  • 338843 11/06/10 16:27 a-onl End_of_Run_8418
  • 338842 11/06/10 16:26 A.shahinyan Roc screen
  • 338841 11/06/10 16:00 adaq Beam Time Accounting
  • 338840 11/06/10 15:58 ent day shift summary
  • 338839 11/06/10 15:46 Punjabi roc, rc, scaler screens
  • 338838 11/06/10 15:28 horn Target: status
  • 338837 11/06/10 15:25 horn Loop 2 fan setting to 10% (24Hz)
  • 338836 11/06/10 15:24 a-onl Start_Run_8418,Run_type=Production,target_type=LD2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam, LD2 target
  • 338835 11/06/10 15:23 horn Target: moved to 15-cm LD2
  • 338834 11/06/10 15:13 ent change to LD2 started
  • 338833 11/06/10 15:09 a-onl End_of_Run_8417
  • 338832 11/06/10 14:42 adaq compton chart
  • 338831 11/06/10 14:24 a-onl Start_Run_8417,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam
  • 338830 11/06/10 14:23 a-onl End_of_Run_8416
  • 338829 11/06/10 14:03 A. Camsonne Reduced pedestal threshold for Gas Cerenkov to 5 channel above pedestal
  • 338828 11/06/10 13:59 a-onl Start_Run_8416,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam
  • 338827 11/06/10 13:33 a-onl Start_Run_8415,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam
  • 338826 11/06/10 13:32 a-onl End_of_Run_8414
  • 338825 11/06/10 13:27 punjabi roc,scaler,rc screens
  • 338824 11/06/10 12:35 R. Michaels run 8413, pedestal run
  • 338823 11/06/10 12:32 a-onl Start_Run_8414,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=kinematic 2 with 4-pass beam
  • 338822 11/06/10 12:31 a-onl Start_Run_8413,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=pedestal run
  • 338821 11/06/10 12:29 a-onl End_of_Run_8412
  • 338820 11/06/10 12:24 Punjabi roc, scaler, rc screens
  • 338819 11/06/10 11:59 auto_snap Radiation and Ozone snapshot
  • 338818 11/06/10 11:59 auto_snap Miscellaneous snapshot
  • 338817 11/06/10 11:59 auto_snap Hall A Tools snapshot
  • 338816 11/06/10 11:59 auto_snap Compton Polarimeter snapshot
  • 338815 11/06/10 11:59 auto_snap Beam Position and Current snapshot
  • 338814 11/06/10 11:28 a-onl Start_Run_8412,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338813 11/06/10 11:27 a-onl End_of_Run_8411
  • 338812 11/06/10 11:27 Punjabi roc, rc, scaler screens for run 8411
  • 338811 11/06/10 11:19 ent ./extract_runend_info.sh: line 17: [: -ge: unary operator expected
  • 338810 11/06/10 11:04 auto_snap Radiation and Ozone snapshot
  • 338809 11/06/10 11:04 auto_snap Miscellaneous snapshot
  • 338808 11/06/10 11:04 auto_snap Hall A Tools snapshot
  • 338807 11/06/10 11:04 auto_snap Compton Polarimeter snapshot
  • 338806 11/06/10 11:04 auto_snap Beam Position and Current snapshot
  • 338805 11/06/10 10:53 ent lost coda daq - start again
  • 338804 11/06/10 10:28 horn Target: status
  • 338803 11/06/10 10:27 a-onl Start_Run_8411,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338802 11/06/10 10:22 a-onl End_of_Run_8410
  • 338801 11/06/10 09:49 punjabi missing start of the run window
  • 338800 11/06/10 09:46 a-onl Start_Run_8410,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338799 11/06/10 09:44 Punjabi roc screens for run 8409
  • 338798 11/06/10 09:42 ent run 8409 ROC17 again
  • 338797 11/06/10 09:40 a-onl End_of_Run_8409
  • 338796 11/06/10 09:07 ent status of run 8408 unclear
  • 338795 11/06/10 08:42 a-onl Start_Run_8409,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338794 11/06/10 08:37 ent adaql2 self-reboot
  • 338793 11/06/10 08:22 adaq Beam Time Accounting
  • 338792 11/06/10 08:17 Ole Hansen Owl Shift Summary
  • 338791 11/06/10 08:06 a-onl Start_Run_8408,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338790 11/06/10 07:59 hkkang end-of-run screenshot
  • 338789 11/06/10 07:56 a-onl End_of_Run_8406
  • 338788 11/06/10 07:17 a-onl Start_Run_8406,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338787 11/06/10 07:17 a-onl End_of_Run_8405
  • 338786 11/06/10 07:16 hkkang end-of-run
  • 338785 11/06/10 06:16 a-onl Start_Run_8405,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338784 11/06/10 06:14 a-onl End_of_Run_8404
  • 338783 11/06/10 06:09 hkkang end-of-run
  • 338782 11/06/10 05:14 a-onl Start_Run_8404,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338781 11/06/10 05:14 a-onl End_of_Run_8403
  • 338780 11/06/10 05:12 hkkang end-of-run
  • 338779 11/06/10 05:10 hoyoung Target screens
  • 338778 11/06/10 04:52 Ole Hansen Compton background high
  • 338777 11/06/10 04:25 Ole Hansen Calorimeter ADC readbacks
  • 338776 11/06/10 04:14 a-onl Start_Run_8403,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338775 11/06/10 04:14 Ole Hansen Resuming production
  • 338774 11/06/10 04:12 Ole Hansen Compton HV on
  • 338773 11/06/10 04:10 hoyoung 15cm LH2 target
  • 338772 11/06/10 04:04 Ole Hansen Moving to LH2 target
  • 338771 11/06/10 03:58 auto_snap Radiation and Ozone snapshot
  • 338770 11/06/10 03:58 auto_snap Miscellaneous snapshot
  • 338769 11/06/10 03:58 auto_snap Hall A Tools snapshot
  • 338768 11/06/10 03:58 auto_snap Compton Polarimeter snapshot
  • 338767 11/06/10 03:58 auto_snap Beam Position and Current snapshot
  • 338766 11/06/10 03:51 Ole Hansen Beam is back
  • 338765 11/06/10 03:44 Ole Hansen Target to empty
  • 338764 11/06/10 03:39 hoyoung Target moved to empty
  • 338763 11/06/10 03:04 auto_snap Radiation and Ozone snapshot
  • 338762 11/06/10 03:04 auto_snap Miscellaneous snapshot
  • 338761 11/06/10 03:04 auto_snap Hall A Tools snapshot
  • 338760 11/06/10 03:03 auto_snap Compton Polarimeter snapshot
  • 338759 11/06/10 03:03 auto_snap Beam Position and Current snapshot
  • 338758 11/06/10 01:43 Ole Hansen Compton HV off
  • 338757 11/06/10 01:42 Ole Hansen Beam update (1+ more hours)
  • 338756 11/06/10 01:26 jroche new BCM calibration in the end of run script
  • 338755 11/06/10 01:25 hkkang end-of-run screenshots
  • 338754 11/06/10 01:20 Ole Hansen Injector trip, no beam
  • 338753 11/06/10 01:20 a-onl End_of_Run_8402
  • 338752 11/06/10 01:07 a-onl Start_Run_8402,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338751 11/06/10 01:06 a-onl End_of_Run_8401
  • 338750 11/06/10 01:05 hkkang end-of-run
  • 338749 11/06/10 01:05 adaq end-of-run summary
  • 338748 11/06/10 00:47 Ole Hansen Compton DAQ failure (fixed)
  • 338747 11/06/10 00:28 hoyoung Target screens
  • 338746 11/06/10 00:14 adaq Beam Time Accounting
  • 338745 11/06/10 00:06 R. Michaels TS11 crash, attempted fix
  • 338744 11/06/10 00:04 a-onl Start_Run_8401,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • Friday

  • 338743 11/05/10 23:53 ent swing shift summary
  • 338742 11/05/10 23:45 a-onl Start_Run_8399,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338741 11/05/10 23:34 a-onl End_of_Run_8398
  • 338740 11/05/10 23:31 A. Shahinyan TS 11 problem
  • 338739 11/05/10 23:18 a-onl Start_Run_8398,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338738 11/05/10 23:16 a-onl End_of_Run_8397
  • 338737 11/05/10 23:02 Carlos End-of-run screenshots
  • 338736 11/05/10 23:01 a-onl Start_Run_8397,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338735 11/05/10 22:59 A. Camsonne TS11 rebboted during run 8396 error message
  • 338734 11/05/10 22:16 a-onl Start_Run_8396,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338733 11/05/10 22:15 a-onl End_of_Run_8395
  • 338732 11/05/10 21:58 ent Compton vertical scan
  • 338731 11/05/10 21:33 A. Camsonne Compton vertical scan during run 8395
  • 338730 11/05/10 21:16 a-onl Start_Run_8395,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338729 11/05/10 21:15 a-onl End_of_Run_8394
  • 338728 11/05/10 21:11 ent Compton vertical scan during run 8394
  • 338727 11/05/10 21:05 Carlos Kinematic change: KIN2-4pass
  • 338726 11/05/10 20:31 a-onl Start_Run_8394,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338725 11/05/10 20:23 A. Camsonne Gate for ROC4 sync ADC restored
  • 338724 11/05/10 20:15 a-onl End_of_Run_8393
  • 338723 11/05/10 20:12 a-onl Start_Run_8393,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338722 11/05/10 20:11 a-onl End_of_Run_8392
  • 338721 11/05/10 20:02 a-onl Start_Run_8392,Run_type=Cosmics,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338720 11/05/10 19:58 auto_snap Radiation and Ozone snapshot
  • 338719 11/05/10 19:58 auto_snap Miscellaneous snapshot
  • 338718 11/05/10 19:58 auto_snap Hall A Tools snapshot
  • 338717 11/05/10 19:58 auto_snap Compton Polarimeter snapshot
  • 338716 11/05/10 19:58 auto_snap Beam Position and Current snapshot
  • 338715 11/05/10 19:44 a-onl End_of_Run_8391
  • 338714 11/05/10 19:35 a-onl Start_Run_8391,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.864,leftarm_theta=26.55,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Kin II 4 pass
  • 338713 11/05/10 19:34 a-onl End_of_Run_8390
  • 338712 11/05/10 19:33 a-onl Start_Run_8390,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Rate check LH2
  • 338711 11/05/10 19:29 horn Target: moved to 15-cm LH2
  • 338710 11/05/10 19:03 auto_snap Radiation and Ozone snapshot
  • 338709 11/05/10 19:03 auto_snap Miscellaneous snapshot
  • 338708 11/05/10 19:03 auto_snap Hall A Tools snapshot
  • 338707 11/05/10 19:03 auto_snap Compton Polarimeter snapshot
  • 338706 11/05/10 19:03 auto_snap Beam Position and Current snapshot
  • 338705 11/05/10 18:53 horn Target: moved to empty position
  • 338704 11/05/10 18:40 a-onl Start_Run_8389,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338703 11/05/10 18:40 a-onl Start_Run_8388,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338701 11/05/10 18:34 a-onl End_of_Run_8387
  • 338700 11/05/10 18:33 a-onl End_of_Run_8388
  • 338699 11/05/10 18:33 a-onl End_of_Run_8389
  • 338698 11/05/10 18:20 A. Camsonne Scope busy
  • 338697 11/05/10 17:48 horn Target: status
  • 338696 11/05/10 17:36 jpchen Hall A target cooldown completed
  • 338695 11/05/10 17:20 a-onl Start_Run_8386,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338694 11/05/10 17:18 a-onl End_of_Run_8385
  • 338693 11/05/10 17:16 a-onl Start_Run_8385,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338692 11/05/10 17:09 a-onl End_of_Run_8384
  • 338691 11/05/10 17:08 R. Michaels double pulsing on LNE solved
  • 338690 11/05/10 17:07 a-onl Start_Run_8384,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338689 11/05/10 17:05 a-onl End_of_Run_8383
  • 338688 11/05/10 17:04 a-onl Start_Run_8383,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes - no write in file
  • 338687 11/05/10 17:00 a-onl Start_Run_8382,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott - et size change to 1000 events of 100 Kbytes
  • 338686 11/05/10 16:53 a-onl End_of_Run_8381
  • 338685 11/05/10 16:48 jroche new calibration for BCM (epics values)
  • 338684 11/05/10 16:44 a-onl Start_Run_8381,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338683 11/05/10 16:39 a-onl End_of_Run_8380
  • 338682 11/05/10 16:22 jpchen LH2 fan back to 46 Hz
  • 338681 11/05/10 16:12 adaq Beam Time Accounting
  • 338680 11/05/10 16:05 R. Michaels gate for helicity scaler restored
  • 338679 11/05/10 16:03 jpchen Loop 1 fan back to 24 Hz
  • 338678 11/05/10 16:03 jpchen LH2 is at 19K, cooling down D2 now
  • 338677 11/05/10 16:01 jpchen Hall C target coolodnw completed, Hall A cooldown in progress
  • 338676 11/05/10 15:58 jpchen 15K supply and target vacuum
  • 338674 11/05/10 15:57 hyde Day Shift Summary
  • 338673 11/05/10 15:56 a-onl Start_Run_8380,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338672 11/05/10 15:34 a-onl Start_Run_8379,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338671 11/05/10 15:31 a-onl Start_Run_8378,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338670 11/05/10 15:15 R. Michaels hall C tgt update, leaky He4 bayonet
  • 338669 11/05/10 14:50 R. Michaels helicity issues, cont.
  • 338668 11/05/10 14:43 a-onl Start_Run_8376,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338667 11/05/10 14:42 a-onl End_of_Run_8375
  • 338666 11/05/10 14:36 a-onl Start_Run_8375,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Checking dead time with D.Abbott
  • 338665 11/05/10 14:24 a-onl End_of_Run_8374
  • 338664 11/05/10 14:23 jroche BCM calibration results
  • 338663 11/05/10 14:23 a-onl Start_Run_8374,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338662 11/05/10 14:21 a-onl End_of_Run_8373
  • 338661 11/05/10 14:19 a-onl Start_Run_8373,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338660 11/05/10 14:13 a-onl End_of_Run_8372
  • 338659 11/05/10 14:12 a-onl Start_Run_8372,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338658 11/05/10 13:43 a-onl End_of_Run_8371
  • 338657 11/05/10 13:42 a-onl Start_Run_8371,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338656 11/05/10 13:25 a-onl Start_Run_8368,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff writing on
  • 338655 11/05/10 13:15 R. Michaels target runs away ... bad vacuum
  • 338654 11/05/10 13:08 hyde target valves closed
  • 338653 11/05/10 12:54 a-onl Start_Run_8367,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338652 11/05/10 12:53 a-onl End_of_Run_8367
  • 338651 11/05/10 12:49 a-onl Start_Run_8365,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338650 11/05/10 12:44 a-onl End_of_Run_8366
  • 338649 11/05/10 12:44 a-onl End_of_Run_8365
  • 338648 11/05/10 12:42 a-onl Start_Run_8361,Run_type=Pedestal,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=HRS pedestal run
  • 338647 11/05/10 12:37 a-onl End_of_Run_8364
  • 338646 11/05/10 12:35 a-onl Start_Run_8364,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics TrigDVCSBuff
  • 338645 11/05/10 12:34 a-onl End_of_Run_8363
  • 338644 11/05/10 12:33 jpchen close loop 1 J-T valve
  • 338643 11/05/10 12:33 a-onl Start_Run_8363,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics ROC1417Buff
  • 338642 11/05/10 12:12 a-onl End_of_Run_8362
  • 338641 11/05/10 12:10 hyde Compton HV off
  • 338640 11/05/10 12:10 a-onl Start_Run_8362,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test random cosmics
  • 338639 11/05/10 12:02 a-onl End_of_Run_8361
  • 338638 11/05/10 12:02 a-onl End_of_Run_8356
  • 338637 11/05/10 12:01 a-onl End_of_Run_8360
  • 338636 11/05/10 12:00 hyde Calo cosmic HV on
  • 338635 11/05/10 11:59 a-onl Start_Run_8360,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser TrigDVCSBuff config - cosmics paddles on
  • 338633 11/05/10 11:58 a-onl End_of_Run_8359
  • 338634 11/05/10 11:58 auto_snap Radiation and Ozone snapshot
  • 338632 11/05/10 11:58 auto_snap Miscellaneous snapshot
  • 338631 11/05/10 11:58 auto_snap Hall A Tools snapshot
  • 338630 11/05/10 11:58 auto_snap Compton Polarimeter snapshot
  • 338629 11/05/10 11:58 auto_snap Beam Position and Current snapshot
  • 338628 11/05/10 11:56 a-onl Start_Run_8359,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser ROC1417buff config - cosmics paddles on
  • 338627 11/05/10 11:55 A. Camsonne Calorimeter HV on for random rate deadtime check
  • 338626 11/05/10 11:54 a-onl End_of_Run_8358
  • 338625 11/05/10 11:53 a-onl Start_Run_8358,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser ROC1417buff config - cosmics paddles on
  • 338624 11/05/10 11:48 a-onl End_of_Run_8357
  • 338623 11/05/10 11:46 a-onl Start_Run_8357,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser ROC1417buff config
  • 338622 11/05/10 11:45 a-onl Start_Run_8356,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser ROC1417buff config
  • 338621 11/05/10 11:43 a-onl End_of_Run_8355
  • 338620 11/05/10 11:42 a-onl End_of_Run_8354
  • 338619 11/05/10 11:41 a-onl Start_Run_8354,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser DVCSBuff config
  • 338618 11/05/10 11:40 A. Camsonne DAQ deadtime at 300 Hz
  • 338617 11/05/10 11:40 a-onl End_of_Run_8353
  • 338616 11/05/10 11:35 A. Camsonne Dead time at 600 Hz pulser
  • 338615 11/05/10 11:33 a-onl Start_Run_8353,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser
  • 338614 11/05/10 11:32 jroche BcmLog_165
  • 338613 11/05/10 11:25 a-onl Start_Run_8352,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser
  • 338612 11/05/10 11:23 a-onl Start_Run_8351,Run_type=Other,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Deadtime DAQ tests with pulser
  • 338611 11/05/10 11:22 a-onl End_of_Run_8343
  • 338610 11/05/10 11:04 hyde Heating from optics target
  • 338609 11/05/10 11:03 auto_snap Radiation and Ozone snapshot
  • 338608 11/05/10 11:03 auto_snap Miscellaneous snapshot
  • 338607 11/05/10 11:03 auto_snap Hall A Tools snapshot
  • 338606 11/05/10 11:03 auto_snap Compton Polarimeter snapshot
  • 338605 11/05/10 11:03 auto_snap Beam Position and Current snapshot
  • 338604 11/05/10 11:02 a-onl End_of_Run_8350
  • 338603 11/05/10 11:01 hyde Hall A Tools screen
  • 338602 11/05/10 10:58 hyde DVCS Calo DC
  • 338601 11/05/10 10:45 hyde spot on target
  • 338600 11/05/10 10:41 a-onl Start_Run_8350,Run_type=Calibration,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Optics - Kin3 4 pass
  • 338599 11/05/10 10:39 R. Michaels target to optics
  • 338598 11/05/10 10:38 hyde BCM Calibration stripchart
  • 338597 11/05/10 10:34 a-onl End_of_Run_8349
  • 338596 11/05/10 10:12 hyde Left HRS collimator
  • 338595 11/05/10 10:03 a-onl Start_Run_8349,Run_type=Calibration,target_type=Empty loop,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Empty target BCM calibration! - Kin3 4 pass
  • 338594 11/05/10 10:02 a-onl End_of_Run_8348
  • 338593 11/05/10 09:51 a-onl Start_Run_8348,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Empty target BCM calibration! - Kin3 4 pass
  • 338592 11/05/10 09:38 a-onl End_of_Run_8347
  • 338591 11/05/10 09:32 a-onl Start_Run_8347,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Empty target BCM calibration? - Kin3 4 pass
  • 338590 11/05/10 09:30 a-onl End_of_Run_8346
  • 338589 11/05/10 09:07 a-onl Start_Run_8346,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Empty target BCM calibration - Kin3 4 pass
  • 338588 11/05/10 09:01 R. Michaels move to empty target
  • 338587 11/05/10 09:00 a-onl End_of_Run_8345
  • 338586 11/05/10 08:36 R. Michaels status of HP heater
  • 338585 11/05/10 08:32 hyde DVCS Calo HV off
  • 338584 11/05/10 08:28 hyde Compton HV on
  • 338583 11/05/10 08:27 hyde Calo DC currents
  • 338582 11/05/10 08:25 a-onl Start_Run_8345,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Optics - Kin3 4 pass
  • 338581 11/05/10 08:22 R. Michaels cryo closes 15K supply to Hall A
  • 338580 11/05/10 08:15 R. Michaels moving to optics target
  • 338579 11/05/10 08:14 R. Michaels hall C target cooldown, hall A warming
  • 338578 11/05/10 08:04 adaq Beam Time Accounting
  • 338577 11/05/10 08:04 Ole Hansen Owl Shift Summary
  • 338576 11/05/10 08:00 hyde Compton HV off
  • 338575 11/05/10 07:51 Ole Hansen Beam off for Hall C cooldown
  • 338574 11/05/10 07:50 cshloka roc,coda,scalers for run 8344
  • 338573 11/05/10 07:49 lselvy Empty Target
  • 338572 11/05/10 07:49 lselvy Beam Off
  • 338571 11/05/10 07:48 a-onl End_of_Run_8344
  • 338570 11/05/10 07:33 a-onl Start_Run_8344,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338569 11/05/10 07:31 cshloka roc,coda,scalers for run 8343
  • 338568 11/05/10 07:16 lselvy Target Update
  • 338567 11/05/10 07:08 Ole Hansen No word from Hall C yet
  • 338566 11/05/10 06:33 a-onl Start_Run_8343,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338565 11/05/10 06:32 a-onl End_of_Run_8342
  • 338564 11/05/10 06:31 cshloka roc,coda,scalrs for run 8342
  • 338563 11/05/10 05:32 a-onl Start_Run_8342,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338562 11/05/10 05:31 a-onl End_of_Run_8341
  • 338561 11/05/10 05:31 cshloka roc,coda,scalers for run 8341
  • 338560 11/05/10 04:26 a-onl Start_Run_8341,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338559 11/05/10 04:25 a-onl End_of_Run_8340
  • 338558 11/05/10 04:25 cshloka roc, coda, scalers at the end of run 8340
  • 338557 11/05/10 03:58 auto_snap Radiation and Ozone snapshot
  • 338556 11/05/10 03:58 auto_snap Miscellaneous snapshot
  • 338555 11/05/10 03:57 auto_snap Hall A Tools snapshot
  • 338554 11/05/10 03:57 auto_snap Compton Polarimeter snapshot
  • 338553 11/05/10 03:57 auto_snap Beam Position and Current snapshot
  • 338552 11/05/10 03:22 a-onl Start_Run_8340,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338551 11/05/10 03:20 a-onl End_of_Run_8339
  • 338550 11/05/10 03:20 cshloka roc,coda,scalers for un 8339
  • 338549 11/05/10 03:03 auto_snap Radiation and Ozone snapshot
  • 338548 11/05/10 03:03 auto_snap Miscellaneous snapshot
  • 338547 11/05/10 03:03 auto_snap Hall A Tools snapshot
  • 338546 11/05/10 03:03 auto_snap Compton Polarimeter snapshot
  • 338545 11/05/10 03:03 auto_snap Beam Position and Current snapshot
  • 338544 11/05/10 02:05 Ole Hansen Script to extract run summary data
  • 338543 11/05/10 02:04 Ole Hansen Script to extract run summary data
  • 338542 11/05/10 01:40 a-onl Start_Run_8339,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338541 11/05/10 01:38 A. Camsonne ROC17 crashed
  • 338540 11/05/10 01:36 a-onl End_of_Run_8338
  • 338539 11/05/10 01:11 Ole Hansen Beam was off for 30 minutes
  • 338538 11/05/10 00:06 LeRose Shift Summary (swing)
  • 338537 11/05/10 00:03 adaq Beam Time Accounting
  • Thursday

  • 338536 11/04/10 23:57 a-onl Start_Run_8338,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338535 11/04/10 23:47 a-onl End_of_Run_8337
  • 338534 11/04/10 23:46 Carlos BCM calibration
  • 338533 11/04/10 23:19 a-onl Start_Run_8337,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338532 11/04/10 23:17 A. Camsonne DVCS DAQ deadtime at 400 Hz
  • 338531 11/04/10 23:13 kelleher Shift Leaders -- Please add to shift summaries
  • 338530 11/04/10 23:02 a-onl Start_Run_8336,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338529 11/04/10 22:54 a-onl End_of_Run_8335
  • 338528 11/04/10 22:28 a-onl Start_Run_8335,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm - Kin3 4 pass
  • 338527 11/04/10 22:23 a-onl End_of_Run_8334
  • 338526 11/04/10 22:11 a-onl Start_Run_8334,Run_type=Production,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production Target LH2 15cm
  • 338525 11/04/10 22:08 LeRose Back to Production Running
  • 338524 11/04/10 22:08 A. Camsonne DVCS clock set to 11 Hz for production data taking
  • 338523 11/04/10 22:08 horn Target: fan setting to 46Hz
  • 338522 11/04/10 22:07 horn Target: moved to 15cm LH2
  • 338521 11/04/10 22:00 A. Camsonne HWP mystery solved - now should be fixed
  • 338520 11/04/10 21:35 A. Camsonne Today DVCS DAQ work
  • 338519 11/04/10 21:32 LeRose Beam Back
  • 338518 11/04/10 21:31 Carlos Block 91
  • 338517 11/04/10 20:43 A. Camsonne Pedestal run HV off 8333
  • 338516 11/04/10 20:42 a-onl End_of_Run_8333
  • 338515 11/04/10 20:42 a-onl Start_Run_8333,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run TrigDVCSBuff HV Off
  • 338514 11/04/10 20:39 a-onl End_of_Run_8332
  • 338513 11/04/10 20:39 a-onl Start_Run_8332,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run TrigDVCSBuff
  • 338512 11/04/10 20:34 a-onl End_of_Run_8331
  • 338511 11/04/10 20:33 a-onl Start_Run_8331,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run ROC1417Buff
  • 338510 11/04/10 20:26 A. Camsonne ARS pedestal run 8330
  • 338509 11/04/10 20:26 LeRose Controlled Access
  • 338508 11/04/10 20:25 a-onl Start_Run_8327,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter block 0
  • 338507 11/04/10 20:24 a-onl Start_Run_8330,Run_type=Pedestal,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Pedestal run
  • 338506 11/04/10 20:05 a-onl End_of_Run_8329
  • 338505 11/04/10 20:05 A. Camsonne Calorimeter LED scan Run 8328
  • 338504 11/04/10 20:03 a-onl Start_Run_8329,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test TrigDVCSBuff
  • 338503 11/04/10 20:02 a-onl End_of_Run_8328
  • 338502 11/04/10 19:57 auto_snap Radiation and Ozone snapshot
  • 338501 11/04/10 19:57 auto_snap Miscellaneous snapshot
  • 338500 11/04/10 19:57 auto_snap Hall A Tools snapshot
  • 338499 11/04/10 19:57 auto_snap Compton Polarimeter snapshot
  • 338498 11/04/10 19:57 auto_snap Beam Position and Current snapshot
  • 338497 11/04/10 19:45 a-onl Start_Run_8328,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter LED scan delay Led det to 0x69
  • 338496 11/04/10 19:42 a-onl End_of_Run_8327
  • 338495 11/04/10 19:36 a-onl End_of_Run_8326
  • 338494 11/04/10 19:35 a-onl Start_Run_8323,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter scan test
  • 338493 11/04/10 19:35 a-onl Start_Run_8326,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter block 0
  • 338492 11/04/10 19:34 a-onl End_of_Run_8325
  • 338491 11/04/10 19:30 a-onl Start_Run_8325,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter scan
  • 338490 11/04/10 19:29 a-onl End_of_Run_8324
  • 338489 11/04/10 19:29 a-onl Start_Run_8324,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter scan test
  • 338488 11/04/10 19:28 a-onl End_of_Run_8323
  • 338487 11/04/10 19:26 a-onl End_of_Run_8322
  • 338486 11/04/10 19:25 a-onl Start_Run_8322,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter scan test
  • 338485 11/04/10 19:25 a-onl End_of_Run_8321
  • 338484 11/04/10 19:24 A. Camsonne Led pulser frequency changed for LED scan
  • 338483 11/04/10 19:19 a-onl Start_Run_8321,Run_type=LED,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Calorimeter scan test
  • 338482 11/04/10 19:09 LeRose Called for a sweep of the Hall
  • 338481 11/04/10 19:04 a-onl End_of_Run_8320
  • 338480 11/04/10 19:04 a-onl Start_Run_8320,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS DAQ checkout E14 ROC14 and 17 with buffering
  • 338479 11/04/10 19:02 auto_snap Radiation and Ozone snapshot
  • 338478 11/04/10 19:02 auto_snap Miscellaneous snapshot
  • 338477 11/04/10 19:02 auto_snap Hall A Tools snapshot
  • 338476 11/04/10 19:02 auto_snap Compton Polarimeter snapshot
  • 338475 11/04/10 19:02 auto_snap Beam Position and Current snapshot
  • 338474 11/04/10 19:02 a-onl End_of_Run_8319
  • 338473 11/04/10 19:01 a-onl Start_Run_8319,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=DVCS DAQ checkout E14
  • 338472 11/04/10 18:56 Carlos Calorimeter ON
  • 338471 11/04/10 18:56 a-onl End_of_Run_8318
  • 338470 11/04/10 18:27 A. Camsonne FPGA programmers assignement and firware version E14 loaded
  • 338469 11/04/10 18:27 A. Camsonne FPGA programmers assignement and firware version E14 loaded
  • 338468 11/04/10 17:12 adaq Helicity checks II
  • 338467 11/04/10 16:02 adaq Beam Time Accounting
  • 338466 11/04/10 15:53 Andrew Tech Checklist
  • 338465 11/04/10 15:51 a-onl End_of_Run_8317
  • 338464 11/04/10 15:38 a-onl Start_Run_8317,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=helicity test: trigerring on the Helicity signal
  • 338463 11/04/10 15:33 Hisham Target status
  • 338462 11/04/10 15:15 a-onl End_of_Run_8316
  • 338461 11/04/10 14:43 a-onl Start_Run_8316,Run_type=Other,target_type=No Target, Empty,beam_e0=4.4,beam_raster=off,beam_current=0,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=helicity test: trigerring on the Helicity signal
  • 338460 11/04/10 13:16 a-onl End_of_Run_8315
  • 338459 11/04/10 12:22 Carlos Calorimeter status
  • 338458 11/04/10 12:22 a-onl Start_Run_8315,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Chek Front end busy - Test validation
  • 338457 11/04/10 12:21 A. Camsonne Switch to validation mode for ARS
  • 338456 11/04/10 12:18 a-onl End_of_Run_8314
  • 338455 11/04/10 12:18 a-onl Start_Run_8314,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Chek Front end busy - Test validation
  • 338454 11/04/10 12:17 a-onl End_of_Run_8313
  • 338453 11/04/10 12:12 a-onl Start_Run_8313,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Chek Front end busy - Test validation
  • 338452 11/04/10 12:07 a-onl End_of_Run_8312
  • 338451 11/04/10 11:57 auto_snap Radiation and Ozone snapshot
  • 338450 11/04/10 11:57 auto_snap Miscellaneous snapshot
  • 338449 11/04/10 11:57 auto_snap Hall A Tools snapshot
  • 338448 11/04/10 11:57 auto_snap Compton Polarimeter snapshot
  • 338447 11/04/10 11:57 auto_snap Beam Position and Current snapshot
  • 338446 11/04/10 11:56 A. Camsonne VME ADC ORed with Fastbus Frontend busy
  • 338445 11/04/10 11:51 a-onl Start_Run_8311,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Chek Front end busy
  • 338444 11/04/10 11:20 R. Michaels helicity checks, part I
  • 338443 11/04/10 11:02 auto_snap Radiation and Ozone snapshot
  • 338442 11/04/10 11:02 auto_snap Miscellaneous snapshot
  • 338441 11/04/10 11:02 auto_snap Hall A Tools snapshot
  • 338440 11/04/10 11:02 auto_snap Compton Polarimeter snapshot
  • 338439 11/04/10 11:02 auto_snap Beam Position and Current snapshot
  • 338438 11/04/10 10:04 hisham target status
  • 338437 11/04/10 09:51 R. Michaels no helicity errors at present
  • 338436 11/04/10 09:51 a-onl End_of_Run_8310
  • 338435 11/04/10 09:47 a-onl Start_Run_8310,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ... whatever
  • 338434 11/04/10 09:43 R. Michaels helicity board changed !
  • 338433 11/04/10 09:34 a-onl Start_Run_8309,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=test ... whatever
  • 338432 11/04/10 08:52 R. Michaels ABU fixed again (i hope)
  • 338431 11/04/10 08:34 Hisham Compton HV is off
  • 338430 11/04/10 08:34 Hisham Target moved to empty
  • 338429 11/04/10 08:23 Carlos Calorimeter HV off
  • 338428 11/04/10 08:12 a-onl End_of_Run_8308
  • 338427 11/04/10 08:04 Ole Hansen BTA still not fixed?
  • 338426 11/04/10 08:02 adaq Beam Time Accounting
  • 338425 11/04/10 08:01 Ole Hansen Owl Shift Summary
  • 338424 11/04/10 07:54 zana roc, RunControl, Scaler run 8308
  • 338423 11/04/10 07:13 a-onl Start_Run_8308,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338422 11/04/10 07:12 a-onl End_of_Run_8307
  • 338421 11/04/10 07:00 zana roc, RunControl, Scaler run 8307
  • 338420 11/04/10 06:48 R. Michaels auto BANU fixed now, I think
  • 338419 11/04/10 06:18 lselvy Target Update
  • 338418 11/04/10 05:59 a-onl Start_Run_8307,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338417 11/04/10 05:58 a-onl End_of_Run_8306
  • 338416 11/04/10 05:24 Ole Hansen Compton improved though could be better
  • 338415 11/04/10 05:00 lselvy Compton Looks Strange Fig
  • 338414 11/04/10 04:59 a-onl Start_Run_8306,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338413 11/04/10 04:58 a-onl End_of_Run_8305
  • 338412 11/04/10 04:53 Ole Hansen Compton looks strange
  • 338411 11/04/10 04:38 zana roc, RunControl, Scaler run 8305
  • 338410 11/04/10 04:17 Ole Hansen Excessive BANU in BTA
  • 338409 11/04/10 04:07 a-onl Start_Run_8305,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338408 11/04/10 04:06 a-onl End_of_Run_8304
  • 338407 11/04/10 03:57 auto_snap Radiation and Ozone snapshot
  • 338406 11/04/10 03:57 auto_snap Miscellaneous snapshot
  • 338405 11/04/10 03:57 auto_snap Hall A Tools snapshot
  • 338404 11/04/10 03:56 auto_snap Compton Polarimeter snapshot
  • 338403 11/04/10 03:56 auto_snap Beam Position and Current snapshot
  • 338402 11/04/10 03:49 zana roc, RunControl and Scaler run 8304
  • 338401 11/04/10 03:15 a-onl Start_Run_8304,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338400 11/04/10 03:14 a-onl End_of_Run_8303
  • 338399 11/04/10 03:04 zana roc, RunControl, Scaler run 8303
  • 338398 11/04/10 03:02 auto_snap Radiation and Ozone snapshot
  • 338397 11/04/10 03:02 auto_snap Miscellaneous snapshot
  • 338396 11/04/10 03:02 auto_snap Hall A Tools snapshot
  • 338395 11/04/10 03:02 auto_snap Compton Polarimeter snapshot
  • 338394 11/04/10 03:02 auto_snap Beam Position and Current snapshot
  • 338393 11/04/10 02:24 a-onl Start_Run_8303,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, ps5=1 kin 3 low energy
  • 338392 11/04/10 02:23 Ole Hansen L-HRS efficiency check
  • 338391 11/04/10 02:21 a-onl End_of_Run_8302
  • 338390 11/04/10 01:52 a-onl Start_Run_8302,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Left HRS Efficiency Check, ps3=1 ps4=1
  • 338389 11/04/10 01:47 a-onl End_of_Run_8301
  • 338388 11/04/10 01:35 zana roc, RunControl, Scaler run 8301
  • 338387 11/04/10 00:59 a-onl Start_Run_8301,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338386 11/04/10 00:58 a-onl End_of_Run_8300
  • 338385 11/04/10 00:57 zana roc, RunControl and Scaler run 8300
  • 338384 11/04/10 00:09 a-onl Start_Run_8300,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338383 11/04/10 00:09 a-onl End_of_Run_8299
  • 338382 11/04/10 00:07 adaq roc, RunControl and Scaler run 8299
  • 338381 11/04/10 00:00 adaq Beam Time Accounting
  • Wednesday

  • 338380 11/03/10 23:54 Yi Qiang Swing Shift Summary
  • 338379 11/03/10 23:33 a-onl Start_Run_8299,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338378 11/03/10 23:32 a-onl End_of_Run_8298
  • 338377 11/03/10 23:21 M Friend HWP change?
  • 338376 11/03/10 22:39 a-onl Start_Run_8298,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338375 11/03/10 22:38 pzhu roc,rc,scaler screen for run 8297
  • 338374 11/03/10 22:36 a-onl End_of_Run_8297
  • 338373 11/03/10 22:13 jroche Helicity sign check
  • 338372 11/03/10 21:59 M Friend Compton polarization history plots
  • 338371 11/03/10 21:51 Yi Qiang ADC Readback of PbF2 Calorimeter
  • 338370 11/03/10 21:44 a-onl Start_Run_8297,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338369 11/03/10 21:44 pzhu roc,rc,scaler screen for run 8296
  • 338368 11/03/10 21:42 a-onl End_of_Run_8296
  • 338367 11/03/10 20:34 a-onl Start_Run_8296,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338366 11/03/10 20:32 pzhu roc,rc,scaler screen for run 8295
  • 338365 11/03/10 20:31 a-onl End_of_Run_8295
  • 338364 11/03/10 19:56 auto_snap Radiation and Ozone snapshot
  • 338363 11/03/10 19:56 auto_snap Miscellaneous snapshot
  • 338362 11/03/10 19:56 auto_snap Hall A Tools snapshot
  • 338361 11/03/10 19:56 auto_snap Compton Polarimeter snapshot
  • 338360 11/03/10 19:56 auto_snap Beam Position and Current snapshot
  • 338359 11/03/10 19:32 a-onl Start_Run_8295,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338358 11/03/10 19:32 a-onl End_of_Run_8289
  • 338357 11/03/10 19:31 pzhu roc,rc,scaler screen for run 8294
  • 338356 11/03/10 19:29 a-onl End_of_Run_8294
  • 338355 11/03/10 19:02 auto_snap Radiation and Ozone snapshot
  • 338354 11/03/10 19:02 auto_snap Miscellaneous snapshot
  • 338353 11/03/10 19:01 auto_snap Hall A Tools snapshot
  • 338352 11/03/10 19:01 auto_snap Compton Polarimeter snapshot
  • 338351 11/03/10 19:01 auto_snap Beam Position and Current snapshot
  • 338350 11/03/10 18:36 a-onl Start_Run_8294,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338349 11/03/10 18:23 pzhu junk run of 8293
  • 338348 11/03/10 18:22 a-onl Start_Run_8293,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=pedestal run
  • 338347 11/03/10 18:22 a-onl End_of_Run_8293
  • 338346 11/03/10 18:20 pzhu roc,rc,scaler screen for pedestal run 8292
  • 338345 11/03/10 18:18 a-onl End_of_Run_8292
  • 338344 11/03/10 17:55 a-onl Start_Run_8292,Run_type=Pedestal,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=pedestal run
  • 338343 11/03/10 17:47 pzhu roc,rc,scaler screen for run 8291
  • 338342 11/03/10 17:45 a-onl End_of_Run_8291
  • 338341 11/03/10 17:33 Carlos Golden run updated for detectorL
  • 338340 11/03/10 17:04 a-onl Start_Run_8291,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338339 11/03/10 16:58 pzhu scaler screen for run 8290
  • 338338 11/03/10 16:54 pzhu roc,rc screens for run 8290
  • 338337 11/03/10 16:52 a-onl End_of_Run_8290
  • 338336 11/03/10 16:06 N. Muangma target screen
  • 338334 11/03/10 16:00 Saha Shift Summary
  • 338335 11/03/10 15:59 adaq Beam Time Accounting
  • 338333 11/03/10 15:45 a-onl Start_Run_8290,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338332 11/03/10 15:41 Punjabi roc,rc, scaler screens for run 8289
  • 338331 11/03/10 14:54 a-onl Start_Run_8289,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338330 11/03/10 14:31 a-onl End_of_Run_8288
  • 338329 11/03/10 13:55 a-onl Start_Run_8288,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338328 11/03/10 13:54 a-onl End_of_Run_8287
  • 338327 11/03/10 13:50 Punjabi rc,roc,scaler s for run 8287
  • 338326 11/03/10 13:06 a-onl Start_Run_8287,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338325 11/03/10 13:05 a-onl End_of_Run_8286
  • 338324 11/03/10 13:05 Punjabi roc,rc,scaler screens for run 8286
  • 338323 11/03/10 12:18 a-onl Start_Run_8286,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338322 11/03/10 12:17 a-onl End_of_Run_8285
  • 338321 11/03/10 12:16 Punjabi roc, scaler, rc screens
  • 338320 11/03/10 11:56 auto_snap Radiation and Ozone snapshot
  • 338319 11/03/10 11:56 auto_snap Miscellaneous snapshot
  • 338318 11/03/10 11:56 auto_snap Hall A Tools snapshot
  • 338317 11/03/10 11:56 auto_snap Compton Polarimeter snapshot
  • 338316 11/03/10 11:56 auto_snap Beam Position and Current snapshot
  • 338315 11/03/10 11:38 a-onl Start_Run_8285,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338314 11/03/10 11:37 a-onl End_of_Run_8284
  • 338313 11/03/10 11:35 Punjabi end of run 8284 roc,rc, scaler screens
  • 338312 11/03/10 11:01 auto_snap Radiation and Ozone snapshot
  • 338311 11/03/10 11:01 auto_snap Miscellaneous snapshot
  • 338310 11/03/10 11:01 auto_snap Hall A Tools snapshot
  • 338309 11/03/10 11:01 auto_snap Compton Polarimeter snapshot
  • 338308 11/03/10 11:01 auto_snap Beam Position and Current snapshot
  • 338307 11/03/10 10:55 a-onl Start_Run_8284,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338306 11/03/10 10:54 a-onl End_of_Run_8283
  • 338305 11/03/10 10:54 Punjabi ROC, RC, Scaler screens for run 8283
  • 338304 11/03/10 10:21 Punjabi VDC and mom spectra
  • 338303 11/03/10 10:16 a-onl Start_Run_8283,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338302 11/03/10 10:15 a-onl End_of_Run_8282
  • 338301 11/03/10 09:35 a-onl Start_Run_8282,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338300 11/03/10 09:35 a-onl End_of_Run_8281
  • 338299 11/03/10 09:34 Punjabi end od run 8281 RC, Scaler, Roc screens
  • 338298 11/03/10 08:52 a-onl Start_Run_8281,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338297 11/03/10 08:52 a-onl End_of_Run_8280
  • 338296 11/03/10 08:44 Punjabi RC, Scaler, ROC screens
  • 338295 11/03/10 08:10 a-onl Start_Run_8280,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338294 11/03/10 08:10 a-onl End_of_Run_8279
  • 338293 11/03/10 08:03 zana roc, RunControl, Scaler run 8279
  • 338292 11/03/10 08:01 adaq Beam Time Accounting
  • 338291 11/03/10 07:59 Ole Hansen Owl Shift Summary
  • 338290 11/03/10 07:30 a-onl Start_Run_8279,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338289 11/03/10 07:30 a-onl End_of_Run_8278
  • 338288 11/03/10 07:27 zana roc, RunControl, Scaler run 8278
  • 338287 11/03/10 06:57 Ole Hansen Resuming production
  • 338286 11/03/10 06:56 a-onl Start_Run_8278,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338285 11/03/10 06:48 lselvy LH2 Target
  • 338284 11/03/10 06:30 Ole Hansen Beam back
  • 338283 11/03/10 06:17 lselvy Target Update
  • 338282 11/03/10 05:55 lselvy Empty Target
  • 338281 11/03/10 04:44 lselvy No Beam
  • 338280 11/03/10 04:36 a-onl End_of_Run_8277
  • 338279 11/03/10 04:17 a-onl Start_Run_8277,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338278 11/03/10 04:16 a-onl End_of_Run_8276
  • 338277 11/03/10 04:14 zana roc, RunControl, Scaler run 8276
  • 338276 11/03/10 03:56 auto_snap Radiation and Ozone snapshot
  • 338275 11/03/10 03:56 auto_snap Miscellaneous snapshot
  • 338274 11/03/10 03:56 auto_snap Hall A Tools snapshot
  • 338273 11/03/10 03:56 auto_snap Compton Polarimeter snapshot
  • 338272 11/03/10 03:56 auto_snap Beam Position and Current snapshot
  • 338271 11/03/10 03:40 a-onl Start_Run_8276,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338270 11/03/10 03:39 a-onl End_of_Run_8275
  • 338269 11/03/10 03:37 zana roc, RunControl, Scaler run 8275
  • 338268 11/03/10 03:05 a-onl Start_Run_8275,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338267 11/03/10 03:04 a-onl End_of_Run_8274
  • 338266 11/03/10 03:03 zana roc, RunControl, Scaler run 8274
  • 338265 11/03/10 03:01 auto_snap Radiation and Ozone snapshot
  • 338264 11/03/10 03:01 auto_snap Miscellaneous snapshot
  • 338263 11/03/10 03:01 auto_snap Hall A Tools snapshot
  • 338262 11/03/10 03:01 auto_snap Compton Polarimeter snapshot
  • 338261 11/03/10 03:01 auto_snap Beam Position and Current snapshot
  • 338260 11/03/10 02:29 a-onl Start_Run_8274,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338259 11/03/10 02:29 a-onl End_of_Run_8273
  • 338258 11/03/10 02:28 zana roc, RunControl and Scaler 8272
  • 338257 11/03/10 01:52 a-onl Start_Run_8273,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338256 11/03/10 01:51 a-onl End_of_Run_8272
  • 338255 11/03/10 01:49 zana roc, RunControl, Scaler run 8272
  • 338254 11/03/10 01:16 a-onl Start_Run_8272,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338253 11/03/10 01:15 a-onl End_of_Run_8271
  • 338252 11/03/10 01:14 zana roc RunControl Scaler run 8271
  • 338251 11/03/10 00:41 a-onl Start_Run_8271,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338250 11/03/10 00:40 a-onl End_of_Run_8270
  • 338249 11/03/10 00:40 zana roc RunControl and scaler run 8270
  • 338248 11/03/10 00:34 M Friend Beam charge asymmetry from Compton DAQ
  • 338247 11/03/10 00:10 adaq Beam Time Accounting
  • 338246 11/03/10 00:05 a-onl Start_Run_8270,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338245 11/03/10 00:04 a-onl End_of_Run_8269
  • 338244 11/03/10 00:04 zana roc, RunControl and Scaler run 8269
  • Tuesday

  • 338243 11/02/10 23:59 Pawel Swing shift summary
  • 338242 11/02/10 23:58 Pawel Swing shift summary
  • 338241 11/02/10 23:28 a-onl Start_Run_8269,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338240 11/02/10 23:27 a-onl End_of_Run_8268
  • 338239 11/02/10 23:26 A. Tobias ROCs, RunControl and Scalars for run 8268 - production
  • 338238 11/02/10 22:58 Pawel calorimeter currents
  • 338237 11/02/10 22:58 mcanan Target Screens
  • 338236 11/02/10 22:39 a-onl Start_Run_8268,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338235 11/02/10 22:38 a-onl End_of_Run_8267
  • 338234 11/02/10 22:36 A. Tobias ROCs, RunControl and Scalars - production
  • 338233 11/02/10 21:52 a-onl Start_Run_8267,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338232 11/02/10 21:51 Pawel Hall B normalization run failed
  • 338231 11/02/10 21:21 a-onl End_of_Run_8266
  • 338230 11/02/10 21:21 Pawel Down for Hall B normalization run
  • 338229 11/02/10 21:17 a-onl Start_Run_8266,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338228 11/02/10 21:17 a-onl End_of_Run_8265
  • 338227 11/02/10 21:14 A. Tobias ROCs, RunControl and Scalars for run 8265 - production
  • 338226 11/02/10 20:34 a-onl Start_Run_8265,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338225 11/02/10 20:32 a-onl End_of_Run_8264
  • 338224 11/02/10 20:32 A. Tobias ROCs, RunControl and Scalars for run 8264 - production
  • 338223 11/02/10 19:55 auto_snap Radiation and Ozone snapshot
  • 338222 11/02/10 19:55 auto_snap Miscellaneous snapshot
  • 338221 11/02/10 19:55 auto_snap Hall A Tools snapshot
  • 338220 11/02/10 19:55 auto_snap Compton Polarimeter snapshot
  • 338219 11/02/10 19:55 auto_snap Beam Position and Current snapshot
  • 338218 11/02/10 19:47 a-onl Start_Run_8264,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338217 11/02/10 19:46 a-onl End_of_Run_8263
  • 338216 11/02/10 19:45 A. Tobias ROCs, RunControl and Scalars for run 8263 - production
  • 338215 11/02/10 19:17 cshloka runlist updated in dvcs wiki
  • 338214 11/02/10 19:02 a-onl Start_Run_8263,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338213 11/02/10 19:01 a-onl End_of_Run_8262
  • 338212 11/02/10 19:01 auto_snap Radiation and Ozone snapshot
  • 338211 11/02/10 19:01 auto_snap Miscellaneous snapshot
  • 338210 11/02/10 19:01 auto_snap Hall A Tools snapshot
  • 338209 11/02/10 19:01 auto_snap Compton Polarimeter snapshot
  • 338208 11/02/10 19:00 auto_snap Beam Position and Current snapshot
  • 338207 11/02/10 18:42 A. Tobias ROCs, RunControl and Scalars for run 8262 - production
  • 338206 11/02/10 18:29 kelleher Large Charge Asymmetry Runs
  • 338205 11/02/10 17:57 a-onl Start_Run_8262,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production, TrigDVCSBuff ps5=1 kin3 low energy
  • 338204 11/02/10 17:52 a-onl End_of_Run_8261
  • 338203 11/02/10 17:50 a-onl Start_Run_8261,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Large Charge Asymmetry Meas: TrigDVCSBuff ps5=1 kin3 low energy
  • 338202 11/02/10 17:44 A. Tobias ROCs, RunControl and Scalars for run 8260 (Large Charge Asymmetry Measurement)
  • 338201 11/02/10 17:43 R. Michaels re: disk full problem and amelioration.
  • 338200 11/02/10 17:24 O.Glamazdin Beam charge asymmetry measurement with Moller
  • 338199 11/02/10 17:15 a-onl Start_Run_8260,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Large Charge Asymmetry Meas: TrigDVCSBuff ps5=1 kin3 low energy
  • 338198 11/02/10 17:08 a-onl End_of_Run_8259
  • 338197 11/02/10 17:06 a-onl End_of_Run_8259
  • 338196 11/02/10 17:04 A. Tobias Pedestals may have shifted. This is a FATAL ERROR. Please PAGE expert inmediately!!
  • 338195 11/02/10 17:02 A. Tobias ROCs, RunControl and Scalars for run 8259 (Large Charge Asymmetry Measurement)
  • 338194 11/02/10 16:43 Pawel Nadel- Large charge asymmetry test plan in progress
  • 338193 11/02/10 16:34 a-onl Start_Run_8259,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338192 11/02/10 16:32 a-onl End_of_Run_8258
  • 338191 11/02/10 16:29 A. Tobias ROCs, RunControl & Scalars for run 8258
  • 338190 11/02/10 16:02 Hisham Day shift summary
  • 338189 11/02/10 16:01 adaq Beam Time Accounting
  • 338188 11/02/10 15:52 a-onl Start_Run_8258,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338187 11/02/10 15:51 a-onl End_of_Run_8257
  • 338186 11/02/10 15:51 jpchen DAQ for run 8256
  • 338185 11/02/10 15:44 jpchen Target screens
  • 338184 11/02/10 15:39 jpchen DAQ problem: disk full.
  • 338183 11/02/10 15:17 Hisham Compton rates
  • 338182 11/02/10 15:06 a-onl Start_Run_8257,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338181 11/02/10 15:03 a-onl End_of_Run_8256
  • 338180 11/02/10 14:31 a-onl Start_Run_8256,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338179 11/02/10 14:02 jpchen DAQ for run8246
  • 338178 11/02/10 13:46 Hisham Beam is back
  • 338177 11/02/10 13:42 Hisham Beam is down
  • 338176 11/02/10 13:05 a-onl Start_Run_8246,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338175 11/02/10 13:05 a-onl Start_Run_8240,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338174 11/02/10 13:04 Hisham ADC Readback
  • 338173 11/02/10 13:04 a-onl End_of_Run_8245
  • 338172 11/02/10 13:04 jpchen DAQ for run 8245
  • 338171 11/02/10 12:17 a-onl Start_Run_8245,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338170 11/02/10 12:15 a-onl End_of_Run_8244
  • 338169 11/02/10 12:14 jpchen DAQ for run 8244
  • 338168 11/02/10 11:59 Hisham Compton rate is back to normal
  • 338167 11/02/10 11:56 A. Camsonne Compton asymmetry changed signed during the night
  • 338166 11/02/10 11:55 auto_snap Radiation and Ozone snapshot
  • 338165 11/02/10 11:55 auto_snap Miscellaneous snapshot
  • 338164 11/02/10 11:55 auto_snap Hall A Tools snapshot
  • 338163 11/02/10 11:55 auto_snap Compton Polarimeter snapshot
  • 338162 11/02/10 11:55 auto_snap Beam Position and Current snapshot
  • 338161 11/02/10 11:42 R. Michaels mod to halla_scaler_process.tcl
  • 338160 11/02/10 11:23 a-onl Start_Run_8244,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338159 11/02/10 11:23 a-onl End_of_Run_8243
  • 338158 11/02/10 11:22 jpchen DAQ for run 8243
  • 338157 11/02/10 11:00 auto_snap Radiation and Ozone snapshot
  • 338156 11/02/10 11:00 auto_snap Miscellaneous snapshot
  • 338155 11/02/10 11:00 auto_snap Hall A Tools snapshot
  • 338154 11/02/10 11:00 auto_snap Compton Polarimeter snapshot
  • 338153 11/02/10 11:00 auto_snap Beam Position and Current snapshot
  • 338152 11/02/10 10:53 A. Camsonne Major tune required to fix the Compton rates - continue takeing data this way
  • 338151 11/02/10 10:53 Carlos Golden run updated for detectorL
  • 338150 11/02/10 10:40 jpchen deadtime negative value
  • 338149 11/02/10 10:37 a-onl Start_Run_8243,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338148 11/02/10 10:36 A. Camsonne Compton laser off for background checks
  • 338147 11/02/10 10:36 a-onl End_of_Run_8242
  • 338146 11/02/10 10:36 jpchen DAQ for run 8242
  • 338145 11/02/10 10:27 Carlos Missing mass squared
  • 338144 11/02/10 10:21 Hisham Compton: run 23723
  • 338143 11/02/10 10:05 Hisham High rates for Compton: run 23722
  • 338142 11/02/10 09:56 A. Camsonne Random pulser on for DVCS trigger
  • 338141 11/02/10 09:51 a-onl Start_Run_8242,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy, added random trigger before run
  • 338140 11/02/10 09:48 a-onl End_of_Run_8241
  • 338139 11/02/10 09:47 jpchen QAD screens for run 8241
  • 338138 11/02/10 09:12 jpchen Gas Cherenkov after threshold change
  • 338137 11/02/10 09:06 Hisham Beam is Up
  • 338136 11/02/10 09:05 a-onl Start_Run_8241,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy, beam half-wave plate flipped before run
  • 338135 11/02/10 09:03 A. Camsonne Half wave plate was out requested to put it in
  • 338134 11/02/10 09:00 a-onl End_of_Run_8240
  • 338133 11/02/10 08:38 Hisham Beam is down
  • 338132 11/02/10 08:33 adaq QAD
  • 338131 11/02/10 08:28 a-onl End_of_Run_8239
  • 338129 11/02/10 08:02 Ole Hansen Owl Shift Summary
  • 338130 11/02/10 08:01 adaq Beam Time Accounting
  • 338128 11/02/10 07:45 a-onl Start_Run_8239,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338127 11/02/10 07:44 a-onl End_of_Run_8238
  • 338126 11/02/10 07:43 zana scaler run 8238
  • 338125 11/02/10 07:43 adaq roc and RunControl run 8238
  • 338124 11/02/10 07:08 a-onl Start_Run_8238,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338123 11/02/10 07:07 a-onl End_of_Run_8237
  • 338122 11/02/10 07:06 zana scaler run 8237
  • 338121 11/02/10 07:06 adaq roc and RunControl run 8237
  • 338120 11/02/10 06:25 a-onl Start_Run_8237,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338119 11/02/10 06:24 a-onl End_of_Run_8236
  • 338118 11/02/10 06:24 zana scaler run 8236
  • 338117 11/02/10 06:23 adaq roc and RunControl run 8236
  • 338116 11/02/10 05:59 lselvy Target Update
  • 338115 11/02/10 05:36 a-onl Start_Run_8236,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338114 11/02/10 05:36 a-onl End_of_Run_8235
  • 338113 11/02/10 05:35 zana scaler run 8235
  • 338112 11/02/10 05:35 adaq roc and RunControl run 8235
  • 338111 11/02/10 04:58 a-onl Start_Run_8235,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338110 11/02/10 04:57 a-onl End_of_Run_8234
  • 338109 11/02/10 04:57 zana scaler run 8234
  • 338108 11/02/10 04:57 adaq roc and RunControl run 8234
  • 338107 11/02/10 04:19 a-onl Start_Run_8234,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338106 11/02/10 04:19 a-onl End_of_Run_8233
  • 338105 11/02/10 04:17 zana scaler run 8233
  • 338104 11/02/10 04:17 adaq roc and RunControl run 8233
  • 338103 11/02/10 03:55 auto_snap Radiation and Ozone snapshot
  • 338102 11/02/10 03:55 auto_snap Miscellaneous snapshot
  • 338101 11/02/10 03:55 auto_snap Hall A Tools snapshot
  • 338100 11/02/10 03:55 auto_snap Compton Polarimeter snapshot
  • 338099 11/02/10 03:55 auto_snap Beam Position and Current snapshot
  • 338098 11/02/10 03:41 a-onl Start_Run_8233,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338097 11/02/10 03:41 a-onl End_of_Run_8232
  • 338096 11/02/10 03:39 zana scaler run 8232
  • 338095 11/02/10 03:39 adaq roc and run control run 8232
  • 338094 11/02/10 03:01 a-onl Start_Run_8232,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338093 11/02/10 03:00 auto_snap Radiation and Ozone snapshot
  • 338092 11/02/10 03:00 auto_snap Miscellaneous snapshot
  • 338091 11/02/10 03:00 auto_snap Hall A Tools snapshot
  • 338090 11/02/10 03:00 a-onl End_of_Run_8231
  • 338089 11/02/10 03:00 auto_snap Compton Polarimeter snapshot
  • 338088 11/02/10 03:00 auto_snap Beam Position and Current snapshot
  • 338087 11/02/10 02:59 zana scaler run 8231
  • 338086 11/02/10 02:58 adaq roc and run control run 8231
  • 338085 11/02/10 02:20 a-onl Start_Run_8231,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338084 11/02/10 02:19 a-onl End_of_Run_8230
  • 338083 11/02/10 02:19 zana scaler run 8230
  • 338082 11/02/10 02:18 adaq roc and terminal run 8230
  • 338081 11/02/10 01:40 a-onl Start_Run_8230,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338080 11/02/10 01:39 a-onl End_of_Run_8229
  • 338079 11/02/10 01:38 zana scaler run 8229
  • 338078 11/02/10 01:38 adaq scaler and runcontrol run 8229
  • 338077 11/02/10 01:02 a-onl Start_Run_8229,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338076 11/02/10 01:02 a-onl End_of_Run_8226
  • 338075 11/02/10 01:01 a-onl End_of_Run_8228
  • 338074 11/02/10 00:58 zana scalers run 8228
  • 338073 11/02/10 00:57 adaq roc and run control run 8228
  • 338072 11/02/10 00:25 a-onl Start_Run_8228,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338071 11/02/10 00:24 a-onl End_of_Run_8227
  • 338070 11/02/10 00:23 zana scaler run 8227
  • 338069 11/02/10 00:22 zana roc and run control run 8227
  • 338068 11/02/10 00:01 S. Golge Swing Shift Summary
  • 338067 11/02/10 00:00 adaq Beam Time Accounting
  • Monday

  • 338066 11/01/10 23:45 a-onl Start_Run_8227,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338065 11/01/10 23:43 flay Run 8226 Trigger/ROC Snapshots
  • 338064 11/01/10 23:03 a-onl Start_Run_8226,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338063 11/01/10 23:03 a-onl End_of_Run_8225
  • 338062 11/01/10 23:02 flay Run 8225 Trigger/ROC Snapshots
  • 338061 11/01/10 22:50 M Friend Compton flags file changed
  • 338060 11/01/10 22:40 A. Tobias Cavity Lock: Compton Stripchart Snapshot
  • 338059 11/01/10 22:23 a-onl Start_Run_8225,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338058 11/01/10 22:22 a-onl End_of_Run_8224
  • 338057 11/01/10 22:21 flay Run 8224 Trigger/ROC Snapshots
  • 338056 11/01/10 21:42 a-onl Start_Run_8224,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338055 11/01/10 21:42 a-onl End_of_Run_8223
  • 338054 11/01/10 21:41 flay Run 8223 Trigger/ROC Snapshots
  • 338053 11/01/10 21:11 A. Camsonne Added sync plots in calo.cfg and update golden run
  • 338052 11/01/10 21:04 a-onl Start_Run_8223,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338051 11/01/10 21:03 a-onl End_of_Run_8222
  • 338050 11/01/10 20:53 A. Camsonne adaqh1 back online
  • 338049 11/01/10 20:46 a-onl Start_Run_8222,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338048 11/01/10 20:45 a-onl End_of_Run_8221
  • 338047 11/01/10 20:40 A. Tobias Target Screens
  • 338046 11/01/10 20:39 A. Tobias Target Status
  • 338045 11/01/10 20:06 a-onl Start_Run_8221,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338044 11/01/10 20:04 a-onl End_of_Run_8220
  • 338043 11/01/10 20:03 flay Run 8220 Trigger/ROC Snapshots
  • 338042 11/01/10 19:55 auto_snap Radiation and Ozone snapshot
  • 338041 11/01/10 19:55 auto_snap Miscellaneous snapshot
  • 338040 11/01/10 19:54 auto_snap Hall A Tools snapshot
  • 338039 11/01/10 19:54 auto_snap Compton Polarimeter snapshot
  • 338038 11/01/10 19:54 auto_snap Beam Position and Current snapshot
  • 338037 11/01/10 19:27 a-onl Start_Run_8220,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338036 11/01/10 19:25 a-onl End_of_Run_8219
  • 338035 11/01/10 19:24 flay Run 8219 Trigger/ROC Snapshots
  • 338034 11/01/10 19:00 auto_snap Radiation and Ozone snapshot
  • 338033 11/01/10 19:00 auto_snap Miscellaneous snapshot
  • 338032 11/01/10 19:00 auto_snap Hall A Tools snapshot
  • 338031 11/01/10 19:00 auto_snap Compton Polarimeter snapshot
  • 338030 11/01/10 19:00 auto_snap Beam Position and Current snapshot
  • 338029 11/01/10 18:40 a-onl Start_Run_8219,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338028 11/01/10 18:39 a-onl End_of_Run_8218
  • 338027 11/01/10 17:57 a-onl Start_Run_8218,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338026 11/01/10 17:56 a-onl End_of_Run_8217
  • 338025 11/01/10 17:16 a-onl Start_Run_8217,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338024 11/01/10 17:15 a-onl End_of_Run_8216
  • 338023 11/01/10 17:14 flay Run 8216 Trigger/ROC Snapshots
  • 338022 11/01/10 16:56 flay Run 8216 Classification
  • 338021 11/01/10 16:51 flay Run 8215: Gas Cherenkov TDC Peaks
  • 338020 11/01/10 16:37 a-onl Start_Run_8216,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338019 11/01/10 16:26 A. Camsonne rebooting adaql2
  • 338018 11/01/10 16:25 flay Run 8215 Trigger/ROC Snapshots
  • 338017 11/01/10 16:13 A. Camsonne Compton fasScaler.crl file updated to use 120 Hz helicity flip rate
  • 338016 11/01/10 16:07 adaq Beam Time Accounting
  • 338015 11/01/10 16:04 flay Run 8214: Cherenkov TDC Double Peaks
  • 338014 11/01/10 16:01 hyde Day Shift Summary
  • 338013 11/01/10 15:55 Hisham Target status snapshot
  • 338012 11/01/10 15:35 hyde Cerenkov Efficiency
  • 338011 11/01/10 15:27 a-onl Start_Run_8215,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production TrigDVCSBuff ps5=1 kin3 low energy
  • 338010 11/01/10 15:24 a-onl End_of_Run_8214
  • 338009 11/01/10 15:17 a-onl Start_Run_8214,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Normalization LeftHRS ps5=1
  • 338008 11/01/10 15:01 a-onl Start_Run_8213,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Normalization LeftHRS ps3=ps4=1
  • 338007 11/01/10 14:53 a-onl End_of_Run_8210
  • 338006 11/01/10 14:52 a-onl Start_Run_8211,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ARS with validation
  • 338005 11/01/10 14:52 a-onl End_of_Run_8211
  • 338004 11/01/10 14:52 a-onl Start_Run_8210,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ARS with validation
  • 338003 11/01/10 14:50 hyde Wien angle +5deg
  • 338002 11/01/10 14:39 a-onl End_of_Run_8209
  • 338001 11/01/10 14:17 a-onl Start_Run_8207,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ARS with validation
  • 338000 11/01/10 14:17 a-onl End_of_Run_8208
  • 337999 11/01/10 14:16 a-onl End_of_Run_8207
  • 337998 11/01/10 14:15 a-onl Start_Run_8206,Run_type=Other,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Test ARS with validation
  • 337997 11/01/10 14:14 a-onl End_of_Run_8206
  • 337996 11/01/10 14:13 a-onl End_of_Run_8204
  • 337995 11/01/10 14:04 hyde helicity = 120 Hz
  • 337994 11/01/10 13:59 hyde controlled access for Cerenkov
  • 337993 11/01/10 13:58 a-onl End_of_Run_8205
  • 337992 11/01/10 13:51 a-onl Start_Run_8205,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Normalization 4 pass Kin 3 LeftHRS ps5=1
  • 337991 11/01/10 13:50 a-onl End_of_Run_8201
  • 337990 11/01/10 13:34 a-onl Start_Run_8204,Run_type=Calibration,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Normalization 4 pass Kin 3 LeftHRS ps3=ps4=1
  • 337989 11/01/10 13:31 a-onl End_of_Run_8203
  • 337988 11/01/10 13:25 a-onl Start_Run_8203,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337987 11/01/10 13:23 hyde Half wave plate out
  • 337986 11/01/10 13:19 a-onl End_of_Run_8202
  • 337985 11/01/10 13:18 Pawel run 8202
  • 337984 11/01/10 12:55 hyde Cerenkov inefficiency = 7%
  • 337983 11/01/10 12:41 hyde Cerenkov vs PR
  • 337982 11/01/10 12:11 R. Michaels info about livetime from TS registers
  • 337981 11/01/10 12:04 a-onl Start_Run_8202,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337980 11/01/10 12:03 a-onl End_of_Run_8200
  • 337979 11/01/10 12:01 Pawel run 8201
  • 337978 11/01/10 11:54 auto_snap Radiation and Ozone snapshot
  • 337977 11/01/10 11:54 auto_snap Miscellaneous snapshot
  • 337976 11/01/10 11:54 auto_snap Hall A Tools snapshot
  • 337975 11/01/10 11:54 auto_snap Compton Polarimeter snapshot
  • 337974 11/01/10 11:54 auto_snap Beam Position and Current snapshot
  • 337973 11/01/10 11:47 Hisham Run 23680: Coda crash at the end of the run
  • 337972 11/01/10 11:03 a-onl Start_Run_8201,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337971 11/01/10 11:00 Pawel run 8200 event recorder
  • 337970 11/01/10 11:00 auto_snap Radiation and Ozone snapshot
  • 337969 11/01/10 10:59 auto_snap Miscellaneous snapshot
  • 337968 11/01/10 10:59 auto_snap Hall A Tools snapshot
  • 337967 11/01/10 10:59 auto_snap Compton Polarimeter snapshot
  • 337966 11/01/10 10:59 auto_snap Beam Position and Current snapshot
  • 337965 11/01/10 10:59 Pawel run 8200
  • 337964 11/01/10 10:42 hyde Cerenkov Run 8161
  • 337963 11/01/10 10:21 Hisham Target status snapshot
  • 337962 11/01/10 10:20 a-onl Start_Run_8200,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337961 11/01/10 10:19 a-onl End_of_Run_8193
  • 337960 11/01/10 10:16 a-onl End_of_Run_8199
  • 337959 11/01/10 10:15 Pawel run 8199
  • 337958 11/01/10 10:09 R. Michaels scaler server mods.
  • 337957 11/01/10 10:02 hyde Cerenkov run 8161
  • 337956 11/01/10 09:36 Hisham Compton rate flatlined -rebooted the Compton DAQ
  • 337955 11/01/10 09:31 a-onl Start_Run_8199,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337954 11/01/10 09:30 a-onl End_of_Run_8198
  • 337953 11/01/10 09:30 Pawel run 8198
  • 337952 11/01/10 09:29 hyde compton run 23677
  • 337951 11/01/10 08:58 hyde Compton
  • 337950 11/01/10 08:37 a-onl Start_Run_8198,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337949 11/01/10 08:35 Pawel run 8197 screens
  • 337948 11/01/10 08:33 a-onl End_of_Run_8197
  • 337947 11/01/10 08:32 hyde Compton flatlined
  • 337945 11/01/10 08:02 Ole Hansen Owl Shift Summary
  • 337946 11/01/10 08:02 adaq Beam Time Accounting
  • 337944 11/01/10 07:55 a-onl Start_Run_8197,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337943 11/01/10 07:54 a-onl End_of_Run_8196
  • 337942 11/01/10 07:53 itard entries end of run
  • 337941 11/01/10 07:07 Carlos ROCs screenshots
  • 337940 11/01/10 07:07 Ole Hansen Fast feedback - off?
  • 337939 11/01/10 06:54 a-onl Start_Run_8196,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337938 11/01/10 06:43 itard entries end of run
  • 337937 11/01/10 06:41 a-onl End_of_Run_8195
  • 337936 11/01/10 06:41 Ole Hansen Beam off
  • 337935 11/01/10 06:40 Ole Hansen CODA recovered
  • 337934 11/01/10 06:39 Ole Hansen Run 8194 probably junk
  • 337933 11/01/10 06:38 a-onl Start_Run_8195,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337932 11/01/10 06:35 a-onl End_of_Run_8194
  • 337931 11/01/10 06:33 itard coda crash
  • 337930 11/01/10 06:21 lselvy Target Update
  • 337929 11/01/10 06:04 a-onl Start_Run_8194,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337928 11/01/10 06:02 itard entries end of run
  • 337927 11/01/10 06:01 a-onl End_of_Run_8182
  • 337926 11/01/10 05:02 a-onl Start_Run_8193,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337925 11/01/10 05:01 itard entries end of run
  • 337924 11/01/10 05:00 a-onl End_of_Run_8192
  • 337923 11/01/10 04:00 a-onl Start_Run_8192,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337922 11/01/10 03:57 itard entries end of run
  • 337921 11/01/10 03:57 a-onl End_of_Run_8191
  • 337920 11/01/10 03:54 auto_snap Radiation and Ozone snapshot
  • 337919 11/01/10 03:54 auto_snap Miscellaneous snapshot
  • 337918 11/01/10 03:54 auto_snap Hall A Tools snapshot
  • 337917 11/01/10 03:54 auto_snap Compton Polarimeter snapshot
  • 337916 11/01/10 03:54 auto_snap Beam Position and Current snapshot
  • 337915 11/01/10 02:59 auto_snap Radiation and Ozone snapshot
  • 337914 11/01/10 02:59 auto_snap Miscellaneous snapshot
  • 337913 11/01/10 02:59 auto_snap Hall A Tools snapshot
  • 337912 11/01/10 02:59 auto_snap Compton Polarimeter snapshot
  • 337911 11/01/10 02:59 auto_snap Beam Position and Current snapshot
  • 337910 11/01/10 02:58 a-onl Start_Run_8191,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337909 11/01/10 02:55 itard entries end of runq
  • 337908 11/01/10 02:55 a-onl End_of_Run_8190
  • 337907 11/01/10 02:19 Ole Hansen Compton crash
  • 337906 11/01/10 01:36 a-onl Start_Run_8190,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337905 11/01/10 01:23 itard entries end of run
  • 337904 11/01/10 01:21 a-onl End_of_Run_8189
  • 337903 11/01/10 00:21 a-onl Start_Run_8189,Run_type=Production,target_type=LH2,beam_e0=4.4,beam_raster=off,beam_current=3,leftarm_p=1.494,leftarm_theta=31.82,leftarm_sieve=out,calo_theta=14.78,calo_distance=1.1,comment_text=Production 4 pass Kin 3 TrigDVCSBuff ps5=1
  • 337902 11/01/10 00:07 A. Camsonne Header updated for ROC18 for Run 8189 included
  • 337901 11/01/10 00:07 A. Camsonne Header updated for ROC18 for Run 8189 included
  • 337900 11/01/10 00:07 a-onl End_of_Run_8188
  • 337899 11/01/10 00:06 A. Tobias Other ROCs for run 8188
  • 337898 11/01/10 00:03 A. Tobias RunControl and Scalars for run 8188
  • 337897 11/01/10 00:02 A. Tobias ROCs for run 8188
  • 337896 11/01/10 00:00 adaq Beam Time Accounting

  • Earlier months at hallaweb.jlab.org/adaq (archive).  The HALOG was generated by version HA1.48
    If you have any comments or problems, please contact:
    Robert Michaels
    rom@jlab.org
    (757) 269-7410